### v1

# Accuracy of the Ensemble Kalman Filter in the Near-Linear Setting

Abstract
The filtering distribution captures the statistics of the state of a dynamical system from partial and noisy observations. Classical particle filters provably approximate this distribution in quite general settings; however they behave poorly for high dimensional problems, suffering weight collapse. This issue is circumvented by the ensemble Kalman filter which is an equal-weight interacting particle system. However, this finite particle system is only proven to approximate the true filter in the linear Gaussian case. In practice, however, it is applied in much broader settings; as a result, establishing its approximation properties more generally is important. There has been recent progress in the theoretical analysis of the algorithm, establishing stability and error estimates in non-Gaussian settings, but the assumptions on the dynamics and observation models rule out the unbounded vector fields that arise in practice and the analysis applies only to the mean field limit of the ensemble Kalman filter. The present work establishes error bounds between the filtering distribution and the finite particle ensemble Kalman filter when the model exhibits linear growth.
Keywords. Ensemble Kalman filter, Stochastic filtering, Weighted total variation metric, Stability estimates, Accuracy estimates, Near-Linear setting.
AMS subject classification. 60G35, 62F15, 65C35, 70F45, 93E11.
1 Introduction
1.1 Literature Review, Contributions and Outline
Algorithms for filtering employ noisy observations arising from a possibly random dynamical system to estimate the distribution of the state of the dynamical system conditional on the observations. The Kalman filter [kalman1960new] determines the filtering distribution exactly for linear Gaussian dynamics and observations. The extended Kalman filter was developed as an extension of the Kalman filtering technique to nonlinear problems and is based on a linearization approximation; see [jazwinski2007stochastic] and [anderson2012optimal]. The linearization approximation leads to an inexact distribution for nonlinear problems, and furthermore requires evaluation of covariance matrices, making the methodology impractical for large-scale geophysical applications [ghil]. Particle filters or sequential Monte Carlo methods [doucet2001sequential, chopin:20] offer an alternative methodology for nonlinear filtering problems, allowing recovery of the exact filtering distribution in the large particle limit. However, this class of methods scales poorly with dimension, and in particular suffer weight collapse, making their application to geophysical problems prohibitive; see for example [snyder2008obstacles, bickel2008sharp, rebeschini2015can, agapiou2017importance].
The ensemble Kalman filter was introduced in the seminal paper [evensen1994sequential]. Its success stems from the low-rank approximation of large covariances, cheaply computed using an ensemble of particles and allowing it to be deployed in geophysical applications. However analysis of its accuracy, in relation to the true filtering distribution, remains in its infancy. The papers [le2009large, mandel2011convergence] studied this issue in the linear Gaussian setting where the mean field limit of the Kalman filter is exact; they demonstrate that the ensemble Kalman filter may be viewed as an interacting particle system approximation of the mean field limit, and establish Monte Carlo type error bounds. The recent article [calvello2025ensemble] overviews the formulation of ensemble Kalman methods using mean field dynamical systems and provides a platform from which the analyses of [le2009large, mandel2011convergence] may be generalized beyond the linear Gaussian setting. In the recent paper [carrillo2022ensemble] the authors establish stability properties of the mean field ensemble Kalman filter and use them to prove accuracy of the filter in a near-Gaussian setting. However the paper does not consider finite particle approximations of the mean field, and the conditions on the dynamics-observation model require boundedness of the vector fields arising. In this paper we address both these issues, establishing error bounds between the finite particle ensemble Kalman filter and the true filtering distribution in settings where the dynamics and observation vector fields grow linearly. We make the following contributions.
  1. 1.
    Theorem 2.1 is a stability result for the mean field ensemble Kalman filter in the setting of dynamical models and Lipschitz obervation operators that grow at most linearly at infinity.
  2. 2.
    Theorem 2.3 quantifies the error between the mean field ensemble Kalman filter and the true filter in the setting of dynamical models and Lipschitz observation operators that are near-linear.
  3. 3.
    Theorem 2.4 quantifies the error between the finite particle ensemble Kalman filter (found as an interacting particle system approximation of the mean field) and the true filter in the setting of near-linear, Lipschitz dynamical models and linear observation operators.
In going beyond the work in [carrillo2022ensemble], the current paper simultaneously addresses a more applicable problem class, by allowing linear growth in the dynamics and observations operators, and confronts the substantial technical challenges which arise from doing so. Bounds on moments of the filtering distribution and mean field ensemble Kalman filter must be established; these bounds grow in the number of iterations which is in contrast to [carrillo2022ensemble] where the $L^{\infty}$ bounds on model and observation operator ensure a uniform bound on moments. A further challenge is presented by the need to establish stability bounds for the conditioning map, giving rise to the true filter, and the transport map giving rise to the ensemble Kalman filter. These results exhibit dependence on moments in the stability constants and require control of the growth given by the dynamics; this is again in contrast to [carrillo2022ensemble] where the $L^{\infty}$ bounds allow for a uniform control.
After discussing notation that will be used throughout the paper in Subsection 1.2, we introduce the filtering problem in Subsection 1.3. In Section 2 we outline the main results of the paper concerning the ensemble Kalman filter. In Subsection 2.1 we formulate the ensemble Kalman filter that we consider in this paper along with the relevant assumptions we will use in the analysis. We state a stability theorem for the mean field formulation of the ensemble Kalman filter in Subsection 2.2, hence Contribution 1. We leverage this result in Subsection 2.3 to derive a theorem quantifying the error between the mean field ensemble Kalman filter and the true filter, Contribution 2. Finally, in Subsection 2.4, we make use of the results of the previous subsections to state a theorem quantifying the error between the (finite particle) ensemble Kalman filter itself and the true filter, yielding Contribution 3. Various technical results, used in the proof of our three main theorems, may be found in the appendices. We conclude with closing remarks in Section 3.
1.2 Notation
[…]
$\|\mathfrak{f}-f\|_{L^{\infty}(^{n})}\leqslant r,\quad\|\mathfrak{g}-g\|_{L^{\infty}(^{d})}\leqslant r.$
We let $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ denote the $C^{0,1}$ semi-norm, namely the Lipschitz constant.
[…]
For a probability measure $\mu\in\mathcal{P}(^{n})$, the notation $\mathcal{M}_{q}(\mu)$ denotes the $q$th polynomial moment under the measure $\mu$, defined as
$\mathcal{M}_{q}(\mu)\coloneqq\int_{{}^{n}}|x|^{q}\mu(\mathrm{d}x).$ (1.1)
The notations $\mathcal{M}(\mu)$ and $\mathcal{C}(\mu)$ denote respectively the mean and covariance under $\mu$:
$\mathcal{M}(\mu)=\mu[x],\qquad\mathcal{C}(\mu)=\mu\Bigl{[}\bigl{(}x-\mathcal{M}(\mu)\bigr{)}\otimes\bigl{(}x-\mathcal{M}(\mu)\bigr{)}\Bigr{]}.$
The notation $\mathcal{P}_{R}(^{n})$ for $R\geqslant 1$ refers to the subset of $\mathcal{P}(^{n})$ of probability measures whose mean and covariance satisfy the bound
$\lvert\mathcal{M}(\mu)\rvert\leqslant R,\qquad\frac{1}{R^{2}}I_{n}\preccurlyeq\mathcal{C}(\mu)\preccurlyeq R^{2}I_{n}.$ (1.2)
[…]
For a probability measure $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ associated with random variable $(u,y)\in^{d_{u}}\times^{d_{y}}$ we use the notation $\mathcal{M}^{u}(\pi)$, $\mathcal{M}^{y}(\pi)$ for the means of the marginal distributions, and the notation $\mathcal{C}^{uu}(\pi)$, $\mathcal{C}^{uy}(\pi)$ and $\mathcal{C}^{yy}(\pi)$ for the blocks of the covariance matrix $\mathcal{C}(\pi)$. That is to say,
$\mathcal{M}(\pi)=\begin{pmatrix}\mathcal{M}^{u}(\pi)\\ \mathcal{M}^{y}(\pi)\end{pmatrix},\qquad\mathcal{C}(\pi)=\begin{pmatrix}\mathcal{C}^{uu}(\pi)&\mathcal{C}^{uy}(\pi)\\ \mathcal{C}^{uy}(\pi)^{\mathsf{T}}&\mathcal{C}^{yy}(\pi)\end{pmatrix}.$ (1.3)
For $h\colon^{d_{u}}\to^{d_{y}}$ we also define ${\mathcal{C}}^{hh}(\pi)$ to be the covariance of vector $h(u)$, for $u$ distributed according to the marginal of $\pi$ on $u$, and ${\mathcal{C}}^{uh}(\pi)$ to be the covariance between $u$ and $h(u)$.
We also introduce the Gaussian projection operator $\mathsf{G}\colon\mathcal{P}(^{n})\to\mathcal{G}(^{n})$ given by $\mathsf{G}\mu=\mathcal{N}\bigl{(}\mathcal{M}(\mu),\mathcal{C}(\mu)\bigr{)}.$ We refer to $\mathsf{G}$ as a projection because $\mathsf{G}\mu$ is the Gaussian distribution closest to $\mu$ with respect to $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ [MR2247587], where $\mathrm{KL}(\mu\|\nu)$ is the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$. Note that $\mathsf{G}$ defines a nonlinear mapping. Throughout this work, we employ the following weighted total variation distance, shown to be useful in the analysis of ensemble Kalman methods in [carrillo2022ensemble]:
Definition 1.1.
Let $g\colon^{n}\to[1,\infty)$ denote $g(v):=1+\lvert v\rvert^{2}$. We define the weighted total variation metric $d_{g}\colon\mathcal{P}(^{n})\times\mathcal{P}(^{n})\to$
$d_{g}(\mu_{1},\mu_{2})=\sup_{\lvert f\rvert\leqslant g}\bigl{\lvert}\mu_{1}[f]-\mu_{2}[f]\bigr{\rvert},$
where the supremum is over all functions $f\colon^{n}\to$ which are bounded from above by $g$ pointwise and in absolute value.
Remark 1.2.
The following remarks concern the distance $d_{g}$:
  • If $\mu_{1},\mu_{2}$ have Lebesgue densities $\rho_{1}$, $\rho_{2}$, then
    $d_{g}(\mu_{1},\mu_{2})=\int g(v)\left\lvert\rho_{1}(v)-\rho_{2}(v)\right\rvert\,\mathrm{d}v.$
  • Unlike the usual total variation distance, the weighted metric in Definition 1.1 enables control of the differences $\lvert\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\rvert$ and $\lVert\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\rVert$. This is the content of Lemma B.6, proved in the appendix, which is used in the proof of a key auxiliary result (Lemma B.2).
1.3 Filtering Distribution
We consider the following stochastic dynamics and data model
$u_{j+1}$ $=\Psi(u_{j})+\xi_{j},\qquad$ $\xi_{j}\sim\mathcal{N}(0,\Sigma),$ (1.4a)
$y_{j+1}$ $=h(u_{j+1})+\eta_{j+1},\qquad$ $\eta_{j+1}\sim\mathcal{N}(0,\Gamma).$ (1.4b)
Here $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}$ is the unknown state, evolving in ${}^{d_{u}}$, and $\{y_{j}\}_{j\in\llbracket 1,J\rrbracket}$ are the observations, evolving in ${}^{d_{y}}$. We assume that the initial state is distributed according to a Gaussian distribution $u_{0}\sim\mathcal{N}(m_{0},C_{0})$ and that the following independence assumption is satisfied:
$u_{0}\perp\!\!\!\!\perp\xi_{0}\perp\!\!\!\!\perp\dotsb\perp\!\!\!\!\perp\xi_{J-1}\perp\!\!\!\!\perp\eta_{1}\perp\!\!\!\!\perp\dotsb\perp\!\!\!\!\perp\eta_{J}.$ (1.5)
The filtering distribution $\mu_{j}$ is the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}:=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ of the data process up to step $j$. Data $Y_{j}^{\dagger}$ may be thought of as arising from a realization of (1.4); but the case of model misspecification, where $Y_{j}^{\dagger}$ does not necessarily arise from (1.4), is also of interest.
It is well-known [MR3363508, reich2015probabilistic] that the true filtering distribution evolves according to
$\mu_{j+1}=\mathsf{L}_{j}\mathsf{P}\mu_{j}.$ (1.6)
where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps on probability measures, referred to respectively as the prediction and analysis steps in the data assimilation community [asch2016data]. The prediction operator $\mathsf{P}$ is linear and defined by the Markov kernel associated with the stochastic dynamics (1.4a). Its action on a probability measure with Lebesgue density $\mu$ reads
$\mathsf{P}\mu(u)=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\mu(v)\,\mathrm{d}v.$ (1.7)
The operator $\mathsf{L}_{j}$ is a nonlinear map which formalizes the incorporation of the new datum $y^{\dagger}_{j+1}$ using Bayes’ theorem. Its action reads
$\mathsf{L}_{j}\mu(u)=\frac{\displaystyle\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u)}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(U)\,\mathrm{d}U}$ (1.8)
The operator $\mathsf{L}_{j}$ effects a reweighting of the measure to which it applies, with more weight assigned to the state values that are consistent with the observation. It is convenient in this work to decompose the analysis map $\mathsf{L}_{j}$ into the composition $\mathsf{B}_{j}\mathsf{Q}$, where the operators $\mathsf{Q}\colon\mathcal{P}(^{d_{u}})\to\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and $\mathsf{B}_{j}\colon\mathcal{P}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ are defined by
$\mathsf{Q}\mu(u,y)=\frac{1}{\sqrt{(2\pi)^{d_{y}}\det\Gamma}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u).$ (1.9)
and
$\mathsf{B}_{j}\mu(u)=\frac{\mu(u,y_{j+1}^{\dagger})}{\displaystyle\int_{{}^{d_{u}}}\mu(U,y_{j+1}^{\dagger})\,\mathrm{d}U}.$ (1.10)
The operator $\mathsf{Q}$ maps a probability measure with density $\mu$ into the density associated with the joint random variable $\bigl{(}U,h(U)+\eta\bigr{)}$, where $U\sim\mu$ is independent of $\eta\sim\mathcal{N}(0,\Gamma)$. The operator $\mathsf{B}_{j}$ performs conditioning on the data $y^{\dagger}_{j+1}$. Map $\mathsf{Q}$ is linear whilst $\mathsf{B}_{j}$ is nonlinear. We may thus write (1.6) in the form
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}.$ (1.11)
2 The Ensemble Kalman Filter
In Subsection 2.1 we define the specific version of the mean field ensemble Kalman filter that we analyze here; other versions may be found in [calvello2025ensemble] and will be amenable to similar analyses. In Subsection 2.2 we prove our main stability theorem, showing that the error between the true filter and its Gaussian projection, on the joint space of state and observations, may be used to control the error between the true filter and the mean field ensemble Kalman filter. In Subsection 2.3 we prove a corollary to this theorem, establishing that the mean field ensemble Kalman filter accurately approximates the true filter for a specific class of non-Gaussian problems. In Subsection 2.4 we deploy the results of the two preceding subsections to state a theorem quantifying the error between the finite particle ensemble Kalman filter itself and the true filtering distribution. All of the theorems allow for linear growth of the vector fields defining the dynamics and the observation processes.
2.1 The Algorithm
The ensemble Kalman filter may be derived as a particle approximation of various mean field dynamics [calvello2025ensemble]. The specific mean field ensemble Kalman filter that we study in this paper reads
$\widehat{u}_{j+1}$ $=\Psi(u_{j})+\xi_{j},\qquad$ $\xi_{j}\sim\mathcal{N}(0,\Sigma),$ (2.1a)
$\widehat{y}_{j+1}$ $=h(\widehat{u}_{j+1})+\eta_{j+1},\qquad$ $\eta_{j+1}\sim\mathcal{N}(0,\Gamma),$ (2.1b)
$u_{j+1}$ $=\widehat{u}_{j+1}+\mathcal{C}^{uy}\left(\widehat{\pi}_{j+1}^{\rm EK}\right)\mathcal{C}^{yy}\left(\widehat{\pi}_{j+1}^{\rm EK}\right)^{-1}\bigl{(}y^{\dagger}_{j+1}-\widehat{y}_{j+1}\bigr{)},$ (2.1c)
where $\widehat{\pi}^{\rm EK}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and independence of the noise terms (1.5) is still assumed to hold, as in (1.5). See Subsection 1.2 for the definition of the covariance matrices that appear in (2.1c). We denote by $\mu^{\rm EK}_{j}$ the law of $u_{j}$. In order to rewrite the evolution of $\mu_{j}^{\rm EK}$ in terms of maps on probability measures we introduce, for a given $y^{\dagger}_{j+1}$, the map $\mathsf{T}_{j}\colon\mathcal{P}^{2}_{\succ 0}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ defined by
$\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi.$ (2.2)
Here the subscript denotes the pushforward and, for given $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$, $z\in^{d_{y}}$,
$\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,z)\colon$${}^{d_{u}}\times^{d_{y}}\to^{d_{u}};$
$(u,y)\mapsto$ $\,\,u+\mathcal{C}^{uy}(\pi)\mathcal{C}^{yy}(\pi)^{-1}\bigl{(}z-y\bigr{)}.$ (2.3)
Evolution of the probability measure $\mu_{j}^{\rm EK}$ may then be written as [calvello2025ensemble]
$\mu_{j+1}^{\rm EK}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}.$ (2.4)
We now discuss the preceding map in relation to (1.11). The specific affine map $\mathscr{T}$ used in (2.1c) is determined by measure $\widehat{\pi}^{\rm EK}_{j+1}$ (here equal to $\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$) and data $y^{\dagger}_{j+1}$. The following observations explain why the law of $u_{j}$ in (2.1) evolves according to (2.4):
  • if $u_{j}\sim\mu_{j}^{\rm EK}$, then $\widehat{u}_{j+1}\sim\mathsf{P}\mu_{j}^{\rm EK}$, by definition of $\mathsf{P}$;
  • random vector $(\widehat{u}_{j+1},\widehat{y}_{j+1})$ is distributed according to $\widehat{\pi}^{\rm EK}_{j+1}=\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$;
  • equation (2.1c) then implies that $u_{j+1}\sim\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$.
As we show in Lemma B.7, operator $\mathsf{T}_{j}$ coincides with conditioning operator $\mathsf{B}_{j}$ over the Gaussians $\mathcal{G}(^{d_{u}}\times^{d_{y}})\subset\mathcal{P}(^{d_{u}}\times^{d_{y}}).$ In the particular case where $\mu_{0}$ is Gaussian, which is a standing assumption in this paper, and the operators $\Psi$ and $h$ are linear, the mean field ensemble Kalman filter (2.4) thus reproduces the exact filtering distribution (1.11), given by the Kalman filter itself. The aim of this paper is to analyze the accuracy of ensemble Kalman methods when $\Psi$ and $h$ are not assumed to be linear.
The ensemble Kalman filter as implemented in practice may be derived as a particle approximation of the mean field dynamics as defined by sample paths in (2.1) or as an evolution on measures in (2.4). For any $\pi$ in the range of $\mathsf{Q}$, $\mathcal{C}^{yy}(\pi)=\mathcal{C}^{hh}(\pi)+\Gamma$ and $\mathcal{C}^{uy}(\pi)=\mathcal{C}^{uh}(\pi)$, using the notation introduced in, and following, (1.3). This is since the noise in the observation component defining $\pi$ is then independent of the state component defining $\pi.$ Using this the particle approximation of (2.1) takes the form
$\widehat{u}^{(i)}_{j+1}$ $=\Psi\bigl{(}u^{(i)}_{j}\bigr{)}+\xi^{(i)}_{j},\qquad$ $\xi^{(i)}_{j}\sim\mathcal{N}(0,\Sigma),$ (2.5a)
$\widehat{y}^{(i)}_{j+1}$ $=h\bigl{(}\widehat{u}^{(i)}_{j+1}\bigr{)}+\eta^{(i)}_{j+1},\qquad$ $\eta^{(i)}_{j+1}\sim\mathcal{N}(0,\Gamma),$ (2.5b)
$u^{(i)}_{j+1}$ $=\widehat{u}^{(i)}_{j+1}+\mathcal{C}^{uh}\left(\widehat{\pi}_{j+1}^{{\rm EK},N}\right)\Bigl{(}\mathcal{C}^{hh}\Bigl{(}\widehat{\pi}_{j+1}^{{\rm EK},N}\Bigr{)}+\Gamma\Bigr{)}^{-1}\bigl{(}y^{\dagger}_{j+1}-\widehat{y}^{(i)}_{j+1}\bigr{)}.$ (2.5c)
where $\widehat{\pi}^{{\rm EK},N}_{j+1}$ is the empirical measure
$\widehat{\pi}^{{\rm EK},N}_{j+1}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\bigl{(}\widehat{u}^{(i)}_{j+1},\widehat{y}^{(i)}_{j+1}\bigr{)}},$
and $\xi^{(i)}_{j}\sim\mathcal{N}(0,\Sigma)$ i.i.d. in both $i$ and $j$ and $\eta^{(i)}_{j+1}\sim\mathcal{N}(0,\Gamma)$ i.i.d. in both $i$ and $j$; furthermore, the set of $\{\xi^{(i)}_{j}\}$ are independent from the set of $\{\eta^{(i)}_{j+1}\}$. Choosing to express the particle approximation of the covariance in observation space through $\mathcal{C}^{hh}$ and $\Gamma$ ensures invertibility, provided $\Gamma$ is invertible. From the particles evolving according to the dynamics in (2.5) we define the empirical measure
$\mu_{j+1}^{{\rm EK},N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{u_{j+1}^{(i)}},$ (2.6)
whose evolution describes the ensemble Kalman filter.
Our theorems in Subsections 2.2 and 2.3 concern relationships between the true filter (1.11) and the mean field ensemble Kalman filter (2.4). We study the setting in which $\Psi$ and $h$ are not assumed to be linear, so that the true filter is not Gaussian, in Subsection 2.2; and then we study small perturbations away from the Gaussian setting that arise when the vector fields $\Psi$ and $h$ are close to affine in Subsection 2.3. The theorem in Subsection 2.4 concerns the relationship between the true filter (1.11) and the ensemble Kalman filter (2.6). In this subsection, we combine existing analysis on the convergence of the ensemble Kalman filter to the mean field ensemble Kalman filter with the results from the previous subsections to state and prove an error estimate between the ensemble Kalman filter and the true filter in the non-linear setting. Specifically, we study the case of a vector field $\Psi$ that is a bounded perturbation away from affine and an affine vector field $h$. To state our theorems we will use the following set of assumptions:
Assumption H.
There exists positive constants $\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\sigma$ and $\gamma$ such that the data $\{y_{j}^{\dagger}\}$, the vector fields $(\Psi,h)$ and the covariances $(\Sigma,\Gamma)$ satisfy:
  1. (H1)
    data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ lies in set $B_{y}\subset^{KJ}$ defined by
    $B_{y}:=\left\{Y^{\dagger}\in^{KJ}:\max_{j\in\llbracket 1,J\rrbracket}\lvert y^{\dagger}_{j}\rvert\leqslant\kappa_{y}\right\};$
  2. (H2)
    function $\Psi\colon^{d_{u}}\to^{d_{u}}$ satisfies $\bigl{\lvert}\Psi(u)\bigr{\rvert}\leqslant\kappa_{\Psi}\bigl{(}1+\lvert u\rvert\bigr{)}$ for all $u\in^{d_{u}}$;
  3. (H3)
    function $h\colon^{d_{u}}\to^{d_{y}}$ satisfies $\bigl{\lvert}h(u)\bigr{\rvert}\leqslant\kappa_{h}\bigl{(}1+\lvert u\rvert\bigr{)}$ for all $u\in^{d_{u}}$;
  4. (H4)
    function $h\colon^{d_{u}}\to^{d_{y}}$ satisfies $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$;
  5. (H5)
    covariance matrices $\Sigma$ and $\Gamma$ are positive definite: $\Sigma\succcurlyeq\sigma I_{d_{u}}$ and $\Gamma\succcurlyeq\gamma I_{d_{y}}$ for positive $\sigma$ and $\gamma$.
Assumption V.
The vector fields $(\Psi,h)$ are affine, i.e. they satisfy the following
  1. (V1)
    The function $\Psi\colon^{d_{u}}\to^{d_{u}}$ satisfies $\Psi(u)\coloneqq Mu+b$, for some $M\in^{d_{u}\times d_{u}}$ and $b\in^{d_{u}}$.
  2. (V2)
    The function $h\colon^{d_{u}}\to^{d_{y}}$ satisfies $h(u)\coloneqq Hu+w$, for some $H\in^{d_{y}\times d_{u}}$ and $w\in^{d_{y}}$.
2.2 Stability Theorem: Mean Field Ensemble Kalman Filter
Roughly speaking, our stability theorem states that if the true filtering distributions $(\mu_{j})_{j\in\llbracket 0,J-1\rrbracket}$ are close to Gaussian, after appropriate lifting to the state/data space, then the distribution $\mu^{\rm EK}_{j}$ given by the mean field ensemble Kalman filter (2.4) is close to the true filtering distribution $\mu_{j}$ given by (1.11) for all $j\in\llbracket 0,J\rrbracket$.
Theorem 2.1 (Stability: Mean Field Ensemble Kalman Filter).
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained respectively from the dynamical systems (1.11) and (2.4), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$. That is,
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\qquad\mu^{\rm EK}_{j+1}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}.$
If Assumption H holds, then there exists $C=C\bigl{(}\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma,J\bigr{)}$ such that
$d_{g}(\mu^{\rm EK}_{J},\mu_{J})\leqslant C\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
Proof of Theorem 2.1 below relies on the following auxiliary results, all proved in Appendix B.
  1. 1.
    For any probability measure $\mu$ with finite first and second order polynomial moments $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$, the means of the probability measures $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their covariances are bounded both from above and from below. The constants in these bounds depend only on the parameters $\kappa_{\Psi}$, $\kappa_{h}$, $\Sigma,\Gamma$ and on $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$. See Lemmas B.1 and B.2.
  2. 2.
    Let $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 1,J\rrbracket}$ denote the probability measures obtained respectively from the dynamical systems (1.11) and (2.4), initialized at the same Gaussian measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$. Then for any $q\in\mathbb{N}$, there exist constants $M_{q},M_{q}^{\rm EK}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
    $\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}(\mu_{j})\leqslant M_{q}\quad\text{and}\quad\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}\bigl{(}\mu^{\rm EK}_{j}\bigr{)}\leqslant M_{q}^{\rm EK}.$
    See Lemmas B.3 and B.5. This will facilitate use of the stability results from items 6 and 7.
  3. 3.
    For any Gaussian measure $\mu\in\mathcal{G}(^{d_{u}}\times^{d_{y}})$, it holds that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma B.7 and also [calvello2025ensemble].
  4. 4.
    The map $\mathsf{P}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{P}}$ depending on the parameters $\kappa_{\Psi}$ and $\Sigma$. See Lemma B.8.
  5. 5.
    The map $\mathsf{Q}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{Q}}$ depending on the parameters $\kappa_{h}$ and $\Gamma$. See Lemma B.9.
  6. 6.
    The map $\mathsf{B}_{j}$ satisfies, for any $\pi\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, the bound
    $\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{B}_{j}\mathsf{G}\pi,\mathsf{B}_{j}\pi)\leqslant C_{\mathsf{B}}d_{g}(\mathsf{G}\pi,\pi),$
    where $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$. This statement concerns the stability of the $\mathsf{B}_{j}$ operator between a measure and its Gaussian approximation. See Lemma B.10.
  7. 7.
    The map $\mathsf{T}_{j}$ satisfies the following bound: for all $R\geqslant 1$, it holds for all probability measures $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$ that
    $\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p),$
    for a constant $L_{\mathsf{T}}=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$. This statement may be viewed as a local Lipschitz continuity result in the case where the second argument of $d_{g}$ is restricted to the range of $\mathsf{Q}\mathsf{P}$. See Lemma B.11.
Proof of Theorem 2.1.
In what follows we refer to the preceding itemized list to clarify the proof. For notational simplicity it is helpful to define the following measure of the difference between the true filtering distribution and its Gaussian projection:
$\varepsilon:=\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$ (2.7)
Assume throughout the following that $j\in\llbracket 0,J-1\rrbracket.$ The proof rests on the following use of the triangle inequality:
$d_{g}(\mu^{\rm EK}_{j+1},\mu_{j+1})$ $=d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$
$\leqslant d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)+d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)+d_{g}\left(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right).$ (2.8a)
We have used the fact that $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by Item 3 (Lemma B.7). Item 2 (Lemmas B.3 and B.5) shows that, for any $q\in\mathbb{N}$, there exist constants $M_{q},M_{q}^{\rm EK}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
$\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}(\mu_{j})\leqslant M_{q}\quad\text{and}\quad\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}\bigl{(}\mu^{\rm EK}_{j}\bigr{)}\leqslant M_{q}^{\rm EK}.$
Therefore by Item 1 (Lemma B.2), there is a constant $R\geqslant 1$, depending on the covariance matrices $\Sigma$, $\Gamma$, the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, and the moment bounds $M_{2}$ and $M_{2}^{\rm EK}$, such that for any $j\in\llbracket 0,J-1\rrbracket$ it holds that $\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$. By Items 4, 5 and 7 (Lemmas B.11, B.9 and B.8), the first term in (2.8a) satisfies
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$ $\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}d_{g}\left(\mu_{j}^{\rm EK},\mu_{j}\right),$ (2.9)
where, for conciseness, we omitted the dependence of the constants on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
Since it holds that $\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\in\mathcal{G}_{R}(^{d_{u}}\times^{d_{y}})$ by definition of $R$, the second term in (2.8a) may be bounded using Item 7 (Lemma B.11) and the definition in (2.7) of $\varepsilon$:
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}d_{g}\left(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}\varepsilon.$
Finally, the third term in (2.8a) can be bounded using item 6 (Lemma B.10) and the definition in (2.7) of $\varepsilon$:
$d_{g}\left(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant C_{\mathsf{B}}d_{g}\left(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant C_{\mathsf{B}}\varepsilon.$
Therefore, letting $\ell=L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}$, we have shown that
$d_{g}(\mu^{\rm EK}_{j+1},\mu_{j+1})$ $\leqslant\ell d_{g}(\mu^{\rm EK}_{j},\mu_{j})+\bigl{(}L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}+C_{\mathsf{B}}\bigr{)}\varepsilon$
and the conclusion follows from the discrete Grönwall lemma, using the fact that $\mu_{0}^{\rm EK}=\mu_{0}.$
2.3 Error Estimate: Mean Field Ensemble Kalman Filter
Theorem 2.1 shows that the mean field ensemble Kalman filter error can be made arbitrarily small if the true filter is arbitrarily close to its Gaussian projection, in state-observation space. This “closeness to Gaussian” assumption can be satisfied in our setting of unbounded vector fields by considering small perturbations of affine vector fields, stated and proved in Proposition 2.2. Combining Proposition 2.2 with Theorem 2.1 gives an error estimate for the mean field ensemble Kalman filter, yielding Theorem 2.3.
Proposition 2.2 (Approximation Result).
Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). Fix $\kappa_{\Psi},\kappa_{h}>0$ and assume that $\Psi_{0}\colon^{d_{u}}\to^{d_{u}}$ and $h_{0}\colon^{d_{u}}\to^{d_{y}}$ are affine functions and hence satisfy Assumptions (V1) and (V2), respectively, while also satisfying Assumptions (H2) and (H3) with $\kappa_{\Psi},\kappa_{h}$. Let $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ denote the true filtering distribution associated with functions $(\Psi,h)$, initialized at the Gaussian probability measure $\mu_{0}=\mathcal{N}(m_{0},C_{0})\in\mathcal{G}(^{d_{u}})$. Then, for any $J\in\mathbb{Z}^{+}$, there is $C=C(m_{0},C_{0},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J)>0$ such that for all $\varepsilon\in[0,1]$ and all $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$, it holds that
$\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j})\leqslant C\varepsilon.$ (2.10)
Proof.
In what follows $(\mu^{0}_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ denote the true filtering distributions associated with functions $(\Psi_{0},h_{0})$ and $(\Psi,h)$, respectively, initialized at the same Gaussian measure $\mathcal{N}(m_{0},C_{0})$. Furthermore, we let $\mathsf{P}_{0}$ and $\mathsf{Q}_{0}$ denote the kernel integral operators (1.7) and (1.9) defined by the specific vector fields $(\Psi_{0},h_{0}).$ By Lemma B.3, the filtering distributions have bounded second moments. Let
[…]
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context, depends only on the constants $m_{0},C_{0},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,j$ (in particular, it does not depend on $\varepsilon$, $\Psi$, $h$), and may change from line to line. Fix $j\in\llbracket 0,J-1\rrbracket$. Note that the filtering distribution defined by $(\Psi_{0},h_{0})$ is Gaussian. Then, using the triangle inequality and Gaussianity of $\mathsf{Q}_{0}\mathsf{P}_{0}\mu_{j}^{0}$ we obtain
$d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j})$ $\leqslant d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{j})+d_{g}(\mathsf{G}\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{j},\mathsf{Q}\mathsf{P}\mu_{j})$
$=d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{j})+d_{g}(\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{j},\mathsf{Q}\mathsf{P}\mu_{j}).$
We note that since the filters have bounded first and second order polynomial moments, by Lemmas B.1 and B.2 we may deduce that there exists $R\geqslant 1$ such that
$\forall j\in\llbracket 0,J-1\rrbracket,\qquad\mathsf{P}_{0}\mu^{0}_{j},\in\mathcal{P}_{R}(^{d_{u}}),\qquad\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{j}\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}}).$ (2.11)
The second part of this display allows direct application of Lemma C.1, which concerns the local Lipschitz continuity result for the Gaussian projection operator $\mathsf{G}$, to obtain
$d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j})$ $\leqslant Cd_{g}(\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{j},\mathsf{Q}\mathsf{P}\mu_{j})$
$\leqslant Cd_{g}(\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{j},\mathsf{Q}\mathsf{P}_{0}\mu^{0}_{j})+Cd_{g}(\mathsf{Q}\mathsf{P}_{0}\mu^{0}_{j},\mathsf{Q}\mathsf{P}\mu_{j}).$
Using (C.2) from Lemma C.2, noting that since $\mathsf{P}_{0}\mu^{0}\in\mathcal{P}_{R}(^{d_{u}})$ by (2.11) it holds that $\mathcal{M}_{2}(\mathsf{P}_{0}\mu^{0}_{j})\leqslant d_{u}R$, and using the Lipschitz continuity of $\mathsf{Q}$ (Lemma B.9) we deduce that
$d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j})$ $\leqslant C\varepsilon(1+d_{u}R)+Cd_{g}(\mathsf{P}_{0}\mu^{0}_{j},\mathsf{P}\mu_{j})$
$\leqslant C\varepsilon(1+d_{u}R)+Cd_{g}\bigl{(}\mathsf{P}_{0}\mu^{0}_{j},\mathsf{P}\mu^{0}_{j}\bigr{)}+Cd_{g}\bigl{(}\mathsf{P}\mu^{0}_{j},\mathsf{P}\mu_{j}\bigr{)}$
$\leqslant C\varepsilon(1+d_{u}R)+C\varepsilon\mathcal{R}+Cd_{g}(\mu^{0}_{j},\mu_{j}),$
where the second inequality follows by the triangle inequality and the third inequality follows from bounding $d_{g}(\mathsf{P}_{0}\mu^{0}_{j},\mathsf{P}\mu_{j})$ using (C.1) from Lemma C.2 and from the Lipschitz continuity of $\mathsf{P}$ (Lemma B.9). The statement then follows from Lemma C.3. ∎
Theorem 2.3 (Error Estimate: Mean Field Ensemble Kalman Filter).
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained respectively from the dynamical systems (1.11) and (2.4) initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$. That is,
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\qquad\mu^{\rm EK}_{j+1}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}.$
Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). Fix $\kappa_{\Psi},\kappa_{h},\ell_{h}>0$ and let the vector fields $\Psi_{0},h_{0}$ be affine functions satisfying Assumptions (V1) and (V2), respectively, while also satisfying Assumptions (H2) and (H3) with $\kappa_{\Psi},\kappa_{h}$. Then for all $\varepsilon\in[0,1]$, there exists a constant  $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{\max\{3+d_{u},4+d_{y}\}}$, such that for any $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$ with $h$ satisfying Assumption (H4), it holds that
$d_{g}(\mu^{\rm EK}_{J},\mu_{J})\leqslant C\varepsilon.$
Proof.
Since Assumption H is satisfied, it is possible to apply the result of Theorem 2.1 to deduce that there exists $C=C\bigl{(}\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$ such that
$d_{g}(\mu^{\rm EK}_{J},\mu_{J})\leqslant C\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
Additionally, since $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$ for $\Psi_{0},h_{0}$ satisfying Assumptions (V1) and (V2), and moreover  Assumptions (H2) and (H3), we may apply Proposition 2.2 to deduce the result. ∎
2.4 Error Estimate: Finite Particle Ensemble Kalman Filter
In this subsection, we combine the results from the work in [le2009large] with stability Theorem 2.1, together with approximation Theorem 2.3, to derive a quantitative error estimate between the finite particle ensemble Kalman filter and the true filter in the non-linear setting. In order to define an appropriate metric we introduce the following class of vector fields.
Assumption P1.
The vector field $\phi:^{d_{u}}\to$ satisfies for any $u,v\in^{d_{u}}$ the condition
$\left|\phi(u)-\phi(v)\right|\leqslant L_{\phi}|u-v|\bigl{(}1+|u|^{{\varsigma}}+|v|^{{\varsigma}}\bigr{)},$
for some ${\varsigma}\geqslant 0$ and for some $L_{\phi}>0$. We note that for any such $\phi$, there exists $R_{\phi}>0$ which depends on $L_{\phi}$ so that $|\phi(u)|\leqslant R_{\phi}\bigl{(}1+|u|^{{\varsigma}+1}\bigr{)}$ for any $u\in^{d_{u}}$.
We will prove various technical lemmas under the Assumption P1; these may be useful beyond the confines of this paper. However for our theorems we use the more specific Assumption P2 which enables the control of first and second moments.
Assumption P2.
The vector field $\phi:^{d_{u}}\to$ satisfies for any $u,v\in^{d_{u}}$ the condition
$\left|\phi(u)-\phi(v)\right|\leqslant L_{\phi}|u-v|\bigl{(}1+|u|+|v|\bigr{)},$
for some $L_{\phi}>0$. Note that for any such $\phi$, there exists $R_{\phi}>0$ depending on $L_{\phi}$ so that $|\phi(u)|\leqslant R_{\phi}\bigl{(}1+|u|^{2}\bigr{)}$ for any $u\in^{d_{u}}$.
Theorem 2.4 (Error Estimate: Finite Particle Ensemble Kalman Filter).
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{{\rm EK},N}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained respectively from the dynamical systems (1.11) and (2.6), initialized at the Gaussian probability measure $\mu_{0}\in~{}\mathcal{G}(^{d_{u}})$ and at the empirical measure $\mu_{0}^{{\rm EK},N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{u_{0}^{(i)}}$ for $u_{0}^{(i)}\sim\mu_{0}$ i.i.d. samples, respectively. That is,
$\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\qquad\mu_{j+1}^{{\rm EK},N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{u_{j+1}^{(i)}},$
where $u_{j+1}^{(i)}$ evolve according to the iteration in (2.5). Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). Assume that the vector field $h$ is linear, and let $\kappa_{h},\ell_{h}$ be positive constants such that Assumptions (H3) and (H4) are satisfied. Furthermore, let the vector field $\Psi_{0}$ be an affine function satisfying Assumption (V1) as well as Assumption (H2) with $\kappa_{\Psi}>0$. Additionally, let $\ell_{\Psi}>0$ be a constant. Then, for all $\varepsilon\in[0,1]$, it holds that for any $\Psi\in B_{L^{\infty}}\bigl{(}\Psi_{0},\varepsilon\bigr{)}$ satisfying $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\ell_{\Psi},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{\max\{3+d_{u},4+d_{y},16\cdot 3^{J-1}\}}$, such that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant C\Bigl{(}\frac{1}{\sqrt{N}}+{\varepsilon}\Bigr{)},$
for any $\phi$ satisfying Assumption P2.
The proof presented hereafter relies on the following elements.
  1. 1.
    We apply the triangle inequality in order to employ two distinct results concerning the mean field ensemble Kalman filter. The proof thus involves quantifying the error between finite particle ensemble Kalman filter and the mean field ensemble Kalman filter, as discussed in Item 2, and the error between the mean field ensemble Kalman filter and the true filter, as discussed in Item 3.
  2. 2.
    The work in [le2009large] establishes a Monte Carlo error estimate, with rate of $1/\sqrt{N}$, between the empirical measure $\mu^{{\rm EK},N}$, representing the particle ensemble Kalman filter, and the measure $\mu^{\rm EK}$ describing the evolution of the mean field ensemble Kalman filter. In particular, this holds under Assumptions (H1) and (H5) on the data and covariances of the noise processes, the linearity of the vector field $h$ and Assumption (H2) on $\Psi$ with the additional assumption that $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$. In Lemma D.1 we mimic the proof of [le2009large, Proposition 4.4] and  [le2009large, Theorem 5.2] to gain insight into the dependence of the constant prefactor multiplying $1/\sqrt{N}$ on the parameters of the Gaussian initial condition and the number of steps $J$.
  3. 3.
    We assume that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5), and assume that $h$ satisfies Assumption (V2). Furthermore, we assume $\Psi$ satisfies Assumption (H2) and $\Psi\in B_{L^{\infty}}(\Psi_{0},\varepsilon)$ with $\Psi_{0}$ satisfying Assumption (V1). These assumptions allow us to apply the the result from Theorem 2.3.
Proof.
Recall that $\mu_{j}^{\rm EK}$ is the mean field ensemble Kalman filter, here initialized at the same Gaussian $\mu_{0}$ as the true filter. We fix a function $\phi$ satisfying Assumption P2 and apply the triangle inequality to deduce that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}+\left(\mathbb{E}~{}\Bigl{|}\mu^{\rm EK}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}.$ (2.12)
Since $\mu^{\rm EK}_{J}$ and $\mu_{J}$ are deterministic probability measures, it holds that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{\rm EK}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}=\Bigl{|}\mu^{\rm EK}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}.$
Since $\phi$ satisfies Assumption P2, it follows that
$\Bigl{|}\mu^{\rm EK}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}\leqslant\sup_{|\phi|\leqslant R_{\phi}(1+|u|^{2})}\Bigl{|}\mu^{\rm EK}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}=R_{\phi}\cdot d_{g}\bigl{(}\mu^{\rm EK}_{J},\mu_{J}\bigr{)}.$
We note that Assumption H holds as $h$ is assumed to be affine; since we additionally assume that $\Psi\in B_{L^{\infty}}\bigl{(}\Psi_{0},\varepsilon\bigr{)}$, where $\Psi_{0}$ is an affine function, we may apply the result of Theorem 2.3 to deduce that
$d_{g}\bigl{(}\mu^{\rm EK}_{J},\mu_{J}\bigr{)}\leqslant C{\varepsilon},$ (2.13)
where $C$ is a constant depending on $\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma$.
The fact that $h$ is assumed to be linear and the additional assumption that $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$ allows us to apply the result of Lemma D.1, so that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}},$ (2.14)
where $C$ is a constant depending on $\mathcal{M}_{16\cdot 3^{J-1}}(\mu_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$. In view of (2.12), combining (2.13) and (2.14) yields the desired result. ∎
3 Discussion and Future Directions
In this paper we have presented the first analysis, in the setting of a filtering problem defined by nonlinear state space dynamics, quantifying the error between the empirical measure obtained by the ensemble Kalman filter and the true filtering distribution. The analysis for the EnKF outlined is based on the proof methodology developed for particle filters in [rebeschini2015can]. This work extends the initial results of [carrillo2022ensemble] to a setting of practical interest, for which dynamical models and observation operators are allowed to be unbounded and nonlinear. The work opens a new avenue for analysis in the quantitative understanding of the performance of the ensemble Kalman methodology applied to nonlinear problems. To this end, we identify the following avenues for further work.
  1. 1.
    As surveyed in [calvello2025ensemble], the ensemble Kalman filtering methodology may be used for solving inverse problems and for sampling; it is of interest to extend this analysis to the ensemble Kalman based inversion algorithms outlined in that paper.
  2. 2.
    There is a substantial body of literature studying the continuous time limits of ensemble Kalman methods; see [calvello2025ensemble] for a review. Performing analysis analogous to that presented here, but in the continuous time setting, would be of interest.
  3. 3.
    For large scale applications, there has been recent wide interest in replacing the dynamical model $\Psi$, representing the solution operator obtained via a high fidelity numerical solver, with a cheap to evaluate surrogate. Multifidelity ensemble methods [bach2023multi] allow the use of a small number of particles evolved according to the high fidelity solver and a large number evolved according to the surrogate. Extending our error analysis to incorporate the effect of model error would be of interest in this context.
Acknowledgments
The research of PM is supported by the projects SWIDIMS (ANR-20-CE40-0022) and CONVIVIALITY (ANR-23-CE40-0003) of the French National Research Agency. The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship, and by the SciAI Center, funded by the Office of Naval Research (ONR), Grant Number N00014-23-1-2729; these grants also support the work of EC. UV is partially supported by the European Research Council (ERC) under the EU Horizon 2020 programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grants ANR-21-CE40-0006 (SINEQ) and ANR-23-CE40-0027 (IPSO). The work was initiated during The Jacques-Louis Lions Lectures given by AMS at Sorbonne Université in December 2023; the authors are grateful to Albert Cohen for hosting this event.
Appendix A Auxiliary Results
We first state and prove a standard result that will be used throughout the paper.
Lemma A.1.
Let $X$ be a random variable taking values in ${}^{d_{u}}$ with finite second moment, and let $m:=\mathbb{E}[X]$. Then it holds that
$\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}=\mathbb{E}\bigl{[}XX^{\mathsf{T}}\bigr{]}-mm^{\mathsf{T}}$ (A.1)
and
$\forall a\in^{d_{u}},\qquad\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}\succcurlyeq\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}.$ (A.2)
Proof.
It holds that
$\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}$ $=\mathbb{E}\Bigl{[}\bigl{(}(X-m)+(m-a)\bigr{)}\bigl{(}(X-m)+(m-a)\bigr{)}^{\mathsf{T}}\Bigr{]}$
$=\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}+(m-a)(m-a)^{\mathsf{T}}.$
Setting $a=0$, we obtain (A.1). The inequality (A.2) follows by noting that the second term in the last expression is positive semidefinite. ∎
Lemma A.2.
Denote by $g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m,S)$ the Lebesgue density of $\mathcal{N}(m,S)$ and by $\mathcal{S}^{n}_{\alpha}$ the set of symmetric $n\times n$ matrices $M$ satisfying
$\frac{1}{\alpha}I_{n}\preccurlyeq M\preccurlyeq\alpha I_{n}.$ (A.3)
Then for all $n\in\mathbb{N}^{+}$ and $\alpha\geqslant 1$, there exists $L_{n,\alpha}>0$ such that for all parameters $(c_{1},m_{1},S_{1})\in\times^{n}\times\mathcal{S}^{n}_{\alpha}$ and $(c_{2},m_{2},S_{2})\in\times^{n}\times\mathcal{S}^{n}_{\alpha}$,
$\lVert h\rVert_{\infty}\leqslant L_{n,\alpha}\lVert h\rVert_{1},\qquad h(y)=c_{1}g(y;m_{1},S_{1})-c_{2}g(y;m_{2},S_{2}).$ (A.4)
Proof.
The lemma as stated may be found in [carrillo2022ensemble, Lemma 14], where a complete proof is given. ∎
Lemma A.3.
Let $\mathsf{P}$ and $\mathsf{Q}$ denote the operators on probability measures given respectively in (1.7) and (1.9). Let Assumptions (H2), (H3), (H4) and (H5) be satisfied and suppose that $\mu\in\mathcal{P}(^{d_{u}})$ satisfies, for some $q>0$, the moment bound
$\mathcal{M}_{q}(\mu)=\int_{{}^{d_{u}}}\lvert x\rvert^{q}\,\mu(\mathrm{d}x)<\infty.$ (A.5)
Then there is $L=L(\mathcal{M}_{q}(\mu),\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma)>0$ such that for all $(u_{1},u_{2},y)\in^{d_{u}}\times^{d_{u}}\times^{d_{y}}$, the probability density of $p=\mathsf{Q}\mathsf{P}\mu$ satisfies
$\bigl{\lvert}p(u_{1},y)-p(u_{2},y)\bigr{\rvert}$ $\leqslant L\lvert u_{1}-u_{2}\rvert\min\left\{\max\left\{\frac{1}{1+\lvert u_{1}\rvert^{q}},\frac{1}{1+\lvert u_{2}\rvert^{q}}\right\},\frac{1}{1+\lvert y\rvert^{q}}\right\}.$ (A.6)
Proof.
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context, depends only on $\mathcal{M}_{q}(\mu),\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma$, and may change from line to line. Sometimes we write the dependence explicitly to indicate which parameters are involved.
Step 1. Bounding the density of $\mathsf{P}\mu$
We first rewrite
$\bigl{(}1+|u|^{q}\bigr{)}\mathsf{P}\mu(u)=C(\Sigma)\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\lvert u-\Psi(v)\rvert_{\Sigma}^{2}\right)\frac{1+\lvert u\rvert^{q}}{1+\lvert v\rvert^{q}}\,\bigl{(}1+\lvert v\rvert^{q}\bigr{)}\mu(\mathrm{d}v).$
We note that
$S(\Sigma,\kappa_{\Psi}):=\sup_{(u,v)\in^{d_{u}}\times^{d_{u}}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\,\frac{1+|u|^{q}}{1+|v|^{q}}<\infty;$
this may be seen observing that
$\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\,\frac{1+|u|^{q}}{1+|v|^{q}}\leqslant\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\,\frac{(1+2^{q-1}|u-\Psi(v)|^{q})+2^{q-1}|\Psi(v)|^{q}}{1+|v|^{q}}.$ (A.7)
The first term on the righthand side of (A.7) may be bounded by noting that $1+2^{q-1}e^{-x^{2}}x^{q}$ is uniformly bounded in $x\in$; on the other hand the second term may be bounded by applying Assumption (H2). Now, using (A.5), we obtain that
$\mathsf{P}\mu(u)\leqslant\frac{C\bigl{(}\Sigma,\kappa_{\Psi},\mathcal{M}_{q}(\mu)\bigr{)}}{1+|u|^{q}}.$ (A.8)
Step 2. Establishing Lipschitz continuity of $u\mapsto\mathsf{P}\mu(u)$
Since $g(x):=\operatorname{e}^{-x^{2}}$ has derivative $2x\operatorname{e}^{-x^{2}}$ and $\lvert x\operatorname{e}^{-x^{2}}\rvert\leqslant\operatorname{e}^{-\frac{x^{2}}{2}}$ for all $x\in$, it holds for some $\xi$ between $|a|$ and $|b|$ that
$\forall(a,b)\in\times,\qquad\left\lvert\operatorname{e}^{-a^{2}}-\operatorname{e}^{-b^{2}}\right\rvert=\lvert b-a\rvert\,\lvert g^{\prime}(\xi)\rvert\leqslant 2\lvert b-a\rvert\left(\operatorname{e}^{-\frac{a^{2}}{2}}+\operatorname{e}^{-\frac{b^{2}}{2}}\right).$ (A.9)
Using this inequality with $a^{2}=\frac{1}{2}\left\lvert u_{1}-\Psi(v)\right\rvert_{\Sigma}^{2}$ and $b^{2}=\frac{1}{2}\left\lvert u_{2}-\Psi(v)\right\rvert_{\Sigma}^{2}$, the triangle inequality, and equivalence of norms, we deduce that, for all $(u_{1},u_{2},v)\in^{d_{u}}\times^{d_{u}}\times^{d_{u}}$,
$\left\lvert\exp\left(-\frac{1}{2}\left\lvert u_{1}-\Psi(v)\right\rvert_{\Sigma}^{2}\right)-\exp\left(-\frac{1}{2}\left\lvert u_{2}-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\right\rvert$
$\hskip 56.9055pt\leqslant C\lvert u_{2}-u_{1}\rvert\left(\exp\left(-\frac{1}{4}\left\lvert u_{1}-\Psi(v)\right\rvert_{\Sigma}^{2}\right)+\exp\left(-\frac{1}{4}\left\lvert u_{2}-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\right)$
for constant $C=C(\Sigma).$ Integrating out the $v$ variable with respect to $\mu$ we obtain that
$\lvert\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\rvert$ $\leqslant C\lvert u_{1}-u_{2}\rvert\int_{{}^{d_{u}}}\exp\left(-\frac{1}{4}\left\lvert u_{1}-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\,\mu(\mathrm{d}v)$
$\quad+C\lvert u_{1}-u_{2}\rvert\int_{{}^{d_{u}}}\exp\left(-\frac{1}{4}\left\lvert u_{2}-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\,\mu(\mathrm{d}v).$
The integrals on the right-hand side can be bounded as in the first step, which leads to the inequality
$\forall(u_{1},u_{2})\in^{d_{u}}\times^{d_{u}},\qquad\lvert\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\rvert\leqslant C\lvert u_{1}-u_{2}\rvert\max\left\{\frac{1}{1+\lvert u_{1}\rvert^{q}},\frac{1}{1+\lvert u_{2}\rvert^{q}}\right\}.$ (A.10)
Step 3. Obtaining a coarse estimate
In view of the elementary inequality (A.9) and the assumed Lipschitz continuity of $h$, it holds that, for constant $C=C(\Gamma,\ell_{h})$
$\Bigl{\lvert}\mathcal{N}\bigl{(}h(u_{1}),\Gamma\bigr{)}(y)-\mathcal{N}\bigl{(}h(u_{2}),\Gamma\bigr{)}(y)\Bigr{\rvert}$
$\leqslant C\lvert u_{1}-u_{2}\rvert\exp\left(-\frac{1}{4}\lvert y-h(u_{1})\rvert_{\Gamma}^{2}\right)+C\lvert u_{1}-u_{2}\rvert\exp\left(-\frac{1}{4}\lvert y-h(u_{2})\rvert_{\Gamma}^{2}\right).$ (A.11)
Using the decomposition
$p(u_{1},y)-p(u_{2},y)$ $=\mathsf{P}\mu(u_{1})\,\mathcal{N}\bigl{(}h(u_{1}),\Gamma\bigr{)}(y)-\mathsf{P}\mu(u_{2})\,\mathcal{N}\bigl{(}h(u_{2}),\Gamma\bigr{)}(y)$
$=\bigl{(}\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\bigr{)}\,\mathcal{N}\bigl{(}h(u_{1}),\Gamma\bigr{)}(y)$
$\quad+\mathsf{P}\mu(u_{2})\Bigl{(}\mathcal{N}\bigl{(}h(u_{1}),\Gamma\bigr{)}(y)-\mathcal{N}\bigl{(}h(u_{2}),\Gamma\bigr{)}(y)\Bigr{)}.$
and employing (A.8), (A.10) and (A.11), we deduce that
$\lvert p(u_{1},y)-p(u_{2},y)\rvert$ $\leqslant C\lvert u_{1}-u_{2}\rvert\max\left\{\frac{1}{1+\lvert u_{1}\rvert^{q}},\frac{1}{1+\lvert u_{2}\rvert^{q}}\right\}$
×$\qquad\times\max\left\{\exp\left(-\frac{1}{4}\lvert y-h(u_{1})\rvert_{\Gamma}^{2}\right),\exp\left(-\frac{1}{4}\lvert y-h(u_{2})\rvert_{\Gamma}^{2}\right)\right\}.$ (A.12)
Note that, for the function $h(u)=u$, the quantity multiplying $\lvert u_{1}-u_{2}\rvert$ on the right-hand does not tend to 0 along the sequence $\left(u_{1}^{(n)},u_{2}^{(n)},y^{(n)}\right)=(0,n,n)$. For our purposes in this work, we need the finer estimate (A.6); establishing this bound is the aim of the next two steps.
Step 4. Bounding the density $p(u,y)$
Recall that $p(u,y)=\mathsf{P}\mu(u)\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(y)$. We prove in this step the inequality
$p(u,y)\leqslant C\min\left\{\frac{1}{1+\lvert u\rvert^{q}},\frac{1}{1+\lvert y\rvert^{q}}\right\},$ (A.13)
or equivalently
$\sup_{(u,y)\in^{d_{u}}\times^{d_{y}}}p(u,y)\,\max\left\{1+\lvert u\rvert^{q},1+\lvert y\rvert^{q}\right\}<\infty.$
To this end, note that
$p(u,y)\max\left\{\lvert u\rvert^{q},\lvert y\rvert^{q}\right\}$ $=\mathsf{P}\mu(u)\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(y)\max\left\{\lvert u\rvert^{q},\lvert y\rvert^{q}\right\}$
$\leqslant\mathsf{P}\mu(u)\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(y)\lvert u\rvert^{q}+\mathsf{P}\mu(u)\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(y)\lvert y\rvert^{q}.$
The first term is bounded uniformly by Step 1, estimate (A.8). For the second term, we use that
$\lvert y\rvert^{q}\leqslant 2^{q-1}\bigl{\lvert}h(u)\bigr{\rvert}^{q}+2^{q-1}\bigl{\lvert}y-h(u)\bigr{\rvert}^{q},$ (A.14)
to obtain
$\mathsf{P}\mu(u)\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(y)\lvert y\rvert^{q}\leqslant C\mathsf{P}\mu(u)\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(y)\Bigl{(}\lvert h(u)\rvert^{q}+\lvert y-h(u)\rvert^{q}\Bigr{)}.$
The first term on the right-hand side is bounded uniformly, again by Step 1, estimate (A.8), and using Assumption (H3). The second term is also bounded uniformly because the function $x\mapsto\mathcal{N}(0,\Gamma)(x)\lvert x\rvert^{q}$ is uniformly bounded in $x$, by a value depending only on $\Gamma$ and $q$.
Step 5. Obtaining the estimate (A.6)
The claimed inequality is equivalent to
$\sup_{u_{1},u_{2},y}\frac{\bigl{\lvert}p(u_{1},y)-p(u_{2},y)\bigr{\rvert}}{\lvert u_{1}-u_{2}\rvert}\max\Bigl{\{}\min\left\{1+\lvert u_{1}\rvert^{q},1+\lvert u_{2}\rvert^{q}\right\},1+\lvert y\rvert^{q}\Bigr{\}}<\infty,$
where the supremum is over ×${}^{d_{u}}\times^{d_{u}}\times^{d_{y}}$. By (A.12), it holds that
$\sup_{u_{1},u_{2},y}\frac{\bigl{\lvert}p(u_{1},y)-p(u_{2},y)\bigr{\rvert}}{\lvert u_{1}-u_{2}\rvert}\min\left\{1+\lvert u_{1}\rvert^{q},1+\lvert u_{2}\rvert^{q}\right\}<\infty,$
so it remains to show that
$\sup_{u_{1},u_{2},y}\frac{\bigl{\lvert}p(u_{1},y)-p(u_{2},y)\bigr{\rvert}}{\lvert u_{1}-u_{2}\rvert}\lvert y\rvert^{q}<\infty.$
By (A.13), it is clear that the supremum is uniformly bounded if restricted to the set $\lvert u_{1}-u_{2}\rvert\geqslant 1$, so it suffices to show that
$\sup_{(u,\delta,y)\in^{d_{u}}\times B(0,1)\times^{d_{y}}}\frac{\bigl{\lvert}p(u,y)-p(u+\delta,y)\bigr{\rvert}}{\lvert\delta\rvert}\lvert y\rvert^{q}<\infty,$
where $B(0,1)$ is the open ball of radius 1 centered radius at the origin in ${}^{d_{u}}$. We use again (A.14) in order to bound
$\frac{\bigl{\lvert}p(u,y)-p(u+\delta,y)\bigr{\rvert}}{\lvert\delta\rvert}\lvert y\rvert^{q}$ $\leqslant C\frac{\bigl{\lvert}p(u,y)-p(u+\delta,y)\bigr{\rvert}}{\lvert\delta\rvert}\bigl{\lvert}h(u)\bigr{\rvert}^{q}$
$\qquad+C\frac{\bigl{\lvert}p(u,y)-p(u+\delta,y)\bigr{\rvert}}{\lvert\delta\rvert}\lvert y-h(u)\rvert^{q}.$
The first term is bounded uniformly by (A.12) and the assumption that $\bigl{\lvert}h(u)\bigr{\rvert}\leqslant\kappa_{h}\bigl{(}1+\lvert u\rvert\bigr{)}$. To conclude the proof, it remains to show that
$\sup_{(u,\delta,y)\in^{d_{u}}\times B(0,1)\times^{d_{y}}}C\frac{\bigl{\lvert}p(u,y)-p(u+\delta,y)\bigr{\rvert}}{\lvert\delta\rvert}\lvert y-h(u)\rvert^{q}<\infty.$ (A.15)
By (A.12) it holds that
$\bigl{\lvert}p(u,y)-p(u+\delta,y)\bigr{\rvert}$
$\qquad\leqslant\frac{C\lvert\delta\rvert}{1+\lvert u\rvert^{q}}\max\left\{\exp\left(-\frac{1}{4}\lvert y-h(u)\rvert_{\Gamma}^{2}\right),\exp\left(-\frac{1}{4}\lvert y-h(u+\delta)\rvert_{\Gamma}^{2}\right)\right\}.$
$\qquad\leqslant\frac{C\lvert\delta\rvert}{1+\lvert u\rvert^{q}}\exp\left(-\frac{1}{8}\lvert y-h(u)\rvert_{\Gamma}^{2}+\frac{1}{4}\Big{\lvert}h(u)-h(u+\delta)\Big{\rvert}_{\Gamma}^{2}\right).$
In the last line we used the inequality $|a+b|^{2}\geqslant\frac{1}{2}|a|^{2}-|b|^{2}$, which follows from Young’s inequality. Since the function $x\mapsto\operatorname{e}^{-\frac{x^{2}}{8}+C}x^{q}$ is bounded uniformly in $x\in$ and by Assumption (H4), the bound (A.15) easily follows, concluding the proof. ∎
Appendix B Technical Results for Theorem 2.1
We show moment bounds in Subsection B.1, we recall that on Gaussian measures the action of the conditioning map and the Kalman transport map are equivalent in Subsection B.2, and we prove stability results in Subsection B.3.
B.1 Moment Bounds
Lemma B.1 (Moment Bounds).
Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumption H, it holds that
$\bigl{\lvert}\mathcal{M}(\mathsf{P}\mu)\bigr{\rvert}\leqslant\kappa_{\Psi}\bigl{(}1+\mathcal{M}_{1}(\mu)\bigr{)},\qquad\Sigma\preccurlyeq\mathcal{C}(\mathsf{P}\mu)\preccurlyeq\Sigma+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}I_{d_{u}}.$ (B.1)
Proof.
Using the definition of $\mathsf{P}$ in (1.7), it holds that
$\mathcal{M}(\mathsf{P}\mu)=\int_{{}^{d_{u}}}u\,\mathsf{P}\mu(u)\,\mathrm{d}u$ $=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\int_{{}^{d_{u}}}u\exp\left(-\frac{1}{2}\big{\lvert}u-\Psi(v)\big{\rvert}_{\Sigma}^{2}\right)\,\mu(\mathrm{d}v)\,\mathrm{d}u$
$=\int_{{}^{d_{u}}}\Psi(v)\,\mu(\mathrm{d}v),$
where application of Fubini’s theorem yields the last equality. The first inequality in (B.1) then follows from Assumption (H2). For the second inequality in (B.1), we first note that for any $a,v\in^{d_{u}}$ it holds that
$\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u$
$\qquad\succcurlyeq\int_{{}^{d_{u}}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u;$ (B.2)
this follows from Lemma A.1 by noting that $\Psi(v)$ is the mean under the Gaussian probability measure $\mathcal{N}\bigl{(}\Psi(v),\Sigma\bigr{)}.$ Letting now $a$ be the mean under the measure $\mathsf{P}\mu$, we observe that, by conditioning on $v$ and using (B.1),
$\mathcal{C}(\mathsf{P}\mu)$ $=\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\,\mathsf{P}\mu(u)\,\mathrm{d}u$
$=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
$\succcurlyeq\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
$=\int_{{}^{d_{u}}}\Sigma\,\mu(\mathrm{d}v)=\Sigma,$
hence $\mathcal{C}(\mathsf{P}\mu)\succcurlyeq\Sigma$. On the other hand, using Lemma A.1 again and noting that, by the Cauchy-Schwarz inequality, $ww^{\mathsf{T}}\preccurlyeq(w^{\mathsf{T}}w)I_{d_{u}}$ for any vector $w\in^{d_{u}}$, we deduce that
$\mathcal{C}(\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d_{u}}}u\otimes u\,\mathsf{P}\mu(u)\,\mathrm{d}u$
$=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\int_{{}^{d_{u}}}u\otimes u\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\mu(\mathrm{d}v)\,\mathrm{d}u$
$=\int_{{}^{d_{u}}}\bigl{(}\Psi(v)\otimes\Psi(v)+\Sigma\bigr{)}\,\mu(\mathrm{d}v)$
$\preccurlyeq\Sigma+\left(\int_{{}^{d_{u}}}|\Psi(v)|^{2}\mu(\mathrm{d}v)\right)I_{d_{u}}\preccurlyeq\Sigma+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}I_{d_{u}},$ (B.3)
which yields the desired result. ∎
Lemma B.2.
Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumption H, it holds that
$\bigl{\lvert}\mathcal{M}^{u}(\mathsf{Q}\mathsf{P}\mu)\bigr{\rvert}$ $\leqslant\kappa_{\Psi}\bigl{(}1+\mathcal{M}_{1}(\mu)\bigr{)},$ (B.4a)
$\bigl{\lvert}\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)\bigr{\rvert}$ $\leqslant\kappa_{h}\sqrt{2\Bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\Bigr{)}}.$ (B.4b)
Furthermore, it holds that
$\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)$ $\preccurlyeq\begin{pmatrix}4\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}I_{d_{u}}+2\Sigma&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&4\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}+\Gamma\end{pmatrix},$ (B.5a)
$\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)$ $\succcurlyeq\frac{\gamma\cdot\min\Bigl{\{}2\sigma,\gamma+4\kappa_{h}^{2}\Bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\Bigr{)}\Bigr{\}}}{2\gamma+8\kappa_{h}^{2}\Bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\Bigr{)}}I_{d_{u}+d_{y}}.$ (B.5b)
Proof.
The inequalities (B.4a) and (B.4b) follow from Assumption H and the fact that
$\mathcal{M}(\mathsf{Q}\mathsf{P}\mu)=\begin{pmatrix}\mathcal{M}(\mathsf{P}\mu)\\ \mathsf{P}\mu[h]\end{pmatrix}.$ (B.6)
Indeed, from Lemma B.1 we know that $\bigl{\lvert}\mathcal{M}(\mathsf{P}\mu)\bigr{\rvert}\leqslant\kappa_{\Psi}\bigl{(}1+\mathcal{M}_{1}(\mu)\bigr{)},$ which leads by Jensen’s inequality to (B.4a). To deduce (B.4b), we note that
$\bigl{|}P\mu[h]\bigr{|}^{2}$ $\leqslant\int_{{}^{d_{u}}}|h(u)|^{2}\,\mathsf{P}\mu(\mathrm{d}u)$
$\leqslant 2\kappa_{h}^{2}+2\kappa_{h}^{2}\int_{{}^{d_{u}}}|u|^{2}\,\mathsf{P}\mu(\mathrm{d}u)$ (B.7a)
$=2\kappa_{h}^{2}+2\kappa_{h}^{2}\int_{{}^{d_{u}}}\Bigl{(}|\Psi(v)|^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}\,\mu(\mathrm{d}v),$ (B.7b)
where (B.7a) follows by applying Assumption (H3) and Young’s inequality and (B.7b) follows by applying the properties of the Gaussian transition density. The result then follows by applying Assumption (H2).
We now turn our attention to establishing inequality (B.5a). To this end, let $\phi\colon^{d_{u}}\to^{d_{u}+d_{y}}$ define the map $\phi(u)=\bigl{(}u,h(u)\bigr{)}$, and let $\phi_{\sharp}$ denote the associated pushforward map on measures. We note that
$\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)=\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)+\begin{pmatrix}0_{d_{u}\times d_{u}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&\Gamma\end{pmatrix}=\begin{pmatrix}\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)&\mathcal{C}^{uy}(\phi_{\sharp}\mathsf{P}\mu)\\ \mathcal{C}^{yu}(\phi_{\sharp}\mathsf{P}\mu)&\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)+\Gamma\end{pmatrix}.$ (B.8)
It holds for any $(a,b)\in^{d_{u}}\times^{d_{y}}$ that
$2\begin{pmatrix}aa^{\mathsf{T}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&bb^{\mathsf{T}}\end{pmatrix}-\begin{pmatrix}a\\ b\end{pmatrix}\otimes\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}a\\ -b\end{pmatrix}\otimes\begin{pmatrix}a\\ -b\end{pmatrix}\succcurlyeq 0_{(d_{u}+d_{y})\times(d_{u}+d_{y})}.$
Hence, it follows that
$\begin{pmatrix}a\\ b\end{pmatrix}\otimes\begin{pmatrix}a\\ b\end{pmatrix}\preccurlyeq 2\begin{pmatrix}aa^{\mathsf{T}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&bb^{\mathsf{T}}\end{pmatrix},$
which enables to deduce, using Lemmas A.1 and B.1 and Assumption H, that
$\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d_{u}}}\begin{pmatrix}u\\ h(u)\end{pmatrix}\otimes\begin{pmatrix}u\\ h(u)\end{pmatrix}\,\mathsf{P}\mu(u)\,\mathrm{d}u$
$\preccurlyeq 2\int_{{}^{d_{u}}}\begin{pmatrix}uu^{\mathsf{T}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&h(u)h(u)^{\mathsf{T}}\end{pmatrix}\,\mathsf{P}\mu(u)\,\mathrm{d}u$
$\preccurlyeq 2\begin{pmatrix}\Sigma+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}I_{d_{u}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}\end{pmatrix},$ (B.9)
where we used (B.3) and the calculation resulting in (B.7a) in the last inequality. Noting this inequality in combination with (B.8) yields the upper bound (B.5a). To establish the lower bound (B.5b), we first note that by the Cauchy–Schwarz inequality, for any $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
$\left\lvert a^{\mathsf{T}}\mathcal{C}^{uy}(\pi)b\right\rvert$ =×$=\int_{{}^{d_{u}}\times^{d_{y}}}\Bigl{(}a^{\mathsf{T}}\bigl{(}u-\mathcal{M}^{u}(\pi)\bigr{)}\Bigr{)}\Bigl{(}b^{\mathsf{T}}\bigl{(}y-\mathcal{M}^{y}(\pi)\bigr{)}\Bigr{)}\pi(\mathrm{d}u\mathrm{d}y)$
$\leqslant\sqrt{a^{\mathsf{T}}\mathcal{C}^{uu}(\pi)a}\,\sqrt{b^{\mathsf{T}}\mathcal{C}^{yy}(\pi)b}.$
Hence, by Young’s inequality, we deduce that for all $\varepsilon\in(0,1)$ and for all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)a-\left(\frac{1}{\varepsilon}-1\right)b^{\mathsf{T}}\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)b$
$\geqslant(1-\varepsilon)a^{\mathsf{T}}\Sigma a-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]},$
where we applied (B.1) and the bound $\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)\preccurlyeq\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}I_{d_{y}}$ in the last inequality. We then use (B.8) to deduce that
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\Sigma a-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}+b^{\mathsf{T}}\Gamma b$
$\geqslant(1-\varepsilon)\sigma\left\lvert a\right\rvert^{2}+\left(\gamma-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\right)\lvert b\rvert^{2}.$
Taking $\varepsilon=\frac{2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}$, so that the coefficient of $\lvert b\rvert^{2}$ is $\frac{\gamma}{2}$, we obtain
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}\geqslant\frac{\gamma\sigma}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}\lvert a\rvert^{2}+\frac{\gamma}{2}\lvert b\rvert^{2}\geqslant\min\left\{\frac{\gamma\sigma}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}},\frac{\gamma}{2}\right\}\left(\lvert a\rvert^{2}+\lvert b\rvert^{2}\right).$ (B.10)
Since this is true for any $(a,b)\in^{d_{u}}\times^{d_{y}}$, it follows that
$\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\succcurlyeq\gamma\frac{\min\left\{2\sigma,\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\right\}}{2\gamma+4\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}I_{d_{u}+d_{y}}.$
In order to finally obtain (B.5b), we use (B.7b) to deduce that
$\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\leqslant 2\kappa_{h}^{2}\Bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\Bigr{)},$
which concludes the proof. ∎
Lemma B.3 (Moment Bound for the True Filtering Distribution).
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (1.11) initialized at the probability measure $\mu_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th order polynomial moment $\mathcal{M}_{q}(\mu_{0})<\infty$. If Assumption H holds then there exists $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$ such that
$\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}(\mu_{j})\leqslant C.$
Proof.
We have $\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}$, for $j\in\llbracket 0,J-1\rrbracket$. Equivalently, using the notation from (1.6), it holds that $\mu_{j+1}=\mathsf{L}_{j}\mathsf{P}\mu_{j}$ for each $j\in\llbracket 0,J-1\rrbracket$. Hence, we note that
$\mu_{j+1}=\frac{\displaystyle\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mathsf{P}\mu_{j}(u)}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mathsf{P}\mu_{j}(U)\,\mathrm{d}U}.$ (B.11)
It is readily observed that
$\mathcal{M}_{q}(\mu_{j+1})$ $=\frac{\displaystyle\int_{{}^{d_{u}}}|u|^{q}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mathsf{P}\mu_{j}(u)\mathrm{d}u}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mathsf{P}\mu_{j}(U)\,\mathrm{d}U}$
$\leqslant\frac{\displaystyle\int_{{}^{d_{u}}}|u|^{q}\,\mathsf{P}\mu_{j}(u)\mathrm{d}u}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\,\mathsf{P}\mu_{j}(U)\,\mathrm{d}U}.$ (B.12)
We first bound from above the numerator of (B.12); indeed, note that
$\int_{{}^{d_{u}}}|u|^{q}\mathsf{P}\mu_{j}(u)\mathrm{d}u$ $=\int_{{}^{d_{u}}}|u|^{q}\left(\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\,\mu_{j}(\mathrm{d}v)\right)\mathrm{d}u$
$=\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}|u|^{q}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\,\mathrm{d}u\right)\mu_{j}(\mathrm{d}v)$ (B.13a)
$\leqslant C\int_{{}^{d_{u}}}\bigl{(}1+\left\lvert\Psi(v)\right\rvert^{q}\bigr{)}\mu_{j}(\mathrm{d}v)$ (B.13b)
$\leqslant C\Bigl{(}1+\mathcal{M}_{q}(\mu_{j})\Bigr{)},$ (B.13c)
where (B.13a) follows from Fubini’s theorem, the inequality (B.13b) from properties of Gaussians, and (B.13c) from application of Assumption H. We note that in (B.13c) the constant $C$ depends on $\kappa_{\Psi},\Sigma$. Now, to obtain a lower bound on the denominator of (B.12), we observe that
$\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mathsf{P}\mu_{j}(u)\,\mathrm{d}u$ $\geqslant\exp\left(-\frac{1}{2}\int_{{}^{d_{u}}}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\mathsf{P}\mu_{j}(u)\right)$
$\geqslant C\exp\left(-\left\lVert\Gamma^{-1}\right\rVert\int_{{}^{d_{u}}}\lvert h(u)\rvert^{2}\mathsf{P}\mu_{j}(u)\right)$
$\geqslant C\exp\Bigl{(}-4\kappa^{2}_{h}\kappa^{2}_{\Psi}\left\lVert\Gamma^{-1}\right\rVert\mathcal{M}_{2}(\mu_{j})\Bigr{)},$ (B.14)
where the first inequality follows by application of Jensen’s inequality and (B.14) follows from the calculation leading to (B.7b). We note that the $C$ in (B.14) is a constant depending on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$. Therefore, by combining (B.14) with (B.13c), it is possible to deduce from (B.12) that
$\mathcal{M}_{q}(\mu_{j+1})\leqslant C\exp\Bigl{(}4\kappa^{2}_{h}\kappa^{2}_{\Psi}\left\lVert\Gamma^{-1}\right\rVert\mathcal{M}_{2}(\mu_{j})\Bigr{)}\Bigl{(}1+\mathcal{M}_{q}(\mu_{j})\Bigr{)},$ (B.15)
where $C$ is a constant depending on $\kappa_{h},\kappa_{\Psi},\kappa_{y},\Sigma,\Gamma$. ∎
Remark B.4.
  • In some situations, the bound (B.15) is overly pessimistic. For example, if $h$ satisfies Assumption (H3) as well as the inequality $\bigl{\lvert}h(u)-y^{\dagger}_{j+1}\bigr{\rvert}\geqslant c_{\ell}\left(|u|-1\right)$ for all $u\in^{d_{u}}$ for some positive $c_{\ell}$, then by [2023arXiv231207373J, Proposition A.3] for all $q>0$ there is $C=C(c_{\ell},q)$ such that
    $\forall\mu\in\mathcal{P}(^{d_{u}}),\qquad\mathsf{L}_{j}\mu\Bigl{[}|x|^{q}\Bigr{]}\leqslant C\mu\Bigl{[}|x|^{q}\Bigr{]}.$
    In this setting, better control of the moments can be achieved than in (B.15).
  • In obtaining (B.15), we did not use any information on $\mathsf{P}\mu_{j}$ other than the moment bound (B.13c). With this approach, the presence of an exponential in the bound (B.15) is to be expected. Indeed, consider the case where $d_{u}=2$, $d_{y}=1$, $\Gamma=\frac{1}{4}$, $y^{\dagger}_{j+1}=0$, and
    $h(u)=u_{1},\qquad u=\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix},\qquad\mu_{R}=\frac{x_{R}^{2}\delta_{(R,0)}+\delta_{(0,x_{R})}}{x_{R}^{2}+1},\qquad x_{R}=\operatorname{e}^{\frac{R^{2}}{2}}.$
    Then, as $R\to\infty$, $\mu_{R}\left[|u|^{2}\right]\sim R^{2}$, while
    $\mathsf{L}_{j}\mu_{R}\left[|u|^{2}\right]=\frac{\int_{{}^{2}}(u_{1}^{2}+u_{2}^{2})\exp\left(-2u_{1}^{2}\right)\,\mu_{R}(\mathrm{d}u_{1}\mathrm{d}u_{2})}{\int_{{}^{2}}\exp\left(-2u_{1}^{2}\right)\,\mu_{R}(\mathrm{d}u_{1}\mathrm{d}u_{2})}=\frac{x_{R}^{2}\operatorname{e}^{-2R^{2}}R^{2}+x_{R}^{2}}{x_{R}^{2}\operatorname{e}^{-2R^{2}}+1}\sim x_{R}^{2}=\operatorname{e}^{R^{2}}.$
    Thus, for large $R$, the second moment of $\mathsf{L}_{j}\mu_{R}$ is approximately equal to the exponential of the second moment of $\mu_{R}$.
Lemma B.5 (Moment Bound for the Approximate Filtering Distribution).
Assume that the probability measures $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (2.4) initialized at the probability measure $\mu^{\rm EK}_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th polynomial order moment $\mathcal{M}_{q}\bigl{(}\mu_{0}^{\rm EK}\bigr{)}<\infty$. If Assumption H holds then there exists $C=C\Bigl{(}\mathcal{M}_{q}\bigl{(}\mu_{0}^{\rm EK}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ such that
$\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}\bigl{(}\mu^{\rm EK}_{j}\bigr{)}\leqslant C.$
Proof.
We begin by noting that $\mu^{\rm EK}_{j+1}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu^{\rm EK}_{j},$ for $j\in\llbracket 0,J-1\rrbracket$. Thus
$\mathcal{M}_{q}\bigl{(}\mu^{\rm EK}_{j+1}\bigr{)}$ $=\int_{{}^{d_{u}}}|u|^{q}\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}(\mathrm{d}u)$
=×$=\int_{{}^{d_{u}}\times^{d_{y}}}\Big{\lvert}\mathscr{T}(u,y;\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK},y^{\dagger}_{j+1})\Big{\rvert}^{q}\,\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}(u,y)\,\mathrm{d}y\,\mathrm{d}u$
=×$=\int_{{}^{d_{u}}\times^{d_{y}}}\Big{\lvert}u+\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK})\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK})^{-1}\bigl{(}y^{\dagger}_{j+1}-y\bigr{)}\Big{\rvert}^{q}\,\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}(u,y)\,\mathrm{d}y\,\mathrm{d}u$
$\leqslant C\int_{{}^{d_{u}}\times^{d_{y}}}|u|^{q}\,\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}(u,y)\,\mathrm{d}y\,\mathrm{d}u$
$\qquad\qquad+C\Bigl{(}1+\mathcal{M}_{2}\bigl{(}\mu^{\rm EK}_{j}\bigr{)}\Bigr{)}^{2q}\cdot\left(1+\int_{{}^{d_{u}}\times^{d_{y}}}|y|^{q}\,\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}(u,y)\,\mathrm{d}y\,\mathrm{d}u\right),$ (B.16)
where in (B.16) the constant $C$ depends on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$ and where the dependence on the second moment of $\mu_{j}^{\rm EK}$ in the second term is derived from (B.5a) and (B.5b). By using the definitions of $\mathsf{Q}$ and $\mathsf{P}$ and by applying reasoning analogous to (B.13c), we deduce that
$\mathcal{M}_{q}\bigl{(}\mu^{\rm EK}_{j+1}\bigr{)}$ $\leqslant C\Bigl{(}1+\mathcal{M}_{2}\bigl{(}\mu^{\rm EK}_{j}\bigr{)}\Bigr{)}^{2q}\cdot\Bigl{(}1+\mathcal{M}_{q}\bigl{(}\mu^{\rm EK}_{j}\bigr{)}\Bigr{)},$
$\leqslant C\Bigl{(}1+\mathcal{M}_{q}\bigl{(}\mu^{\rm EK}_{j}\bigr{)}\Bigr{)}^{2q+1}$
where $C$ is a constant depending on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$. Iteration gives the desired result. ∎
Lemma B.6.
For $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ with finite second moments, it holds that
$\bigl{\lvert}\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\bigr{\rvert}$ $\leqslant\frac{1}{2}d_{g}(\mu_{1},\mu_{2}),$
$\bigl{\lVert}\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\bigr{\rVert}$ $\leqslant\left(1+\frac{1}{2}\lvert\mathcal{M}(\mu_{1})+\mathcal{M}(\mu_{2})\rvert\right)\,d_{g}(\mu_{1},\mu_{2}).$
Proof.
The lemma as stated may be found in [carrillo2022ensemble, Lemma 20], where a complete proof is given. ∎
B.2 Action of $\mathsf{T}_{j}$ on Gaussians
Lemma B.7 ($\mathsf{B}_{j}\mathsf{G}=\mathsf{T}_{j}\mathsf{G}$).
Fix $y^{\dagger}_{j+1}\in^{d_{y}}$ and let $\pi$ be a Gaussian measure over ×${}^{d_{u}}\times^{d_{y}}$. Then the probability measure $\mathsf{B}_{j}\pi$, with $\mathsf{B}_{j}$ defined in (1.10), coincides with the probability measure $\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi$, where $\mathscr{T}$ is defined in (2.3).
Proof.
The lemma as stated may be found in [carrillo2022ensemble, Lemma 21], where a complete proof is given. ∎
B.3 Stability Results
Lemma B.8 (Map $\mathsf{P}$ is Lischitz).
Suppose that $\Psi$ and $\Sigma$ satisfy Assumptions (H2) and (H5), respectively. Then it holds that
$\forall(\mu,\nu)\in\mathcal{P}(^{d_{u}})\times\mathcal{P}(^{d_{u}}),\qquad d_{g}(\mathsf{P}\mu,\mathsf{P}\nu)\leqslant\Bigl{(}1+2\kappa_{\Psi}^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}\,d_{g}(\mu,\nu).$
Proof.
By definition of $\mathsf{P}$, it holds that
$\mathsf{P}\mu(\mathrm{d}u)=\int_{{}^{d_{u}}}p(v,\mathrm{d}u)\,\mu(\mathrm{d}v),\qquad\text{ where }p(v,\mathrm{d}u):=\frac{\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\,\mathrm{d}u.$
Let $g(u)=1+|u|^{2}$, as in Definition 1.1, and take any function $f\colon^{d_{u}}\to$ such that $|f|\leqslant g$. Assumption (H2) implies that
$\left\lvert\int_{{}^{d_{u}}}f(u)\,p(v,\mathrm{d}u)\right\rvert\leqslant\int_{{}^{d_{u}}}g(u)\,p(v,\mathrm{d}u)$ $=1+\left\lvert\Psi(v)\right\rvert^{2}+\operatorname*{tr}(\Sigma)$
$\leqslant 1+2\kappa_{\Psi}^{2}+2\kappa_{\Psi}^{2}\lvert v\rvert^{2}+\operatorname*{tr}(\Sigma)\leqslant\Bigl{(}1+2\kappa_{\Psi}^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}g(v).$
By Fubini’s theorem, it follows that
$\Bigl{\lvert}\mathsf{P}\mu[f]-\mathsf{P}\nu[f]\Bigr{\rvert}$ $=\biggl{\lvert}\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}f(u)\,p(v,\mathrm{d}u)\right)\bigl{(}\mu(\mathrm{d}v)-\nu(\mathrm{d}v)\bigr{)}\biggr{\rvert}$
$=\Bigl{(}1+2\kappa_{\Psi}^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}\Bigl{\lvert}\mu[g]-\nu[g]\Bigr{\rvert}\leqslant\Bigl{(}1+2\kappa_{\Psi}^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}\,d_{g}(\mu,\nu),$
which concludes the proof. ∎
Lemma B.9 (Map $\mathsf{Q}$ is Lipschitz).
Suppose that $h$ and $\Gamma$ satisfy Assumptions (H3) and (H5), respectively. Then it holds that
$\forall(\mu,\nu)\in\mathcal{P}(^{d_{u}})\times\mathcal{P}(^{d_{u}}),\qquad d_{g}(\mathsf{Q}\mu,\mathsf{Q}\nu)\leqslant\Bigl{(}1+2\kappa_{h}^{2}+\operatorname*{tr}(\Gamma)\Bigr{)}d_{g}(\mu,\nu).$ (B.17)
Proof.
Take $f\colon^{d_{u}}\times^{d_{y}}\to$ satisfying $|f|\leqslant g$, where $g(u,v)=1+|u|^{2}+|y|^{2}$. By Fubini’s theorem, it holds that
$\mathsf{Q}\mu[f]-\mathsf{Q}\nu[f]=\int_{{}^{d_{u}}}\Pi f(u)\,\Bigl{(}\mu(\mathrm{d}u)-\nu(\mathrm{d}u)\Bigr{)},\qquad\text{ where }\Pi f(u):=\int_{{}^{d_{y}}}f(u,y)\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(\mathrm{d}y).$
The function $\Pi f\colon^{d_{u}}\to$ satisfies
$\forall u\in^{d_{u}},\qquad\left\lvert\Pi f(u)\right\rvert$ $\leqslant\int_{{}^{d_{y}}}\bigl{\lvert}f(u,y)\bigr{\rvert}\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(\mathrm{d}y)$
$\leqslant\int_{{}^{d_{y}}}\left(1+|u|^{2}+|y|^{2}\right)\,\mathcal{N}\bigl{(}h(u),\Gamma\bigr{)}(\mathrm{d}y)$
$=1+|u|^{2}+\bigl{\lvert}h(u)\bigr{\rvert}^{2}+\operatorname*{tr}(\Gamma)\leqslant\Bigl{(}1+2\kappa_{h}^{2}+\operatorname*{tr}(\Gamma)\Bigr{)}\left(1+\lvert u\rvert^{2}\right).$
Therefore, we deduce that
$\Bigl{\lvert}\mathsf{Q}\mu[f]-\mathsf{Q}\nu[f]\Bigr{\rvert}\leqslant(1+2\kappa_{h}^{2}+\operatorname*{tr}(\Gamma)\Bigr{)}\,d_{g}(\mu,\nu),$
which concludes the proof. ∎
Lemma B.10 (Stability of Map $\mathsf{B}_{j}$).
Under Assumption H, for any probability measure $\mu\in\mathcal{P}(^{d_{u}})$ with $\mathcal{M}_{2}(\mu)<\infty$ there exists a constant $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}>0$ such that
$\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)\leqslant C_{\mathsf{B}}d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
Proof.
For ease of exposition we write $y^{\dagger}=y^{\dagger}_{j+1}$. We define the $y$-marginal densities
$\alpha_{\mu}(y):=\int_{{}^{d_{u}}}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\,,\qquad\beta_{\mu}(y):=\int_{{}^{d_{u}}}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\,.$
It holds that
$d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)=\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\left|\frac{\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})}{\alpha_{\mu}(y^{\dagger})}-\frac{\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})}{\beta_{\mu}(y^{\dagger})}\right|\,\mathrm{d}u$
$\qquad\qquad\leqslant\frac{1}{\alpha_{\mu}(y^{\dagger})}\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})-\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})\bigr{\rvert}\,\mathrm{d}u$
$\qquad\qquad\qquad+\left|\frac{\alpha_{\mu}(y^{\dagger})-\beta_{\mu}(y^{\dagger})}{\alpha_{\mu}(y^{\dagger})\beta_{\mu}(y^{\dagger})}\right|\,\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})\,\mathrm{d}u.$ (B.18)
Step 1: bounding $\alpha_{\mu}(y^{\dagger})$ and $\beta_{\mu}(y^{\dagger})$ from below
The distribution $\alpha_{\mu}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ is Gaussian with mean $\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)$ and covariance
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)=\Gamma+\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\otimes h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]}-\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]}\otimes\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]},$ (B.19)
which by Assumption H is bounded from below by $\Gamma$ and from above by $\Gamma+2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}$. To deduce this we have again made use of the fact that $aa^{\mathsf{T}}\preccurlyeq(a^{\mathsf{T}}a)I_{d}$ for any vector $a\in^{d_{u}}$, by the Cauchy-Schwarz inequality. Therefore, noting that (B.4b) implies that $\left\lvert\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)\right\rvert\leqslant C$ for a constant $C$ depending only on the parameters $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma$, we obtain
$\alpha_{\mu}(y)$ $=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\bigl{(}\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)\bigr{)}}}\exp\left(-\frac{1}{2}\bigl{(}y-\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)\bigr{)}^{\mathsf{T}}\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)^{-1}\bigl{(}y-\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)\bigr{)}\right)$
$\geqslant\frac{\exp\Bigl{(}-\frac{1}{2}\bigl{(}|y|+C\bigr{)}^{2}\left\lVert\Gamma^{-1}\right\rVert\Bigr{)}}{\sqrt{(2\pi)^{d_{y}}\det\bigl{(}\Gamma+2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}\bigr{)}}}.$ (B.20)
The function $\beta_{\mu}$ can be bounded from below using similar reasoning. Indeed, by Assumption H we have that for all $y\in^{d_{y}}$,
$\beta_{\mu}(y)$ $=\int_{{}^{d_{u}}}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u=\int_{{}^{d_{u}}}\frac{\exp\left(-\frac{1}{2}\bigl{(}y-h(u)\bigr{)}^{\mathsf{T}}\Gamma^{-1}\bigl{(}y-h(u)\bigr{)}\right)}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma)}}\mathsf{P}\mu(u)\,\mathrm{d}u$
$\geqslant\frac{C\exp\Bigl{(}-\left\lVert\Gamma^{-1}\right\rVert|y|^{2}-\left\lVert\Gamma^{-1}\right\rVert\kappa^{2}_{h}\int_{{}^{d_{u}}}|u|^{2}\mathsf{P}\mu\mathrm{d}u\Bigr{)}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma)}}$ (B.21a)
$\geqslant C\exp\Bigl{(}-\left\lVert\Gamma^{-1}\right\rVert|y|^{2}\Bigr{)},$ (B.21b)
where we applied Jensen’s inequality in (B.21a) and the constant in (B.21b) depends on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$.
Step 2: bounding the first term in (B.18)
We begin by noting that for fixed $u\in^{d_{u}}$, the functions $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)$ and $y\mapsto\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)$ are Gaussians up to constant factors. The covariance matrix of the former is $\Gamma$; it is possible to deduce that the covariance of the latter is given by
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)-\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu),$ (B.22)
where we have used the formula for the covariance of the conditional distribution of a Gaussian. We note that $\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ is positive semi-definite, hence by (B.19) and its upper bound it follows that the matrix (B.22) is bounded from above by ${2}\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}+\Gamma$. With the same notation as in (B.8), we find that
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)-\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ (B.23)
$\qquad=\Gamma+\Bigl{(}\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)-\mathcal{C}^{yu}(\phi_{\sharp}\mathsf{P}\mu)\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\phi_{\sharp}\mathsf{P}\mu)\Bigr{)}\succcurlyeq\Gamma;$ (B.24)
here, to deduce that the term in brackets is positive semi-definite, we observe that the term is the Schur complement of the block $\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)$ of the matrix $\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$. Hence, (B.22) is bounded from below by $\Gamma$. Therefore the integral in the first term of (B.18) may be bounded from above by applying the result of Lemma A.2 in Appendix A with parameter $\alpha=\alpha\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma\bigr{)}$, yielding
$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\left|\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\right|\mathrm{d}u$
$\qquad\leqslant C\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
$\qquad\leqslant C\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\left(1+|u|^{2}+|y|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
$\qquad=Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu),$ (B.25)
where $C$ depends on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$.
Step 3: bounding the second term in (B.18)
By (B.25), it holds that
$\bigl{\lvert}\alpha_{\mu}(y)-\beta_{\mu}(y)\bigr{\rvert}$ $=\left|\int_{{}^{d_{u}}}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\right|$
$\leqslant\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
Now, since $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)/\mathsf{P}\mu(u)$ is a Gaussian density with covariance $\Gamma$, which is bounded uniformly from above by $\bigl{(}(2\pi)^{d_{y}}\det(\Gamma)\bigr{)}^{-1/2}$, we also have that
$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$ $=\beta_{\mu}(y)+\int_{{}^{d_{u}}}|u|^{2}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$
$\leqslant\beta_{\mu}(y)+\int_{{}^{d_{u}}}\frac{|u|^{2}\mathsf{P}\mu(u)}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}\,\mathrm{d}u$
$=\beta_{\mu}(y)+\frac{\operatorname*{tr}\bigl{(}\mathcal{C}(\mathsf{P}\mu)\bigr{)}+\left\lvert\mathcal{M}(\mathsf{P}\mu)\right\rvert^{2}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}.$
We proceed by noting the result of Lemma B.1, which gives that the first and second moments of $\mathsf{P}\mu$ are bounded from above by a constant which depends on $\mathcal{M}_{2}(\mu),\kappa_{\Psi},\Sigma$.
Step 4: concluding the proof
Putting together the above bounds, we conclude that
$d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)\leqslant\frac{C\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}}{\alpha_{\mu}(y^{\dagger}_{j+1})}\left(1+\frac{1}{\beta_{\mu}(y^{\dagger}_{j+1})}\right)d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
Applying the inequalities (B.20) and (B.21b) yields the desired result. ∎
Lemma B.11 (Stability of Map $\mathsf{T}_{j}$).
Suppose that Assumption H is satisfied. Then, for all $R\geqslant 1$, it holds that for any $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, there is $L_{\mathsf{T}}:=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$, such that
$\forall j\in\llbracket 1,J\rrbracket,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p).$
Proof.
By the results of Lemma B.2, it holds that the probability measure $\mathsf{Q}\mathsf{P}\mu$ belongs to $\mathcal{P}_{\widetilde{R}}(^{d_{u}}\times^{d_{y}})$ for some $\widetilde{R}\geqslant 1$. Define
$r=\max\Bigl{\{}R,\widetilde{R},\kappa_{y}\Bigr{\}}.$
We denote by $\mathscr{T}^{\pi}$ and $\mathscr{T}^{p}$ the affine maps corresponding to evaluation of covariance information at the probability measures $\pi$ and $p=\mathsf{Q}\mathsf{P}\mu$, respectively. Namely, letting $K=\mathcal{C}(\pi)$, $S=\mathcal{C}(p)$ and $y^{\dagger}=y^{\dagger}_{j+1}$, we write
$\mathscr{T}^{\pi}(u,y)$ $=u+A_{\pi}(y^{\dagger}-y),\qquad$ $A_{\pi}:=K_{uy}K_{yy}^{-1},$
$\mathscr{T}^{p}(u,y)$ $=u+A_{p}(y^{\dagger}-y),\qquad$ $A_{p}:=S_{uy}S_{yy}^{-1}.$
By a straightforward application of the triangle inequality, it holds that
$d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant d_{g}(\mathscr{T}^{\pi}_{\sharp}\pi,\mathscr{T}^{\pi}_{\sharp}p)+d_{g}(\mathscr{T}^{\pi}_{\sharp}p,\mathscr{T}^{p}_{\sharp}p).$ (B.26)
Before bounding each term on the right-hand side of B.26, we derive some inequalities that we will use in the proof. It holds that
$\lVert A_{\pi}\rVert\leqslant\lVert K_{uy}\rVert\lVert K_{yy}^{-1}\rVert\leqslant\lVert K\rVert\lVert K^{-1}\rVert\leqslant r^{4},$ (B.27)
and similarly for $A_{p}$. This can easily be seen by noting that the matrix norm (induced by the Euclidean vector norm) of any submatrix is bounded from above by the norm of the full matrix. Using this bound again and assuming without loss of generality that $r>1$, we deduce that
$\lVert A_{\pi}-A_{p}\rVert=\lVert K_{uy}K_{yy}^{-1}-S_{uy}S_{yy}^{-1}\rVert$ $\leqslant\lVert(K_{uy}-S_{uy})K_{yy}^{-1}\rVert+\lVert S_{uy}\left(K_{yy}^{-1}-S_{yy}^{-1}\right)\rVert$
$\leqslant r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}-S_{yy}^{-1}\rVert\Bigr{)}$
$=r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}(S_{yy}-K_{yy})S_{yy}^{-1}\rVert\Bigr{)}$
$\leqslant r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}\rVert\lVert S_{yy}-K_{yy}\rVert\lVert S_{yy}^{-1}\rVert\Bigr{)}$
$\leqslant 2r^{6}\lVert K-S\rVert\leqslant 2r^{6}(1+2r)\,d_{g}(\pi,p),$ (B.28)
where the last line follows from the result in Lemma B.6.
Bounding the first term in (B.26)
Let $f\colon^{d_{u}}\to$ be a function such that $\lvert f\rvert\leqslant g$ with $g(u)=1+\lvert u\rvert^{2}$. It follows from the definition of the pushforward measure that
$\bigl{\lvert}\mathscr{T}^{\pi}_{\sharp}\pi[f]-\mathscr{T}^{\pi}_{\sharp}p[f]\bigr{\rvert}=\bigl{\lvert}\pi[f\circ\mathscr{T}^{\pi}]-p[f\circ\mathscr{T}^{\pi}]\bigr{\rvert}.$
For any $(u,y)\in^{d_{u}}\times^{d_{y}}$, we have that
$\bigl{\lvert}f\circ\mathscr{T}^{\pi}(u,y)\bigr{\rvert}$ $=\left\lvert f\Bigl{(}u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}\right\rvert\leqslant g\Bigl{(}u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}$
$=1+\bigl{\lvert}u+A_{\pi}(y^{\dagger}-y)\bigr{\rvert}^{2}\leqslant 1+3\lvert u\rvert^{2}+3\lvert A_{\pi}y^{\dagger}\rvert^{2}+3\lvert A_{\pi}y\rvert^{2}$
$\leqslant 3\left(1+\lvert A_{\pi}y^{\dagger}\rvert^{2}\right)\max\left\{1,\lVert A_{\pi}\rVert^{2}\right\}g(u,y).$
Hence, by (B.27) we conclude that
$d_{g}(\mathscr{T}^{\pi}_{\sharp}\pi,\mathscr{T}^{\pi}_{\sharp}p)\leqslant 3\left(1+r^{10}\right)r^{8}\,d_{g}(\pi,p).$ (B.29)
Bounding the second term in (B.26)
Let $f\colon^{d_{u}}\to$ again be a function such that the inequality $\lvert f\rvert\leqslant g$, with $g(u)=1+\lvert u\rvert^{2}$, is satisfied. We note that
$\Bigl{\lvert}\mathscr{T}^{\pi}_{\sharp}p[f]-\mathscr{T}^{p}_{\sharp}p[f]\Bigr{\rvert}=\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}]-p[f\circ\mathscr{T}^{p}]\Bigr{\rvert}=\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}.$
The right-hand side may be rewritten as
$\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}$ $=\left\lvert\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f\left(u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\right)-f\left(u+A_{p}\bigl{(}y^{\dagger}-y\bigr{)}\right)\,p(u,y)\,\mathrm{d}u\,\mathrm{d}y\right\rvert.$
A change of variable yields
$\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\bigl{(}f(u+A_{\pi}z)-f(u+A_{p}z)\bigr{)}\,p(u,y^{\dagger}-z)\,\mathrm{d}u\,\mathrm{d}z$
$=\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f(v)\Bigl{(}p(v-A_{\pi}z,y^{\dagger}-z)-p(v-A_{p}z,y^{\dagger}-z)\Bigr{)}\mathrm{d}v\,\mathrm{d}z.$ (B.30)
Since $\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty$ by assumption, from Lemma A.3 it follows that
$\Bigl{\lvert}p(v-A_{\pi}z,y^{\dagger}-z)-p(v-A_{p}z,y^{\dagger}-z)\Bigr{\rvert}$
$\qquad\leqslant C\lvert A_{\pi}z-A_{p}z\rvert\cdot\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\cdot$
$\qquad\qquad\cdot\frac{1}{1+\lvert y^{\dagger}-z\rvert^{\max\{3+d_{u},4+d_{y}\}}}.$
We apply this inequality to bound for fixed $z\in^{d_{y}}$ the inner integral in (B.30). Keeping only the terms that depend on $v$, we obtain that
$\int_{{}^{d_{u}}}\lvert f(v)\rvert\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\,\mathrm{d}v$
$\qquad\leqslant\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\mathrm{d}v+\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\mathrm{d}v.$
Since $\lvert f(v)\rvert\leqslant 1+\lvert v\rvert^{2}$, proceeding by substitution gives that this expression is bounded from above by
$C\Bigl{(}1+\lvert A_{\pi}z\rvert^{2}+\lvert A_{p}z\rvert^{2}\Bigr{)}\leqslant Cr^{8}\left(1+\lvert z\rvert^{2}\right),$
where the last inequality follows from (B.27). The remaining integral in the $z$ variable can be bounded in a similar way:
$d_{g}(\mathscr{T}^{\pi}_{\sharp}p,\mathscr{T}^{p}_{\sharp}p)$ $\leqslant Cr^{8}\int_{{}^{d_{y}}}\frac{1+\lvert z\rvert^{2}}{1+\lvert y^{\dagger}-z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\lvert A_{\pi}z-A_{p}z\rvert\,\mathrm{d}z$
$\leqslant Cr^{11}\lVert A_{\pi}-A_{p}\rVert\leqslant Cr^{18}d_{g}(\pi,p),$ (B.31)
where the last inequality follows from (B.28). Combining (B.29) and (B.31) yields the desired result. ∎
Appendix C Technical Results for Approximation Result in Proposition 2.2
In Lemma C.1 we recall the local Lipschitz continuity result for the operator $\mathsf{G}$ established in [carrillo2022ensemble]. Lemma C.3 establishes that the filtering distribution is a locally Lipschitz function of $(\Psi,h)$, viewed as a mapping from Banach space equipped with the $\|\cdot\|_{\infty}$ norm into the space of probability measures metrized using the $d_{g}$ distance. This is preceded by Lemma C.2 which establishes bounds used to prove this Lipschitz property. The two lemmas do not require $\Psi_{0}$ and $h_{0}$ to be affine, but simply require that they both satisfy Assumptions (H2) and (H3). This is in contrast with the more specific setting of Proposition 2.2, which imposes an affine assumption on $(\Psi_{0},h_{0})$.
Lemma C.1.
For all $R\geqslant 1$, there exists $L_{\mathsf{G}}=L_{\mathsf{G}}(R,n)$ so that for any $\mu_{1},\mu_{2}\in\mathcal{P}_{R}(^{n})$ it holds that
$d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})\leqslant L_{\mathsf{G}}(R,n)\cdot d_{g}(\mu_{1},\mu_{2}).$
Proof.
The lemma as stated may be found in [carrillo2022ensemble, Lemma 28], where a complete proof is given. ∎
Lemma C.2.
Suppose that the matrices $(\Sigma,\Gamma)$ satisfy Assumption (H5). Fix $\kappa_{\Psi},\kappa_{h}>0$ and assume that $\Psi_{0}\colon^{d_{u}}\to^{d_{u}}$ and $h_{0}\colon^{d_{u}}\to^{d_{y}}$ are functions satisfying Assumptions (H2) and (H3). Then the following statements hold:
  • There is a constant $C_{p}=C_{p}(\kappa_{\Psi},\Sigma)$ such that for all $\varepsilon\in[0,1]$ and all $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$,
    $\forall\mu\in\mathcal{P}(^{d_{u}}),\qquad d_{g}(\mathsf{P}_{0}\mu,\mathsf{P}\mu)\leqslant C_{p}\varepsilon\cdot\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}$ (C.1)
  • There is a constant $C_{q}=C_{q}(\kappa_{h},\Gamma)$ such that for all $\varepsilon\in[0,1]$ and all $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$,
    $\forall\mu\in\mathcal{P}(^{d_{u}}),\qquad d_{g}(\mathsf{Q}_{0}\mu,\mathsf{Q}\mu)\leqslant C_{q}\varepsilon\cdot\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}$ (C.2)
  • There is $C_{b}=C_{b}(\kappa_{h},\Gamma)$ such that for all $y^{\dagger}\in^{d_{y}}$, all $\varepsilon\in[0,1]$, all $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$, and all probability measures $(\mu_{0},\mu)\in\mathcal{P}(^{d_{u}})\times\mathcal{P}(^{d_{u}})$,
    $d_{g}\Bigl{(}\mathsf{B}(\mathsf{Q}_{0}\mu_{0};y^{\dagger}),\mathsf{B}(\mathsf{Q}\mu,y^{\dagger})\Bigr{)}\leqslant\exp(C_{b}\mathcal{R})\Bigl{(}\varepsilon+d_{g}(\mu_{0},\mu)\Bigr{)},$ (C.3)
    where $\mathcal{R}\in[1,\infty]$ is given by
    $\mathcal{R}=\max\left\{\left\lvert y^{\dagger}\right\rvert^{2},1+\mathcal{M}_{2}(\mu_{0}),1+\mathcal{M}_{2}(\mu)\right\}.$
Here $\mathsf{P}_{0}$ and $\mathsf{Q}_{0}$ denote the maps associated to $(\Psi_{0},h_{0})$, and $\mathsf{P}$ and $\mathsf{Q}$ are the maps associated to $(\Psi,h)$.
Proof.
Throughout this proof, $C_{p}$ denotes a constant depending only on $(\kappa_{\Psi},\Sigma)$, and $C_{q}$ is a constant that depends only on $(\kappa_{h},\Gamma)$. Both may change from line to line.
Proof of (C.1)
It holds that
$\mathsf{P}_{0}\mu(u)-\mathsf{P}\mu(u)=C_{p}\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}u-\Psi_{0}(v)\bigr{\rvert}_{\Sigma}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}u-\Psi(v)\bigr{\rvert}_{\Sigma}^{2}\right)\,\mu(\mathrm{d}v).$
By the elementary inequality (A.9), the integrand on the right-hand side is bounded in absolute value by
$2\bigl{\lvert}\Psi_{0}(v)-\Psi(v)\bigr{\rvert}\left(\exp\left(-\frac{1}{4}\bigl{\lvert}u-\Psi_{0}(v)\bigr{\rvert}_{\Sigma}^{2}\right)+\exp\left(-\frac{1}{4}\bigl{\lvert}u-\Psi(v)\bigr{\rvert}_{\Sigma}^{2}\right)\right).$
By Young’s inequality, it holds that $|a+b|^{2}\geqslant\frac{1}{2}|a|^{2}-|b|^{2}$, and so this is bounded by
$4\bigl{\lvert}\Psi_{0}(v)-\Psi(v)\bigr{\rvert}\left(\exp\left(-\frac{1}{8}\bigl{\lvert}u-\Psi_{0}(v)\bigr{\rvert}_{\Sigma}^{2}+\frac{1}{4}\bigl{\lvert}\Psi_{0}(v)-\Psi(v)\bigr{\rvert}_{\Sigma}^{2}\right)\right)\leqslant C_{p}\varepsilon\exp\left(-\frac{1}{8}\bigl{\lvert}u-\Psi_{0}(v)\bigr{\rvert}_{\Sigma}^{2}\right).$
It follows, by Fubini’s theorem, that
$d_{g}(\mathsf{P}_{0}\mu,\mathsf{P}\mu)$ $\leqslant C_{p}\varepsilon\int_{{}^{d_{u}}}\int_{{}^{d_{u}}}\bigl{(}1+|u|^{2}\bigr{)}\exp\left(-\frac{1}{8}\bigl{\lvert}u-\Psi_{0}(v)\bigr{\rvert}_{\Sigma}^{2}\right)\mathrm{d}u\,\mu(\mathrm{d}v)$
$\leqslant C_{p}\varepsilon\int_{{}^{d_{u}}}\Bigl{(}1+\bigl{\lvert}\Psi_{0}(v)\bigr{\rvert}^{2}\Bigr{)}\,\mu(\mathrm{d}v)\leqslant C_{p}\varepsilon\,(1+\kappa_{\Psi}^{2})\int_{{}^{d_{u}}}\left(1+|v|^{2}\right)\,\mu(\mathrm{d}v).$
This concludes the proof of (C.1).
Proof of (C.2)
Recall that
$d_{g}(\mathsf{Q}_{0}\mu,\mathsf{Q}\mu)=C_{q}\int_{{}^{d_{u}}}\int_{{}^{d_{y}}}g(u,y)\left(\exp\left(-\frac{1}{2}\bigl{\lvert}y-h_{0}(u)\bigr{\rvert}_{\Gamma}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\right)\,\mathrm{d}y\,\mu(\mathrm{d}u),$
where $g(u,y)=1+|u|^{2}+|v|^{2}$. Using the same reasoning as above we obtain that
$\left\lvert\exp\left(-\frac{1}{2}\bigl{\lvert}y-h_{0}(u)\bigr{\rvert}_{\Gamma}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\right\rvert\leqslant C_{q}\varepsilon\exp\left(-\frac{1}{8}\bigl{\lvert}y-h_{0}(u)\bigr{\rvert}_{\Gamma}^{2}\right).$ (C.4)
Therefore, we deduce that
$d_{g}(\mathsf{Q}_{0}\mu,\mathsf{Q}\mu)\leqslant C_{q}\varepsilon\int_{{}^{d_{u}}}\Bigl{(}1+\bigl{\lvert}h_{0}(u)\bigr{\rvert}^{2}\Bigr{)}\,\mu(\mathrm{d}u)\leqslant C_{q}\varepsilon(1+\kappa_{h}^{2})\int_{{}^{d_{u}}}\bigl{(}1+|u|^{2}\bigr{)}\,\mu(\mathrm{d}u),$
which proves (C.2).
Proof of (C.3)
We assume for simplicity that $\mu_{0}$ and $\mu$ have densities, but note that this is not required. Let $\nu_{0}=\mathsf{B}(\mathsf{Q}_{0}\mu_{0};y^{\dagger})$ and $\nu=\mathsf{B}(\mathsf{Q}\mu,y^{\dagger})$ and recall that
$\nu_{0}(u)=\frac{\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h_{0}(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu_{0}(u)}{\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h_{0}(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu_{0}(U)\,\mathrm{d}U}=:\frac{f_{0}(u)}{\int_{{}^{d_{u}}}f_{0}(U)\,\mathrm{d}U}=:\frac{f_{0}(u)}{Z_{0}}$
and similarly
$\nu(u)=\frac{\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u)}{\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(U)\,\mathrm{d}U}=:\frac{f(u)}{\int_{{}^{d_{u}}}f(U)\,\mathrm{d}U}=:\frac{f_{0}(u)}{Z}.$
Note that
$d_{g}(\nu_{0},\nu)$ $=\int_{{}^{d_{u}}}\bigl{(}1+|u|^{2}\bigr{)}\left\lvert\frac{f_{0}(u)}{Z_{0}}-\frac{f(u)}{Z}\right\rvert\,\mathrm{d}u$
$=\frac{1}{Z}\int_{{}^{d_{u}}}\bigl{(}1+|u|^{2}\bigr{)}\bigl{\lvert}f_{0}(u)-f(u)\bigr{\rvert}\,\mathrm{d}u+\left\lvert\frac{1}{Z_{0}}-\frac{1}{Z}\right\rvert\int_{{}^{d_{u}}}\bigl{(}1+|u|^{2}\bigr{)}\,f_{0}(u)\,\mathrm{d}u.$
In order to bound the first term, we write
$f_{0}(u)-f(u)$ $=\left(\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h_{0}(u)\bigr{\rvert}_{\Gamma}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\right)\mu_{0}(u)$
$\qquad+\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\Bigl{(}\mu_{0}(u)-\mu(u)\Bigr{)}.$ (C.5)
Using (C.4), we obtain that
$\bigl{\lvert}f_{0}(u)-f(u)\bigr{\rvert}\leqslant C_{q}\varepsilon\exp\left(-\frac{1}{8}\bigl{\lvert}y^{\dagger}-h_{0}(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu_{0}(u)+\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\Bigl{\lvert}\mu_{0}(u)-\mu(u)\Bigr{\rvert},$ (C.6)
and so
$\int_{{}^{d_{u}}}\bigl{(}1+|u|^{2}\bigr{)}\bigl{\lvert}f_{0}(u)-f(u)\bigr{\rvert}\,\mathrm{d}u$ $\leqslant C_{q}\varepsilon\mathcal{R}+d_{g}(\mu_{0},\mu).$
Therefore it holds that
$d_{g}(\nu_{0},\nu)\leqslant\mathcal{R}\left(\frac{C_{q}\varepsilon}{Z}+\frac{\lvert Z_{0}-Z\rvert}{Z_{0}Z}\right)+\frac{1}{Z}d_{g}(\mu_{0},\mu).$
By (C.6) it holds that
$\lvert Z_{0}-Z\rvert\leqslant\int_{{}^{d_{u}}}\lvert f_{0}(U)-f(U)\rvert\,\mathrm{d}U\leqslant C_{q}\varepsilon+d_{g}(\mu_{0},\mu),$
and so we obtain finally
$d_{g}(\nu_{0},\nu)\leqslant\mathcal{R}\left(\frac{1}{Z}+\frac{1}{Z_{0}Z}\right)\Bigl{(}C_{q}\varepsilon+d_{g}(\mu_{0},\mu)\Bigr{)}.$
Furthermore $Z_{0}$ is bounded from below because, by Jensen’s inequality,
$Z_{0}$ $=\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y^{\dagger}-h_{0}(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu_{0}(U)\,\mathrm{d}U\geqslant\exp\left(-\frac{1}{2}\int_{{}^{d_{u}}}\bigl{\lvert}y^{\dagger}-h_{0}(U)\bigr{\rvert}_{\Gamma}^{2}\,\mu_{0}(U)\,\mathrm{d}U\right)$
$\geqslant\exp\left(-\bigl{\lvert}y^{\dagger}\bigr{\rvert}_{\Gamma}^{2}-C_{q}\kappa_{h}^{2}\int_{{}^{d_{u}}}\bigl{(}1+|u|^{2}\bigr{)}\mu_{0}(U)\,\mathrm{d}U\right)=\exp\left(-C_{q}\mathcal{R}\right).$
The same bound holds for $Z$, and so we obtain the result. ∎
Lemma C.3.
Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices  $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). Fix $\kappa_{\Psi},\kappa_{h}>0$ and assume that $\Psi_{0}\colon^{d_{u}}\to^{d_{u}}$ and $h_{0}\colon^{d_{u}}\to^{d_{y}}$ are functions satisfying Assumptions (H2) and (H3), respectively. Let $(\mu^{0}_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ denote the true filtering distributions associated with functions $(\Psi_{0},h_{0})$ and $(\Psi,h)$, respectively, initialized at the same Gaussian probability measure $\mu_{0}=\mathcal{N}(m_{0},C_{0})\in~{}\mathcal{G}(^{d_{u}})$. For all $J\in\mathbb{Z}^{+}$ there is $C=C(m_{0},C_{0},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J)>0$ such that for all $\varepsilon\in[0,1]$ and all $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$, it holds that
$\max_{j\in\llbracket 0,J\rrbracket}d_{g}(\mu^{0}_{j},\mu_{j})\leqslant C\varepsilon.$ (C.7)
Proof.
By Lemma B.3, the filtering distributions have bounded second moments. Let
[…]
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context, depends only on the constants $m_{0},C_{0},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,k$ (but neither on $\varepsilon$, nor on $\Psi$ and $h$) and may change from line to line.
The statement is obviously true for $J=0$. Reasoning by induction, we assume that the statement is true up to $J=k$ and show that there is $C>0$ such that
$\forall\varepsilon\in[0,1],\qquad\forall(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)},\qquad d_{g}(\mu^{0}_{k+1},\mu_{k+1})\leqslant C\varepsilon.$
To this end, let $\mathsf{P}_{0}$ and $\mathsf{Q}_{0}$ denote the maps associated to $(\Psi_{0},h_{0})$, fix $\varepsilon\in[0,1]$, and fix $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$. Using (C.3), then the triangle inequality, and finally (C.1) and Lemma B.8, we have that
$d_{g}(\mu^{0}_{k+1},\mu_{k+1})$ $=d_{g}\bigl{(}\mathsf{B}_{k}\mathsf{Q}_{0}\mathsf{P}_{0}\mu^{0}_{k},\mathsf{B}_{k}\mathsf{Q}\mathsf{P}\mu_{k}\bigr{)}$
$\leqslant\operatorname{e}^{C\mathcal{R}}\Bigl{(}\varepsilon+d_{g}\bigl{(}\mathsf{P}_{0}\mu^{0}_{k},\mathsf{P}\mu_{k}\bigr{)}\Bigr{)}$
$\leqslant\operatorname{e}^{C\mathcal{R}}\Bigl{(}\varepsilon+d_{g}\bigl{(}\mathsf{P}_{0}\mu^{0}_{k},\mathsf{P}\mu^{0}_{k}\bigr{)}+d_{g}\bigl{(}\mathsf{P}\mu^{0}_{k},\mathsf{P}\mu_{k}\bigr{)}\Bigr{)}$
$\leqslant\operatorname{e}^{C\mathcal{R}}\Bigl{(}\varepsilon+C\varepsilon\mathcal{R}+Cd_{g}(\mu^{0}_{k},\mu_{k})\Bigr{)}.$
Using the induction hypothesis gives us the desired bound. ∎
Appendix D Technical Result for Theorem 2.4
In [le2009large] machinery is established to prove Monte Carlo error estimates between the finite particle ensemble Kalman filter and its mean field limit. We use such results as a component in proving Theorem 2.4 and, in so-doing, explicit dependence on moments must be tracked. In this section we detail proofs of the result in  [le2009large, Proposition 4.4] and [le2009large, Theorem 5.2] , following the analysis closely and, in addition, tracking dependence on moments. This leads to the following error estimate stating the desired Monte Carlo error estimate.
Lemma D.1.
Assume that the probability measures $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{{\rm EK},N}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained, respectively, from the dynamical systems (2.4) and  (2.6), initialized at the Gaussian probability measure $\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$ and at the empirical measure $\mu_{0}^{{\rm EK},N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{u_{0}^{(i)}}$ for $u_{0}^{(i)}\sim\mu^{\rm EK}_{0}$ i.i.d. samples. That is,
$\mu^{\rm EK}_{j+1}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK},\qquad\mu^{{\rm EK},N}_{j+1}=\frac{1}{N}\sum_{i=1}^{N}\delta_{u_{j+1}^{(i)}},$
where $u_{j+1}^{(i)}$ evolve according to the iteration in (2.5). Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). We assume that the vector field $h$ is a linear transformation, i.e. that Assumption (V2) is satisfied. Furthermore, if the vector field $\Psi$ additionally satisfies $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, then for all $\phi$ satisfying Assumption P1, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu^{\rm EK}_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\Sigma,\Gamma\bigr{)}$ where $q\coloneqq\max\{16\cdot 3^{J-1},8(\varsigma+1)3^{J-1}\}$ such that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}}.$
Proof.
To prove the proposition, we apply the coupling argument used in [le2009large]. Using similar notation to the one in [le2009large], to the interacting $N$-particle system $\{u^{(i)}_{j}\}_{n=1}^{N}$ evolving according to the ensemble Kalman dynamics (2.5) , we couple $N$ copies of the mean field dynamics $\{\overline{u}^{(j)}_{j}\}_{n=1}^{N}$ evolving according to the mean field ensemble Kalman dynamics (2.1). The mean field replicas are synchronously coupled to the interacting particle system, in the sense that they are initialized at the same initial condition and driven by the same noises; namely, the two particle systems are initialized at i.i.d. samples $u_{0}^{(i)}=\overline{u}_{0}^{(i)}\sim\mu_{0}^{\rm EK}$ for $i=1,\ldots,N$. For simplicity of notation, the forecast particles are denoted by the letter $v$, and we drop the hat notation from the forecast and simulated observations. Furthermore, we add a bar $\overline{\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}}$ to all the variables related to the synchronously coupled mean field particles, including the probability measures $\overline{\pi}^{\rm EK}_{j}$. We define for $\pi\in\mathcal{P}(^{d_{u}\times d_{y}})$ the Kalman gain
$\mathcal{K}(\pi):=\operatorname{\mathcal{C}}^{uh}\left(\pi\right)\left(\operatorname{\mathcal{C}}^{hh}\left(\pi\right)+\Gamma\right)^{-1}.$
With this notation, the interacting particle system, and synchronously coupled system read as follows:
[…]
Note that the synchronously coupled particles are independent and identically distributed. With this set-up we proceed with the proof. We first define
$D_{j}^{N,p}\coloneqq\Bigl{(}\mathbb{E}\left\lvert u^{(i)}_{j}-\overline{u}^{(i)}_{j}\right\rvert^{p}\Bigr{)}^{1/p}.$
We now let $\phi$ satisfy Assumption P1. By applying the triangle inequality, we deduce that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{1/2}+\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}.$ (D.3)
The proof of [le2009large, Theorem 5.2] uses the law of large numbers in $L^{2}$ to establish the bound for the second term of the right hand side of D.3
$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{J}^{\rm EK}\bigr{)},R_{\phi}\Bigr{)}}{\sqrt{N}};$ (D.4)
furthermore, in the proof of [le2009large, Theorem 5.2], the following bound on the first term of the right hand side of D.3 may be deduced:
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{1/2}\leqslant L_{\phi}D_{J}^{N,4}\cdot\Bigl{(}1+\bigl{|}\mathcal{M}^{1/4}_{4\varsigma}\bigl{(}\mu_{J}^{\rm EK}\bigr{)}\bigr{|}\Bigr{)}+L_{\phi}|D_{J}^{N,2(\varsigma+1)}|^{\varsigma+1}.$ (D.5)
Under the assumptions of the proposition, the result of [le2009large, Proposition 4.4] shows by induction that for any $j\in\mathbb{N}$, it holds that  [^1] The result of [le2009large, Proposition 4.4] holds more generally for functions $\Psi$ that are locally Lipschitz and that grow at most polynomially at infinity [le2009large, Assumption B]. However, to obtain (D.6) from the result of [le2009large, Proposition 4.4] we use that $\Psi$ is assumed to be globally Lipschitz.
$D_{j}^{N,p}\leqslant CD_{j-1}^{N,p}$ $+CD_{j-1}^{N,2\vee(pr)}\Bigl{(}D_{j-1}^{N,2\vee(pr^{\prime})}+\bigl{(}\mathcal{M}\bigl{(}\mathsf{P}\mu_{j-1}^{\rm EK}\bigr{)}\bigr{)}^{2\vee(pr^{\prime})}\Bigr{)}\cdot\Bigl{(}D_{j-1}^{N,2\vee(pr^{\prime\prime})}+\overline{S}_{j}^{pr^{\prime\prime}}\Bigr{)}$ (D.6)
$\quad+C\Bigl{(}\mathbb{E}\bigl{|}Z_{j}^{N}\bigr{|}^{pq}\Bigr{)}^{1/pq}\cdot\bigl{(}D_{j-1}^{N,pq^{\prime}}+\overline{S}_{j}^{pq^{\prime}}\bigr{)},$
for any $j\in\mathbb{N}$, where $C$ is a constant depending on $\Sigma,\Gamma,\ell_{\Psi},\kappa_{\Psi},\kappa_{h},\mathcal{M}_{2}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)}$. We note that the specific form of $C$ may be deduced using the proof of [le2009large, Lemma 4.1]. Furthermore, we note that
$\frac{1}{q}+\frac{1}{q^{\prime}}=1,\qquad\frac{1}{r}+\frac{1}{r^{\prime}}+\frac{1}{r^{\prime\prime}}=1,$
and that $\overline{S}_{j}^{t}$ and $Z_{j}^{N}$ are defined as
$\overline{S}_{j}^{t}$ $\coloneqq\Bigl{(}\mathbb{E}\bigl{|}{y}^{\dagger}_{j}-h\Bigl{(}\overline{v}^{(i)}_{j}\Bigr{)}-\eta^{(i)}_{j}\bigr{|}^{t}\Bigr{)}^{1/t},$
$Z_{j}^{N}$ $\coloneqq\left\|\mathcal{C}\Bigl{(}\frac{1}{N}\sum_{i=1}^{N}\delta_{\overline{v}^{(i)}_{j}}\Bigr{)}-\mathcal{C}\bigl{(}\mathsf{P}\mu_{j-1}^{\rm EK}\bigr{)}\right\|.$
By applying the triangle inequality it is straightforward to deduce that
$\overline{S}_{j}^{t}\leqslant\left(\int_{{}^{d_{u}}}\bigl{|}{y}^{\dagger}_{j}-h(u)\bigr{|}^{t}\mathsf{P}\mu_{j-1}^{\rm EK}(\mathrm{d}u)\right)^{1/t}+\bigl{(}\mathbb{E}|\eta_{j}|^{t}\bigr{)}^{1/t}\leqslant C\Bigl{(}\mathcal{M}_{t}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)},\kappa_{\Psi},\kappa_{h},\kappa_{y},\Sigma,\Gamma\Bigr{)},$ (D.7)
for any $t\in\mathbb{N}$. On the other hand, [le2009large, Lemma 3.3] shows that
$\Bigl{(}\mathbb{E}\bigl{|}Z_{j}^{N}\bigr{|}^{t}\Bigr{)}^{1/t}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2t}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)},\kappa_{\Psi},\Sigma\Bigr{)}}{\sqrt{N}},$ (D.8)
for any $N\in\mathbb{N}$ and any $t\in\mathbb{N}$. Combining the results from (D.7) and (D.8) with the one in (D.6), and choosing for convenience $q=q^{\prime}=2$ and $r=r^{\prime}=r^{\prime\prime}=3$ for any $j\in\mathbb{N}$ we obtain
$D_{j}^{N,p}\leqslant CD_{j-1}^{N,p}+CD_{j-1}^{N,3p}\Bigl{(}D_{j-1}^{N,3p}+1\Bigr{)}\cdot\Bigl{(}D_{j-1}^{N,3p}+1\Bigr{)}+\frac{C}{\sqrt{N}}\cdot\bigl{(}D_{j-1}^{N,2p}+1\bigr{)},$ (D.9)
for a constant $C$ depending on $\mathcal{M}_{4p}{\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)}},\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$. Since $D_{0}^{N,p}=0$ for any $p\in\mathbb{N}$, by iterating in $j$ the result in (D.9), we find that
$D_{J}^{N,4}\leqslant\frac{C}{\sqrt{N}},$
where $C$ is a constant depending on $\mathcal{M}_{16\cdot 3^{J-1}}{\bigl{(}\mu_{0}^{\rm EK}\bigr{)}},J,\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$. Hence, by combining the results of (D.5) and (D.4) with the decomposition (D.3), we deduce that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}},$
where $C$ is a constant depending on $\mathcal{M}_{q}{\bigl{(}\mu_{0}^{\rm EK}\bigr{)}},J,R_{\phi},L_{\phi},\varsigma,\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$ for a constant $q$ defined as $q\coloneqq\max\{16\cdot 3^{J-1},8(\varsigma+1)3^{J-1}\}$. ∎
[…]

### v2

@@
 There has been recent progress in the theoretical analysis of the algorithm, establishing stability and error estimates in non-Gaussian settings, but the assumptions on the dynamics and observation models rule out the unbounded vector fields that arise in practice and the analysis applies only to the mean field limit of the ensemble Kalman filter.
-The present work establishes error bounds between the filtering distribution and the finite particle ensemble Kalman filter when the model exhibits linear growth.
+The present work establishes error bounds between the filtering distribution and the finite particle ensemble Kalman filter when the dynamics and observation vector fields may be unbounded, allowing linear growth.
 Keywords. Ensemble Kalman filter, Stochastic filtering, Weighted total variation metric, Stability estimates, Accuracy estimates, Near-Linear setting.
@@
 Particle filters or sequential Monte Carlo methods [doucet2001sequential, chopin:20] offer an alternative methodology for nonlinear filtering problems, allowing recovery of the exact filtering distribution in the large particle limit.
-However, this class of methods scales poorly with dimension, and in particular suffer weight collapse, making their application to geophysical problems prohibitive; see for example [snyder2008obstacles, bickel2008sharp, rebeschini2015can, agapiou2017importance].
+However, this class of methods scales poorly with dimension, and in particular suffers weight collapse, making its application to geophysical problems prohibitive; see for example [snyder2008obstacles, bickel2008sharp, rebeschini2015can, agapiou2017importance].
 The ensemble Kalman filter was introduced in the seminal paper [evensen1994sequential].
-Its success stems from the low-rank approximation of large covariances, cheaply computed using an ensemble of particles and allowing it to be deployed in geophysical applications.
+Its success stems from the low-rank approximation of large covariances, cheaply computed using an ensemble of particles, which allows the methodology to be deployed in geophysical applications.
 However analysis of its accuracy, in relation to the true filtering distribution, remains in its infancy.
@@
 In the recent paper [carrillo2022ensemble] the authors establish stability properties of the mean field ensemble Kalman filter and use them to prove accuracy of the filter in a near-Gaussian setting.
-However the paper does not consider finite particle approximations of the mean field, and the conditions on the dynamics-observation model require boundedness of the vector fields arising.
+However, the paper does not consider finite particle approximations of the mean field, and the conditions on the dynamics-observation model require boundedness of the vector fields arising.
 In this paper we address both these issues, establishing error bounds between the finite particle ensemble Kalman filter and the true filtering distribution in settings where the dynamics and observation vector fields grow linearly.
@@
     Theorem 2.4 quantifies the error between the finite particle ensemble Kalman filter (found as an interacting particle system approximation of the mean field) and the true filter in the setting of near-linear, Lipschitz dynamical models and linear observation operators.
-In going beyond the work in [carrillo2022ensemble], the current paper simultaneously addresses a more applicable problem class, by allowing linear growth in the dynamics and observations operators, and confronts the substantial technical challenges which arise from doing so.
+In going beyond the work in [carrillo2022ensemble], the current paper simultaneously addresses a more applicable problem class, by allowing linear growth in the dynamics and observation operators, and confronts the substantial technical challenges which arise from doing so.
 Bounds on moments of the filtering distribution and mean field ensemble Kalman filter must be established; these bounds grow in the number of iterations which is in contrast to [carrillo2022ensemble] where the $L^{\infty}$ bounds on model and observation operator ensure a uniform bound on moments.
@@
 $\|\mathfrak{f}-f\|_{L^{\infty}(^{n})}\leqslant r,\quad\|\mathfrak{g}-g\|_{L^{\infty}(^{d})}\leqslant r.$
-We let $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ denote the $C^{0,1}$ semi-norm, namely the Lipschitz constant.
+We write $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ for the $C^{0,1}$ semi-norm, referring in particular to the Lipschitz constant.
 […]
@@
 $\mathcal{M}_{q}(\mu)\coloneqq\int_{{}^{n}}|x|^{q}\mu(\mathrm{d}x).$ (1.1)
-The notations $\mathcal{M}(\mu)$ and $\mathcal{C}(\mu)$ denote respectively the mean and covariance under $\mu$:
+We denote by $\mathcal{M}(\mu)$ and $\mathcal{C}(\mu)$ the mean and covariance under $\mu$, respectively:
 $\mathcal{M}(\mu)=\mu[x],\qquad\mathcal{C}(\mu)=\mu\Bigl{[}\bigl{(}x-\mathcal{M}(\mu)\bigr{)}\otimes\bigl{(}x-\mathcal{M}(\mu)\bigr{)}\Bigr{]}.$
-The notation $\mathcal{P}_{R}(^{n})$ for $R\geqslant 1$ refers to the subset of $\mathcal{P}(^{n})$ of probability measures whose mean and covariance satisfy the bound
-$\lvert\mathcal{M}(\mu)\rvert\leqslant R,\qquad\frac{1}{R^{2}}I_{n}\preccurlyeq\mathcal{C}(\mu)\preccurlyeq R^{2}I_{n}.$ (1.2)
+We use $\mathcal{P}_{R}(^{n})$ with $R\geqslant 1$ to denote the subset of $\mathcal{P}(^{n})$ of probability measures with mean and covariance satisfying the bounds
+$\lvert\mathcal{M}(\mu)\rvert\leqslant R,\qquad\frac{1}{R^{2}}I_{n}\preccurlyeq\mathcal{C}(\mu)\preccurlyeq R^{2}I_{n},$ (1.2)
 […]
-For a probability measure $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ associated with random variable $(u,y)\in^{d_{u}}\times^{d_{y}}$ we use the notation $\mathcal{M}^{u}(\pi)$, $\mathcal{M}^{y}(\pi)$ for the means of the marginal distributions, and the notation $\mathcal{C}^{uu}(\pi)$, $\mathcal{C}^{uy}(\pi)$ and $\mathcal{C}^{yy}(\pi)$ for the blocks of the covariance matrix $\mathcal{C}(\pi)$.
-That is to say,
+For $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ defined on the joint space of state and data associated with random variable $(u,y)\in^{d_{u}}\times^{d_{y}}$ we denote by $\mathcal{M}^{u}(\pi)$, $\mathcal{M}^{y}(\pi)$ the means of the marginal distributions, and by $\mathcal{C}^{uu}(\pi)$, $\mathcal{C}^{uy}(\pi)$ and $\mathcal{C}^{yy}(\pi)$ the blocks of the covariance matrix $\mathcal{C}(\pi)$.
+In particular, we write
 $\mathcal{M}(\pi)=\begin{pmatrix}\mathcal{M}^{u}(\pi)\\ \mathcal{M}^{y}(\pi)\end{pmatrix},\qquad\mathcal{C}(\pi)=\begin{pmatrix}\mathcal{C}^{uu}(\pi)&\mathcal{C}^{uy}(\pi)\\ \mathcal{C}^{uy}(\pi)^{\mathsf{T}}&\mathcal{C}^{yy}(\pi)\end{pmatrix}.$ (1.3)
 For $h\colon^{d_{u}}\to^{d_{y}}$ we also define ${\mathcal{C}}^{hh}(\pi)$ to be the covariance of vector $h(u)$, for $u$ distributed according to the marginal of $\pi$ on $u$, and ${\mathcal{C}}^{uh}(\pi)$ to be the covariance between $u$ and $h(u)$.
-We also introduce the Gaussian projection operator $\mathsf{G}\colon\mathcal{P}(^{n})\to\mathcal{G}(^{n})$ given by $\mathsf{G}\mu=\mathcal{N}\bigl{(}\mathcal{M}(\mu),\mathcal{C}(\mu)\bigr{)}.$ We refer to $\mathsf{G}$ as a projection because $\mathsf{G}\mu$ is the Gaussian distribution closest to $\mu$ with respect to $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ [MR2247587], where $\mathrm{KL}(\mu\|\nu)$ is the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$.
-Note that $\mathsf{G}$ defines a nonlinear mapping.
-Throughout this work, we employ the following weighted total variation distance, shown to be useful in the analysis of ensemble Kalman methods in [carrillo2022ensemble]:
+Given this notation, we also introduce the set $\mathcal{P}^{2}_{\succ 0}(^{d_{u}\times d_{y}})$ of probability measures $\pi$ with finite second moment satisfying $\mathcal{C}^{yy}(\pi)\succ 0$.
+Throughout, we will denote operators acting on the space of probability measures via the $\mathsf{mathsf}$ font.
+We introduce the Gaussian projection operator $\mathsf{G}\colon\mathcal{P}(^{n})\to\mathcal{G}(^{n})$ defined by $\mathsf{G}\mu=\mathcal{N}\bigl{(}\mathcal{M}(\mu),\mathcal{C}(\mu)\bigr{)}.$ We refer to $\mathsf{G}$ as a projection because $\mathsf{G}\mu$ is the Gaussian distribution closest to $\mu$ with respect to $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ [MR2247587], where $\mathrm{KL}(\mu\|\nu)$ denotes the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$.
+We note that $\mathsf{G}$ defines a nonlinear mapping.
+Throughout the paper we make use of the following weighted total variation distance as in [carrillo2022ensemble]:
 Definition 1.1.
-Let $g\colon^{n}\to[1,\infty)$ denote $g(v):=1+\lvert v\rvert^{2}$. We define the weighted total variation metric $d_{g}\colon\mathcal{P}(^{n})\times\mathcal{P}(^{n})\to$
-$d_{g}(\mu_{1},\mu_{2})=\sup_{\lvert f\rvert\leqslant g}\bigl{\lvert}\mu_{1}[f]-\mu_{2}[f]\bigr{\rvert},$
-where the supremum is over all functions $f\colon^{n}\to$ which are bounded from above by $g$ pointwise and in absolute value.
+Let $g\colon^{n}\to[1,\infty)$ define the function $g(v):=1+\lvert v\rvert^{2}$. The weighted total variation metric $d_{g}\colon\mathcal{P}(^{n})\times\mathcal{P}(^{n})\to$ is defined by
+$d_{g}(\mu_{1},\mu_{2})=\sup_{|f|\leqslant g}\bigl{|}\mu_{1}[f]-\mu_{2}[f]\bigr{|},$
+where the supremum is taken over all functions $f\colon^{n}\to$ bounded from above by $g$ pointwise and in absolute value.
 Remark 1.2.
@@
   • Unlike the usual total variation distance, the weighted metric in Definition 1.1 enables control of the differences $\lvert\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\rvert$ and $\lVert\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\rVert$.
-This is the content of Lemma B.6, proved in the appendix, which is used in the proof of a key auxiliary result (Lemma B.2).
+This is the content of Lemma B.6, stated in Subsection B.1, which is used in the proof of a key auxiliary result (Lemma B.11).
 1.3 Filtering Distribution
-We consider the following stochastic dynamics and data model
+Here we introduce the hidden Markov model that gives rise to the filtering distribution.
+We employ notation similar to that established in [calvello2025ensemble, carrillo2022ensemble].
+We consider $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}\subset^{d_{u}}$ to be unknown states to be determined from associated observations $\{y_{j}\}_{j\in\llbracket 1,J\rrbracket}\subset^{d_{y}}$.
+We assume the states and observations to be governed by the following stochastic dynamics and data model:
 $u_{j+1}$ $=\Psi(u_{j})+\xi_{j},\qquad$ $\xi_{j}\sim\mathcal{N}(0,\Sigma),$ (1.4a)
 $y_{j+1}$ $=h(u_{j+1})+\eta_{j+1},\qquad$ $\eta_{j+1}\sim\mathcal{N}(0,\Gamma).$ (1.4b)
-Here $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}$ is the unknown state, evolving in ${}^{d_{u}}$, and $\{y_{j}\}_{j\in\llbracket 1,J\rrbracket}$ are the observations, evolving in ${}^{d_{y}}$.
-We assume that the initial state is distributed according to a Gaussian distribution $u_{0}\sim\mathcal{N}(m_{0},C_{0})$ and that the following independence assumption is satisfied:
+We assume that the initial state is a Gaussian random variable $u_{0}\sim\mathcal{N}(m_{0},C_{0})$ and that the following independence condition is satisfied:
 $u_{0}\perp\!\!\!\!\perp\xi_{0}\perp\!\!\!\!\perp\dotsb\perp\!\!\!\!\perp\xi_{J-1}\perp\!\!\!\!\perp\eta_{1}\perp\!\!\!\!\perp\dotsb\perp\!\!\!\!\perp\eta_{J}.$ (1.5)
-The filtering distribution $\mu_{j}$ is the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}:=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ of the data process up to step $j$.
-Data $Y_{j}^{\dagger}$ may be thought of as arising from a realization of (1.4); but the case of model misspecification, where $Y_{j}^{\dagger}$ does not necessarily arise from (1.4), is also of interest.
-It is well-known [MR3363508, reich2015probabilistic] that the true filtering distribution evolves according to
-$\mu_{j+1}=\mathsf{L}_{j}\mathsf{P}\mu_{j}.$ (1.6)
-where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps on probability measures, referred to respectively as the prediction and analysis steps in the data assimilation community [asch2016data].
-The prediction operator $\mathsf{P}$ is linear and defined by the Markov kernel associated with the stochastic dynamics (1.4a).
-Its action on a probability measure with Lebesgue density $\mu$ reads
-$\mathsf{P}\mu(u)=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\mu(v)\,\mathrm{d}v.$ (1.7)
-The operator $\mathsf{L}_{j}$ is a nonlinear map which formalizes the incorporation of the new datum $y^{\dagger}_{j+1}$ using Bayes’ theorem. Its action reads
-$\mathsf{L}_{j}\mu(u)=\frac{\displaystyle\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u)}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(U)\,\mathrm{d}U}$ (1.8)
-The operator $\mathsf{L}_{j}$ effects a reweighting of the measure to which it applies, with more weight assigned to the state values that are consistent with the observation.
-It is convenient in this work to decompose the analysis map $\mathsf{L}_{j}$ into the composition $\mathsf{B}_{j}\mathsf{Q}$, where the operators $\mathsf{Q}\colon\mathcal{P}(^{d_{u}})\to\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and $\mathsf{B}_{j}\colon\mathcal{P}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ are defined by
-$\mathsf{Q}\mu(u,y)=\frac{1}{\sqrt{(2\pi)^{d_{y}}\det\Gamma}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u).$ (1.9)
-and
-$\mathsf{B}_{j}\mu(u)=\frac{\mu(u,y_{j+1}^{\dagger})}{\displaystyle\int_{{}^{d_{u}}}\mu(U,y_{j+1}^{\dagger})\,\mathrm{d}U}.$ (1.10)
-The operator $\mathsf{Q}$ maps a probability measure with density $\mu$ into the density associated with the joint random variable $\bigl{(}U,h(U)+\eta\bigr{)}$, where $U\sim\mu$ is independent of $\eta\sim\mathcal{N}(0,\Gamma)$.
-The operator $\mathsf{B}_{j}$ performs conditioning on the data $y^{\dagger}_{j+1}$.
-Map $\mathsf{Q}$ is linear whilst $\mathsf{B}_{j}$ is nonlinear.
-We may thus write (1.6) in the form
+We define the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}:=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ as the filtering distribution, which we denote by $\mu_{j}$.
+We also refer to $\mu_{j}$ as the true filter.
+In the context of this paper, data $Y_{j}^{\dagger}$ is assumed to be a collection of realizations from (1.4).
+As formulated in [MR3363508, reich2015probabilistic], the evolution of the filtering distribution may be written as
+$\mu_{j+1}=\mathsf{L}_{j}\mathsf{P}\mu_{j},$ (1.6)
+where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps from $\mathcal{P}(^{d_{u}})$ into itself and which effect what are referred to in the data assimilation community as the prediction and analysis steps, respectively [asch2016data].
+The prediction operator $\mathsf{P}$ is linear and is determined by the Markov kernel arising from (1.4a).
+On the other hand, the operator $\mathsf{L}_{j}$ is nonlinear and encodes the incorporation of the new data point $y^{\dagger}_{j+1}$ using Bayes’ theorem.
+These operators are defined via action on probability measures with Lebesgue density $\mu$ as
+$\mathsf{P}\mu(u)$ $=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\mu(v)\,\mathrm{d}v,$ (1.7a)
+$\mathsf{L}_{j}\mu(u)$ $=\frac{\displaystyle\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u)}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(U)\,\mathrm{d}U}$ (1.7b)
+We may rewrite the analysis map $\mathsf{L}_{j}$ as the composition $\mathsf{B}_{j}\mathsf{Q}$, where the operators $\mathsf{Q}\colon\mathcal{P}(^{d_{u}})\to\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and $\mathsf{B}_{j}\colon\mathcal{P}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ are defined by
+$\mathsf{Q}\mu(u,y)$ $=\frac{1}{\sqrt{(2\pi)^{d_{y}}\det\Gamma}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u),$ (1.8a)
+$\mathsf{B}_{j}\mu(u)$ $=\frac{\mu(u,y_{j+1}^{\dagger})}{\displaystyle\int_{{}^{d_{u}}}\mu(U,y_{j+1}^{\dagger})\,\mathrm{d}U}.$ (1.8b)
+Linear operator $\mathsf{Q}$ maps a probability measure with density $\mu$ into the law of random variable $\bigl{(}U,h(U)+\eta\bigr{)}$, with $U\sim\mu$ independent of $\eta\sim\mathcal{N}(0,\Gamma)$.
+Nonlinear operator $\mathsf{B}_{j}$ effects conditioning on the datum $y^{\dagger}_{j+1}$.
+Hence, we may reformulate (1.6) as
 $\mu_{j+1}=\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}.$ (1.9)
 2 The Ensemble Kalman Filter
-In Subsection 2.1 we define the specific version of the mean field ensemble Kalman filter that we analyze here; other versions may be found in [calvello2025ensemble] and will be amenable to similar analyses.
-In Subsection 2.2 we prove our main stability theorem, showing that the error between the true filter and its Gaussian projection, on the joint space of state and observations, may be used to control the error between the true filter and the mean field ensemble Kalman filter.
-In Subsection 2.3 we prove a corollary to this theorem, establishing that the mean field ensemble Kalman filter accurately approximates the true filter for a specific class of non-Gaussian problems.
+We begin in Subsection 2.1 by introducing the specific form of the mean field ensemble Kalman filter that we consider in this paper; other versions leading to implementable algorithms and for which similar analysis may be developed can be found in [calvello2025ensemble].
+In this subsection we also outline the various assumptions that will be employed for the results of the paper.
+We note that the subsections that follow proceed with increasingly restrictive assumptions on dynamics and observation operator.
+However, all of the theorems allow for linear growth of the vector fields defining the dynamics and the observation processes.
+In Subsection 2.2 we state and prove a stability theorem, which shows that the error between the true filter and the mean field ensemble Kalman filter may be controlled by the error between the true filter and its Gaussian projection on the joint space of state and observations.
+In Subsection 2.3 we establish an error estimate which shows that the mean field ensemble Kalman filter provides an accurate approximation of the true filtering distribution for dynamics and observation operators that are close-to-linear.
 In Subsection 2.4 we deploy the results of the two preceding subsections to state a theorem quantifying the error between the finite particle ensemble Kalman filter itself and the true filtering distribution.
-All of the theorems allow for linear growth of the vector fields defining the dynamics and the observation processes.
 2.1 The Algorithm
-The ensemble Kalman filter may be derived as a particle approximation of various mean field dynamics [calvello2025ensemble].
-The specific mean field ensemble Kalman filter that we study in this paper reads
+The ensemble Kalman filter as implemented in practice may be derived as a particle approximation of various mean field dynamics [calvello2025ensemble].
+The particular mean field ensemble Kalman filter that we consider in this paper may be written as
 $\widehat{u}_{j+1}$ $=\Psi(u_{j})+\xi_{j},\qquad$ $\xi_{j}\sim\mathcal{N}(0,\Sigma),$ (2.1a)
@@
 $u_{j+1}$ $=\widehat{u}_{j+1}+\mathcal{C}^{uy}\left(\widehat{\pi}_{j+1}^{{\rm EK}}\right)\mathcal{C}^{yy}\left(\widehat{\pi}_{j+1}^{{\rm EK}}\right)^{-1}\bigl{(}y^{\dagger}_{j+1}-\widehat{y}_{j+1}\bigr{)},$ (2.1c)
-where $\widehat{\pi}^{\rm EK}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and independence of the noise terms (1.5) is still assumed to hold, as in (1.5).
-See Subsection 1.2 for the definition of the covariance matrices that appear in (2.1c).
+where $\widehat{\pi}^{{\rm EK}}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and where the independence condition (1.5) holds.
+We refer the reader back to Subsection 1.2 for the definition of the covariance matrices appearing in (2.1c).
 We denote by $\mu^{{\rm EK}}_{j}$ the law of $u_{j}$.
-In order to rewrite the evolution of $\mu_{j}^{\rm EK}$ in terms of maps on probability measures we introduce, for a given $y^{\dagger}_{j+1}$, the map $\mathsf{T}_{j}\colon\mathcal{P}^{2}_{\succ 0}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ defined by
+We aim to formulate the evolution of $\mu_{j}^{{\rm EK}}$ in terms of maps on probability measures, hence we introduce, for a given $y^{\dagger}_{j+1}$, the map $\mathsf{T}_{j}\colon\mathcal{P}^{2}_{\succ 0}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ defined by
 $\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi.$ (2.2)
-Here the subscript denotes the pushforward and, for given $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$, $z\in^{d_{y}}$,
-$\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,z)\colon$${}^{d_{u}}\times^{d_{y}}\to^{d_{u}};$
-$(u,y)\mapsto$ $\,\,u+\mathcal{C}^{uy}(\pi)\mathcal{C}^{yy}(\pi)^{-1}\bigl{(}z-y\bigr{)}.$ (2.3)
-Evolution of the probability measure $\mu_{j}^{\rm EK}$ may then be written as [calvello2025ensemble]
+Here for given $\pi\in{\mathcal{P}^{2}_{\succ 0}}(^{d_{u}}\times^{d_{y}})$, $z\in^{d_{y}}$, the transport map $\mathscr{T}$ is defined as
+$\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,z)\colon$${}^{d_{u}}\times^{d_{y}}\to^{d_{u}};$ (2.3a)
+$(u,y)\mapsto$ $\,\,u+\mathcal{C}^{uy}(\pi)\mathcal{C}^{yy}(\pi)^{-1}\bigl{(}z-y\bigr{)}.$ (2.3b)
+Evolution of the probability measure $\mu_{j}^{{\rm EK}}$ may then be written (see [calvello2025ensemble]) as
 $\mu_{j+1}^{{\rm EK}}=\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}}.$ (2.4)
-We now discuss the preceding map in relation to (1.11).
-The specific affine map $\mathscr{T}$ used in (2.1c) is determined by measure $\widehat{\pi}^{\rm EK}_{j+1}$ (here equal to $\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$) and data $y^{\dagger}_{j+1}$.
-The following observations explain why the law of $u_{j}$ in (2.1) evolves according to (2.4):
-  • if $u_{j}\sim\mu_{j}^{\rm EK}$, then $\widehat{u}_{j+1}\sim\mathsf{P}\mu_{j}^{\rm EK}$, by definition of $\mathsf{P}$;
-  • random vector $(\widehat{u}_{j+1},\widehat{y}_{j+1})$ is distributed according to $\widehat{\pi}^{\rm EK}_{j+1}=\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$;
-  • equation (2.1c) then implies that $u_{j+1}\sim\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$.
-As we show in Lemma B.7, operator $\mathsf{T}_{j}$ coincides with conditioning operator $\mathsf{B}_{j}$ over the Gaussians $\mathcal{G}(^{d_{u}}\times^{d_{y}})\subset\mathcal{P}(^{d_{u}}\times^{d_{y}}).$ In the particular case where $\mu_{0}$ is Gaussian, which is a standing assumption in this paper, and the operators $\Psi$ and $h$ are linear, the mean field ensemble Kalman filter (2.4) thus reproduces the exact filtering distribution (1.11), given by the Kalman filter itself.
-The aim of this paper is to analyze the accuracy of ensemble Kalman methods when $\Psi$ and $h$ are not assumed to be linear.
+We note that a specific instance of map $\mathscr{T}$ defined in (2.3) appears in (2.1c) and is determined by probability measure $\pi:=\widehat{\pi}^{{\rm EK}}_{j+1}$ (here equal to $\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}}$) and data $z:=y^{\dagger}_{j+1}$.
+We refer to [carrillo2022ensemble] for a step-by-step argument detailing why the law of $u_{j}$ in (2.1) evolves according to (2.4).
+The operator $\mathsf{T}_{j}$ is equivalent, over the Gaussians $\mathcal{G}(^{d_{u}}\times^{d_{y}})\subset\mathcal{P}(^{d_{u}}\times^{d_{y}}),$ to the conditioning operator $\mathsf{B}_{j}$; see Lemma B.7.
+We recall that in the particular case where $\mu_{0}$ is Gaussian, which we assume throughout the paper, and the operators $\Psi$ and $h$ are linear, the mean field ensemble Kalman filter (2.4) coincides with the filtering distribution (1.9).
+However, this paper focuses on the aim of analyzing the accuracy of ensemble Kalman methods when $\Psi$ and $h$ are not assumed to be linear.
 The ensemble Kalman filter as implemented in practice may be derived as a particle approximation of the mean field dynamics as defined by sample paths in (2.1) or as an evolution on measures in (2.4).
-For any $\pi$ in the range of $\mathsf{Q}$, $\mathcal{C}^{yy}(\pi)=\mathcal{C}^{hh}(\pi)+\Gamma$ and $\mathcal{C}^{uy}(\pi)=\mathcal{C}^{uh}(\pi)$, using the notation introduced in, and following, (1.3).
-This is since the noise in the observation component defining $\pi$ is then independent of the state component defining $\pi.$ Using this the particle approximation of (2.1) takes the form
+To this end, we observe that for any $\pi$ in the range of $\mathsf{Q}$, $\mathcal{C}^{yy}(\pi)=\mathcal{C}^{hh}(\pi)+\Gamma$ and $\mathcal{C}^{uy}(\pi)=\mathcal{C}^{uh}(\pi)$, using the notation introduced in, and following, (1.3).
+This is since the noise in the observation component defining $\pi$ is then independent of the state component defining $\pi.$ Using this observation, the particle approximation of (2.1) takes the form
 $\widehat{u}^{(i)}_{j+1}$ $=\Psi\bigl{(}u^{(i)}_{j}\bigr{)}+\xi^{(i)}_{j},\qquad$ $\xi^{(i)}_{j}\sim\mathcal{N}(0,\Sigma),$ (2.5a)
@@
 whose evolution describes the ensemble Kalman filter.
-Our theorems in Subsections 2.2 and 2.3 concern relationships between the true filter (1.11) and the mean field ensemble Kalman filter (2.4).
-We study the setting in which $\Psi$ and $h$ are not assumed to be linear, so that the true filter is not Gaussian, in Subsection 2.2; and then we study small perturbations away from the Gaussian setting that arise when the vector fields $\Psi$ and $h$ are close to affine in Subsection 2.3.
+Our theorems in Subsections 2.2 and 2.3 regard the relationship between the true filter (1.9) and the mean field ensemble Kalman filter (2.4).
+In Subsection 2.2 we consider the setting in which $\Psi$ and $h$ exhibit linear growth but are not assumed to be linear, hence the true filter is not Gaussian; in Subsection 2.3 we then consider small perturbations of the Gaussian setting that arise when the vector fields $\Psi$ and $h$ are close to affine.
 The theorem in Subsection 2.4 concerns the relationship between the true filter (1.9) and the ensemble Kalman filter (2.6).
-In this subsection, we combine existing analysis on the convergence of the ensemble Kalman filter to the mean field ensemble Kalman filter with the results from the previous subsections to state and prove an error estimate between the ensemble Kalman filter and the true filter in the non-linear setting.
+In this subsection, we combine existing analysis on the convergence of the ensemble Kalman filter to the mean field ensemble Kalman filter with the results from the previous subsections to state and prove an error estimate between the ensemble Kalman filter and the true filter in the nonlinear setting.
 Specifically, we study the case of a vector field $\Psi$ that is a bounded perturbation away from affine and an affine vector field $h$.
-To state our theorems we will use the following set of assumptions:
+To state our theorems we will use the following set of assumptions.
 Assumption H.
-There exists positive constants $\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\sigma$ and $\gamma$ such that the data $\{y_{j}^{\dagger}\}$, the vector fields $(\Psi,h)$ and the covariances $(\Sigma,\Gamma)$ satisfy:
+There exist positive constants $\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\sigma$ and $\gamma$ such that the data $\{y_{j}^{\dagger}\}$, the vector fields $(\Psi,h)$ and the covariances $(\Sigma,\Gamma)$ satisfy:
   1. (H1)
@@
     $B_{y}:=\left\{Y^{\dagger}\in^{KJ}:\max_{j\in\llbracket 1,J\rrbracket}\lvert y^{\dagger}_{j}\rvert\leqslant\kappa_{y}\right\};$
+  2. (H2)
+    covariance matrices $\Sigma$ and $\Gamma$ are positive definite: $\Sigma\succcurlyeq\sigma I_{d_{u}}$ and $\Gamma\succcurlyeq\gamma I_{d_{y}}$ for positive $\sigma$ and $\gamma$.
   3. (H3)
@@
     function $h\colon^{d_{u}}\to^{d_{y}}$ satisfies $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$;
-  5. (H5)
-    covariance matrices $\Sigma$ and $\Gamma$ are positive definite: $\Sigma\succcurlyeq\sigma I_{d_{u}}$ and $\Gamma\succcurlyeq\gamma I_{d_{y}}$ for positive $\sigma$ and $\gamma$.
 Assumption V.
@@
 2.2 Stability Theorem: Mean Field Ensemble Kalman Filter
-Roughly speaking, our stability theorem states that if the true filtering distributions $(\mu_{j})_{j\in\llbracket 0,J-1\rrbracket}$ are close to Gaussian, after appropriate lifting to the state/data space, then the distribution $\mu^{\rm EK}_{j}$ given by the mean field ensemble Kalman filter (2.4) is close to the true filtering distribution $\mu_{j}$ given by (1.11) for all $j\in\llbracket 0,J\rrbracket$.
+In this section we consider the setting of dynamics and observation operator satisfying the linear growth assumptions contained in Assumption H.
+Informally the result of Theorem 2.1 shows that, under these assumptions, and if the true filtering distributions $(\mu_{j})_{j\in\llbracket 0,J-1\rrbracket}$ are close to Gaussian after lifting to the joint state and data space, then the distribution $\mu^{{\rm EK}}_{j}$ given by the mean field ensemble Kalman filter (2.4) is close to the filtering distribution $\mu_{j}$ given by (1.9) for all $j\in\llbracket 0,J\rrbracket$.
 Theorem 2.1 (Stability: Mean Field Ensemble Kalman Filter).
-Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained respectively from the dynamical systems (1.11) and (2.4), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$.
+Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 0,J\rrbracket}$ are given respectively by the dynamical systems (1.9) and (2.4), initialized at the Gaussian probability measure $\mu_{0}=\mu_{0}^{{\rm EK}}\in~{}\mathcal{G}(^{d_{u}})$.
 That is,
@@
   1. 1.
-    For any probability measure $\mu$ with finite first and second order polynomial moments $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$, the means of the probability measures $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their covariances are bounded both from above and from below.
+    For any probability measure $\mu$ with finite first and second order polynomial moments $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$, the means of $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their covariances are bounded both from above and from below.
 The constants in these bounds depend only on the parameters $\kappa_{\Psi}$, $\kappa_{h}$, $\Sigma,\Gamma$ and on $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$.
@@
     Let $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 1,J\rrbracket}$ denote the probability measures obtained respectively from the dynamical systems (1.9) and (2.4), initialized at the same Gaussian measure $\mu_{0}=\mu_{0}^{{\rm EK}}\in~{}\mathcal{G}(^{d_{u}})$.
-Then for any $q\in\mathbb{N}$, there exist constants $M_{q},M_{q}^{\rm EK}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
+Then for any integer $q\geqslant 2$, there exist constants $M_{q},M_{q}^{{\rm EK}}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
     $\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}(\mu_{j})\leqslant M_{q}\quad\text{and}\quad\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}\bigl{(}\mu^{{\rm EK}}_{j}\bigr{)}\leqslant M_{q}^{{\rm EK}}.$
@@
   3. 3.
-    For any Gaussian measure $\mu\in\mathcal{G}(^{d_{u}}\times^{d_{y}})$, it holds that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma B.7 and also [calvello2025ensemble].
+    For any Gaussian measure $\mu\in\mathcal{G}(^{d_{u}}\times^{d_{y}})$, we have that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma B.7 and [calvello2025ensemble, carrillo2022ensemble].
   4. 4.
-    The map $\mathsf{P}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{P}}$ depending on the parameters $\kappa_{\Psi}$ and $\Sigma$.
+    The map $\mathsf{P}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$.
+The Lipschitz constant $L_{\mathsf{P}}$ depends only on the parameters $\kappa_{\Psi}$ and $\Sigma$.
 See Lemma B.8.
   5. 5.
-    The map $\mathsf{Q}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{Q}}$ depending on the parameters $\kappa_{h}$ and $\Gamma$.
+    The map $\mathsf{Q}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$.
+The Lipschitz constant $L_{\mathsf{Q}}$ depends only on the parameters $\kappa_{h}$ and $\Gamma$.
 See Lemma B.9.
   6. 6.
-    The map $\mathsf{B}_{j}$ satisfies, for any $\pi\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, the bound
+    The map $\mathsf{B}_{j}$ satisfies that, for all $\pi\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, it holds
     $\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{B}_{j}\mathsf{G}\pi,\mathsf{B}_{j}\pi)\leqslant C_{\mathsf{B}}d_{g}(\mathsf{G}\pi,\pi),$
     where $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$.
-This statement concerns the stability of the $\mathsf{B}_{j}$ operator between a measure and its Gaussian approximation.
+Namely, this item regards the stability of the operator $\mathsf{B}_{j}$ between a probability measure and its Gaussian projection.
 See Lemma B.10.
   7. 7.
-    The map $\mathsf{T}_{j}$ satisfies the following bound: for all $R\geqslant 1$, it holds for all probability measures $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$ that
+    The map $\mathsf{T}_{j}$ satisfies the following bound: for all $R\geqslant 1$, it holds for all $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$ that
     $\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p),$
-    for a constant $L_{\mathsf{T}}=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$.
-This statement may be viewed as a local Lipschitz continuity result in the case where the second argument of $d_{g}$ is restricted to the range of $\mathsf{Q}\mathsf{P}$.
+    for a constant $L_{\mathsf{T}}=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},{\ell_{h}},\Sigma,\Gamma\bigr{)}$.
+Namely, this item may be regarded as a local Lipschitz continuity result.
 See Lemma B.11.
 Proof of Theorem 2.1.
-In what follows we refer to the preceding itemized list to clarify the proof.
-For notational simplicity it is helpful to define the following measure of the difference between the true filtering distribution and its Gaussian projection:
+Within the following argument we make reference to the list of items outlined above.
+We begin by defining for ease of exposition the following measure of difference between the true filter and its Gaussian projection:
 $\varepsilon:=\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$ (2.7)
-Assume throughout the following that $j\in\llbracket 0,J-1\rrbracket.$ The proof rests on the following use of the triangle inequality:
+We assume throughout the argument that $j\in\llbracket 0,J-1\rrbracket.$ The proof rests upon the following application of the triangle inequality:
 $d_{g}(\mu^{{\rm EK}}_{j+1},\mu_{j+1})$ $=d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$
 $\leqslant d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)+d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)+d_{g}\left(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right).$ (2.8a)
-We have used the fact that $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by Item 3 (Lemma B.7).
-Item 2 (Lemmas B.3 and B.5) shows that, for any $q\in\mathbb{N}$, there exist constants $M_{q},M_{q}^{\rm EK}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
+Here we have used the fact that $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by Item 3 (Lemma B.7).
+Item 2 (Lemmas B.3 and B.5) shows that, for any integer $q\geqslant 2$, there exist constants $M_{q},M_{q}^{{\rm EK}}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
 $\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}(\mu_{j})\leqslant M_{q}\quad\text{and}\quad\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}\bigl{(}\mu^{{\rm EK}}_{j}\bigr{)}\leqslant M_{q}^{{\rm EK}}.$
-Therefore by Item 1 (Lemma B.2), there is a constant $R\geqslant 1$, depending on the covariance matrices $\Sigma$, $\Gamma$, the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, and the moment bounds $M_{2}$ and $M_{2}^{\rm EK}$, such that for any $j\in\llbracket 0,J-1\rrbracket$ it holds that $\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$.
-By Items 4, 5 and 7 (Lemmas B.11, B.9 and B.8), the first term in (2.8a) satisfies
-$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$ $\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}d_{g}\left(\mu_{j}^{\rm EK},\mu_{j}\right),$ (2.9)
-where, for conciseness, we omitted the dependence of the constants on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
-Since it holds that $\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\in\mathcal{G}_{R}(^{d_{u}}\times^{d_{y}})$ by definition of $R$, the second term in (2.8a) may be bounded using Item 7 (Lemma B.11) and the definition in (2.7) of $\varepsilon$:
-$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}d_{g}\left(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}\varepsilon.$
-Finally, the third term in (2.8a) can be bounded using item 6 (Lemma B.10) and the definition in (2.7) of $\varepsilon$:
+Therefore by Item 1 (Lemma B.2), there is a constant $R\geqslant 1$, depending on the covariances $\Sigma$, $\Gamma$, the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, and the moment bounds $M_{2}$ and $M_{2}^{{\rm EK}}$, such that for any $j\in\llbracket 0,J-1\rrbracket$ it holds that $\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}}\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$.
+In view of Items 4, 5 and 7 (Lemmas B.11, B.9 and B.8), the first term in (2.8a) satisfies
+$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$ $\leqslant L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}d_{g}\left(\mu_{j}^{{\rm EK}},\mu_{j}\right),$ (2.9)
+where, for conciseness, we have omitted the dependence of the constants in the bound on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
+By definition of $R$ it holds that $\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\in\mathcal{G}_{R}(^{d_{u}}\times^{d_{y}})$, hence the second term in (2.8a) may be bounded using Item 7 (Lemma B.11) and the definition in of $\varepsilon$ in (2.7).
+Indeed, it holds that
+$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}d_{g}\left(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}\varepsilon.$
+Using Item 6 (Lemma B.10) and the definition of $\varepsilon$ in (2.7) we establish the following bound on the third term in (2.8a):
 $d_{g}\left(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant C_{\mathsf{B}}d_{g}\left(\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant C_{\mathsf{B}}\varepsilon.$
-Therefore, letting $\ell=L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}$, we have shown that
-$d_{g}(\mu^{\rm EK}_{j+1},\mu_{j+1})$ $\leqslant\ell d_{g}(\mu^{\rm EK}_{j},\mu_{j})+\bigl{(}L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}+C_{\mathsf{B}}\bigr{)}\varepsilon$
-and the conclusion follows from the discrete Grönwall lemma, using the fact that $\mu_{0}^{\rm EK}=\mu_{0}.$
+Therefore, setting $\ell=L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}$, we have that
+$d_{g}(\mu^{{\rm EK}}_{j+1},\mu_{j+1})$ $\leqslant\ell d_{g}(\mu^{{\rm EK}}_{j},\mu_{j})+\bigl{(}L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}+C_{\mathsf{B}}\bigr{)}\varepsilon.$
+The conclusion then follows from the discrete Grönwall lemma, using the fact that $\mu_{0}^{{\rm EK}}=\mu_{0}.$
 2.3 Error Estimate: Mean Field Ensemble Kalman Filter
-Theorem 2.1 shows that the mean field ensemble Kalman filter error can be made arbitrarily small if the true filter is arbitrarily close to its Gaussian projection, in state-observation space.
-This “closeness to Gaussian” assumption can be satisfied in our setting of unbounded vector fields by considering small perturbations of affine vector fields, stated and proved in Proposition 2.2.
+It is possible to deduce from the result of Theorem 2.1 that the error between the true filter and the mean field ensemble Kalman filter can be arbitrarily small if the true filter is arbitrarily close to its Gaussian projection, in state-observation space.
+This condition on the true filter, which we refer to as “closeness to Gaussian”, can be satisfied in the setting of unbounded vector fields by considering small perturbations of affine vector fields, as we prove in Proposition 2.2.
 Combining Proposition 2.2 with Theorem 2.1 gives an error estimate for the mean field ensemble Kalman filter, yielding Theorem 2.3.
@@
 In what follows $(\mu^{0}_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ denote the true filtering distributions associated with functions $(\Psi_{0},h_{0})$ and $(\Psi,h)$, respectively, initialized at the same Gaussian measure $\mathcal{N}(m_{0},C_{0})$.
-Furthermore, we let $\mathsf{P}_{0}$ and $\mathsf{Q}_{0}$ denote the kernel integral operators (1.7) and (1.9) defined by the specific vector fields $(\Psi_{0},h_{0}).$ By Lemma B.3, the filtering distributions have bounded second moments.
+Furthermore, we let $\mathsf{P}_{0}$ and $\mathsf{Q}_{0}$ denote the kernel integral operators (1.7a) and (1.8a) defined by the specific vector fields $(\Psi_{0},h_{0}).$ By Lemma B.3, the filtering distributions have bounded second moments.
 Let
@@
 Theorem 2.3 (Error Estimate: Mean Field Ensemble Kalman Filter).
-Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained respectively from the dynamical systems (1.11) and (2.4) initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$.
+Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 0,J\rrbracket}$ are given respectively by the dynamical systems (1.9) and (2.4) initialized at the Gaussian measure $\mu_{0}=\mu_{0}^{{\rm EK}}\in~{}\mathcal{G}(^{d_{u}})$.
 That is,
@@
 Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H2).
-Fix $\kappa_{\Psi},\kappa_{h},\ell_{h}>0$ and let the vector fields $\Psi_{0},h_{0}$ be affine functions satisfying Assumptions (V1) and (V2), respectively, while also satisfying Assumptions (H2) and (H3) with $\kappa_{\Psi},\kappa_{h}$.
-Then for all $\varepsilon\in[0,1]$, there exists a constant  $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{\max\{3+d_{u},4+d_{y}\}}$, such that for any $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$ with $h$ satisfying Assumption (H4), it holds that
+Let the vector fields $\Psi_{0},h_{0}$ be affine functions satisfying Assumptions (V1) and (V2), respectively, while also satisfying Assumptions (H3) and (H4) with some constants $\kappa_{\Psi},\kappa_{h}>0$.
+Additionally, let $\ell_{\Psi}>0$ be a constant.
+Then there exists a constant  $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{2\cdot\max\{3+d_{u},4+d_{y}\}}$, such that for all $\varepsilon\in[0,1]$ and all $(\Psi,h)\in B_{L^{\infty}}\bigl{(}(\Psi_{0},h_{0}),\varepsilon\bigr{)}$ with $h$ satisfying Assumption (H5), it holds that
 $d_{g}(\mu^{{\rm EK}}_{J},\mu_{J})\leqslant C\varepsilon.$
 Proof.
-Since Assumption H is satisfied, it is possible to apply the result of Theorem 2.1 to deduce that there exists $C=C\bigl{(}\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$ such that
+Since Assumption H is satisfied, it is possible to apply the result of Theorem 2.1 to deduce that there exists $C=C\bigl{(}\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$ such that
 $d_{g}(\mu^{{\rm EK}}_{J},\mu_{J})\leqslant C\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
@@
 2.4 Error Estimate: Finite Particle Ensemble Kalman Filter
-In this subsection, we combine the results from the work in [le2009large] with stability Theorem 2.1, together with approximation Theorem 2.3, to derive a quantitative error estimate between the finite particle ensemble Kalman filter and the true filter in the non-linear setting.
+In this subsection, we combine the results from the work in [le2009large] with stability Theorem 2.1, together with approximation Theorem 2.3, to derive a quantitative error estimate between the finite particle ensemble Kalman filter and the true filter in the nonlinear setting.
 In order to define an appropriate metric we introduce the following class of vector fields.
@@
 Additionally, let $\ell_{\Psi}>0$ be a constant.
-Then, for all $\varepsilon\in[0,1]$, it holds that for any $\Psi\in B_{L^{\infty}}\bigl{(}\Psi_{0},\varepsilon\bigr{)}$ satisfying $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\ell_{\Psi},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{\max\{3+d_{u},4+d_{y},16\cdot 3^{J-1}\}}$, such that
-$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant C\Bigl{(}\frac{1}{\sqrt{N}}+{\varepsilon}\Bigr{)},$
-for any $\phi$ satisfying Assumption P2.
+Then, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\ell_{\Psi},\Sigma,\Gamma,J\bigr{)}$, where $q\coloneqq{2\cdot\max\{3+d_{u},4+d_{y},2\cdot 4^{J}\}}$, such that for all $\varepsilon\in[0,1]$ and all $\Psi\in B_{L^{\infty}}\bigl{(}\Psi_{0},\varepsilon\bigr{)}$ satisfying $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, the following bound holds for any $\phi$ satisfying Assumption P2:
+$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant C\Bigl{(}\frac{1}{\sqrt{N}}+{\varepsilon}\Bigr{)}.$
 The proof presented hereafter relies on the following elements.
@@
 In particular, this holds under Assumptions (H1) and (H2) on the data and covariances of the noise processes, the linearity of the vector field $h$ and Assumption (H3) on $\Psi$ with the additional assumption that $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$.
-In Lemma D.1 we mimic the proof of [le2009large, Proposition 4.4] and  [le2009large, Theorem 5.2] to gain insight into the dependence of the constant prefactor multiplying $1/\sqrt{N}$ on the parameters of the Gaussian initial condition and the number of steps $J$.
+In Lemma D.1 we provide a self-contained proof of [le2009large, Theorem 5.2] to gain insight into the dependence of the constant prefactor multiplying $1/\sqrt{N}$ on the parameters of the Gaussian initial condition and the number of steps $J$.
   3. 3.
@@
 $d_{g}\bigl{(}\mu^{{\rm EK}}_{J},\mu_{J}\bigr{)}\leqslant C{\varepsilon},$ (2.13)
-where $C$ is a constant depending on $\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma$.
+where $C$ is a constant depending on $\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma$.
 The fact that $h$ is assumed to be linear and the additional assumption that $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$ allows us to apply the result of Lemma D.1, so that
 $\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}},$ (2.14)
-where $C$ is a constant depending on $\mathcal{M}_{16\cdot 3^{J-1}}(\mu_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
+where $C$ is a constant depending on $\mathcal{M}_{4^{J+1}}(\mu_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
 In view of (2.12), combining (2.13) and (2.14) yields the desired result.
@@
 3 Discussion and Future Directions
-In this paper we have presented the first analysis, in the setting of a filtering problem defined by nonlinear state space dynamics, quantifying the error between the empirical measure obtained by the ensemble Kalman filter and the true filtering distribution.
-The analysis for the EnKF outlined is based on the proof methodology developed for particle filters in [rebeschini2015can].
-This work extends the initial results of [carrillo2022ensemble] to a setting of practical interest, for which dynamical models and observation operators are allowed to be unbounded and nonlinear.
-The work opens a new avenue for analysis in the quantitative understanding of the performance of the ensemble Kalman methodology applied to nonlinear problems.
-To this end, we identify the following avenues for further work.
+In this paper we have presented the first analysis, in the setting of a filtering problem defined by nonlinear state space dynamics, quantifying the error between the empirical measure obtained by the finite particle ensemble Kalman filter and the true filtering distribution.
+The analysis for the EnKF outlined is based on the proof methodology developed for particle filters in [rebeschini2015can] and developed for the mean field ensemble Kalman filter in [carrillo2022ensemble].
+Our work goes beyond [carrillo2022ensemble] by considering finite particle filters and by allowing for dynamical models and observation operators that are unbounded.
+The work builds substantially on the new avenue for analysis of the ensemble Kalman methodology that was introduced in [carrillo2022ensemble] but leaves open numerous avenues of investigation for further work; we detail these.
   1. 1.
-    As surveyed in [calvello2025ensemble], the ensemble Kalman filtering methodology may be used for solving inverse problems and for sampling; it is of interest to extend this analysis to the ensemble Kalman based inversion algorithms outlined in that paper.
+    As surveyed in [calvello2025ensemble], the ensemble Kalman filtering methodology may be used for solving inverse problems and for sampling; it is of interest to extend the analysis in our paper to the ensemble Kalman based inversion algorithms outlined in that paper.
   2. 2.
@@
 The research of PM is supported by the projects SWIDIMS (ANR-20-CE40-0022) and CONVIVIALITY (ANR-23-CE40-0003) of the French National Research Agency.
-The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship, and by the SciAI Center, funded by the Office of Naval Research (ONR), Grant Number N00014-23-1-2729; these grants also support the work of EC.
+The work of AMS is supported by a Department of Defense Vannevar Bush Faculty Fellowship, and by the SciAI Center, funded by the Office of Naval Research (ONR), Grant Number N00014-23-1-2729.
+The work of EC was supported by the Resnick Sustainability Institute, and also by the SciAI Center and by the Vannevar Bush Faculty Fellowship held by AMS.
 UV is partially supported by the European Research Council (ERC) under the EU Horizon 2020 programme (grant agreement No 810367), and by the Agence Nationale de la Recherche under grants ANR-21-CE40-0006 (SINEQ) and ANR-23-CE40-0027 (IPSO).
 The work was initiated during The Jacques-Louis Lions Lectures given by AMS at Sorbonne Université in December 2023; the authors are grateful to Albert Cohen for hosting this event.
+[…]
 Appendix A Auxiliary Results
@@
 Lemma A.1.
-Let $X$ be a random variable taking values in ${}^{d_{u}}$ with finite second moment, and let $m:=\mathbb{E}[X]$. Then it holds that
-$\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}=\mathbb{E}\bigl{[}XX^{\mathsf{T}}\bigr{]}-mm^{\mathsf{T}}$ (A.1)
+Let $X$ be a random vector in ${}^{d_{u}}$ with finite second moment. Then it holds that
+$\mathbb{E}\Bigl{[}\bigl{(}X-\mathbb{E}[X]\bigr{)}\bigl{(}X-\mathbb{E}[X]\bigr{)}^{\mathsf{T}}\Bigr{]}=\mathbb{E}\bigl{[}XX^{\mathsf{T}}\bigr{]}-\mathbb{E}[X]\mathbb{E}[X]^{\mathsf{T}}$ (A.1)
 and
-$\forall a\in^{d_{u}},\qquad\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}\succcurlyeq\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}.$ (A.2)
+$\forall a\in^{d_{u}},\qquad\mathbb{E}\Bigl{[}\bigl{(}X-a\bigr{)}\bigl{(}X-a\bigr{)}^{\mathsf{T}}\Bigr{]}\succcurlyeq\mathbb{E}\Bigl{[}\bigl{(}X-\mathbb{E}[X]\bigr{)}\bigl{(}X-\mathbb{E}[X]\bigr{)}^{\mathsf{T}}\Bigr{]}.$ (A.2)
 Proof.
-It holds that
-$\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}$ $=\mathbb{E}\Bigl{[}\bigl{(}(X-m)+(m-a)\bigr{)}\bigl{(}(X-m)+(m-a)\bigr{)}^{\mathsf{T}}\Bigr{]}$
-$=\mathbb{E}\Bigl{[}(X-m)(X-m)^{\mathsf{T}}\Bigr{]}+(m-a)(m-a)^{\mathsf{T}}.$
-Setting $a=0$, we obtain (A.1). The inequality (A.2) follows by noting that the second term in the last expression is positive semidefinite. ∎
+The statements follow from the following equalities:
+$\mathbb{E}\Bigl{[}(X-a)(X-a)^{\mathsf{T}}\Bigr{]}$ $=\mathbb{E}\Bigl{[}\Bigl{(}\bigl{(}X-\mathbb{E}[X]\bigr{)}+\bigl{(}\mathbb{E}[X]-a\bigr{)}\Bigr{)}\Bigl{(}\bigl{(}X-\mathbb{E}[X]\bigr{)}+\bigl{(}\mathbb{E}[X]-a\bigr{)}\Bigr{)}^{\mathsf{T}}\Bigr{]}$
+$=\mathbb{E}\Bigl{[}\bigl{(}X-\mathbb{E}[X]\bigr{)}\bigl{(}X-\mathbb{E}[X]\bigr{)}^{\mathsf{T}}\Bigr{]}+\bigl{(}\mathbb{E}[X]-a\bigr{)}\bigl{(}\mathbb{E}[X]-a\bigr{)}^{\mathsf{T}}.\qed$
 Lemma A.2.
-Denote by $g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m,S)$ the Lebesgue density of $\mathcal{N}(m,S)$ and by $\mathcal{S}^{n}_{\alpha}$ the set of symmetric $n\times n$ matrices $M$ satisfying
-$\frac{1}{\alpha}I_{n}\preccurlyeq M\preccurlyeq\alpha I_{n}.$ (A.3)
-Then for all $n\in\mathbb{N}^{+}$ and $\alpha\geqslant 1$, there exists $L_{n,\alpha}>0$ such that for all parameters $(c_{1},m_{1},S_{1})\in\times^{n}\times\mathcal{S}^{n}_{\alpha}$ and $(c_{2},m_{2},S_{2})\in\times^{n}\times\mathcal{S}^{n}_{\alpha}$,
-$\lVert h\rVert_{\infty}\leqslant L_{n,\alpha}\lVert h\rVert_{1},\qquad h(y)=c_{1}g(y;m_{1},S_{1})-c_{2}g(y;m_{2},S_{2}).$ (A.4)
+Let $g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m,S)$ be the Lebesgue density of $\mathcal{N}(m,S)$ and $\mathcal{S}^{n}_{\alpha}$ the set of symmetric $n\times n$ matrices $M$ satisfying
+$\frac{1}{\tau}I_{n}\preccurlyeq M\preccurlyeq\tau I_{n}.$
+Then for any $n\in\mathbb{N}^{+}$ and $\tau\geqslant 1$, there is $L_{n,\tau}>0$ such that for all parameters $(c_{1},m_{1},S_{1})\in\times^{n}\times\mathcal{S}^{n}_{\tau}$ and $(c_{2},m_{2},S_{2})\in\times^{n}\times\mathcal{S}^{n}_{\tau}$ it holds that
+$\lVert h\rVert_{\infty}\leqslant L_{n,\tau}\lVert h\rVert_{1},\qquad h(y)=c_{1}g(y;m_{1},S_{1})-c_{2}g(y;m_{2},S_{2}).$
 Proof.
-The lemma as stated may be found in [carrillo2022ensemble, Lemma 14], where a complete proof is given. ∎
+The lemma as stated may be found in [carrillo2022ensemble, Lemma A.4], where a complete proof is given. ∎
 Lemma A.3.
-Let $\mathsf{P}$ and $\mathsf{Q}$ denote the operators on probability measures given respectively in (1.7) and (1.9).
+Let $\mathsf{P}$ and $\mathsf{Q}$ denote the operators on probability measures given respectively in (1.7a) and (1.8a).
 Let Assumptions (H3), (H4), (H5) and (H2) be satisfied and suppose that $\mu\in\mathcal{P}(^{d_{u}})$ satisfies, for some $q>0$, the moment bound
@@
 Appendix B Technical Results for Theorem 2.1
-We show moment bounds in Subsection B.1, we recall that on Gaussian measures the action of the conditioning map and the Kalman transport map are equivalent in Subsection B.2, and we prove stability results in Subsection B.3.
+We establish moment bounds in Subsection B.1, we recall that on Gaussian measures the action of the conditioning map and the Kalman transport map are equivalent in Subsection B.2, and we prove stability results in Subsection B.3.
 B.1 Moment Bounds
 Lemma B.1 (Moment Bounds).
-Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumption H, it holds that
+Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$.
+Under Assumptions (H3) and (H2), it holds that
 $\bigl{\lvert}\mathcal{M}(\mathsf{P}\mu)\bigr{\rvert}\leqslant\kappa_{\Psi}\bigl{(}1+\mathcal{M}_{1}(\mu)\bigr{)},\qquad\Sigma\preccurlyeq\mathcal{C}(\mathsf{P}\mu)\preccurlyeq\Sigma+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}I_{d_{u}}.$ (B.1)
 Proof.
-Using the definition of $\mathsf{P}$ in (1.7), it holds that
+The proof of this lemma follows the steps of the proof of [carrillo2022ensemble, Lemma B.1]; however, here different bounds reflecting the assumptions on $\Psi$ and $h$ in this paper are required.
+Using the definition of $\mathsf{P}$ in (1.7a), it holds that
 $\mathcal{M}(\mathsf{P}\mu)=\int_{{}^{d_{u}}}u\,\mathsf{P}\mu(u)\,\mathrm{d}u$ $=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\int_{{}^{d_{u}}}u\exp\left(-\frac{1}{2}\big{\lvert}u-\Psi(v)\big{\rvert}_{\Sigma}^{2}\right)\,\mu(\mathrm{d}v)\,\mathrm{d}u$
@@
 The first inequality in (B.1) then follows from Assumption (H3).
-For the second inequality in (B.1), we first note that for any $a,v\in^{d_{u}}$ it holds that
-$\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u$
-$\qquad\succcurlyeq\int_{{}^{d_{u}}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u;$ (B.2)
-this follows from Lemma A.1 by noting that $\Psi(v)$ is the mean under the Gaussian probability measure $\mathcal{N}\bigl{(}\Psi(v),\Sigma\bigr{)}.$ Letting now $a$ be the mean under the measure $\mathsf{P}\mu$, we observe that, by conditioning on $v$ and using (B.1),
-$\mathcal{C}(\mathsf{P}\mu)$ $=\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\,\mathsf{P}\mu(u)\,\mathrm{d}u$
-$=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
-$\succcurlyeq\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
-$=\int_{{}^{d_{u}}}\Sigma\,\mu(\mathrm{d}v)=\Sigma,$
-hence $\mathcal{C}(\mathsf{P}\mu)\succcurlyeq\Sigma$.
-On the other hand, using Lemma A.1 again and noting that, by the Cauchy-Schwarz inequality, $ww^{\mathsf{T}}\preccurlyeq(w^{\mathsf{T}}w)I_{d_{u}}$ for any vector $w\in^{d_{u}}$, we deduce that
+Obtaining the lower bound of the second inequality in (B.1) may be done identically to the proof of [carrillo2022ensemble, Lemma B.1].
+We turn our attention to obtaining the upper bound.
+Using Lemma A.1 and noting that $ww^{\mathsf{T}}\preccurlyeq(w^{\mathsf{T}}w)I_{d_{u}}$ for any vector $w\in^{d_{u}}$ by the Cauchy-Schwarz inequality, we deduce that
 $\mathcal{C}(\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d_{u}}}u\otimes u\,\mathsf{P}\mu(u)\,\mathrm{d}u$
@@
 Lemma B.2.
-Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumption H, it holds that
+Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$.
+Under Assumptions (H3), (H4) and (H2), it holds that
 $\bigl{\lvert}\mathcal{M}^{u}(\mathsf{Q}\mathsf{P}\mu)\bigr{\rvert}$ $\leqslant\kappa_{\Psi}\bigl{(}1+\mathcal{M}_{1}(\mu)\bigr{)},$ (B.3a)
@@
 Proof.
-The inequalities (B.4a) and (B.4b) follow from Assumption H and the fact that
+The inequalities (B.3a) and (B.3b) follow from the assumptions and the fact that
 $\mathcal{M}(\mathsf{Q}\mathsf{P}\mu)=\begin{pmatrix}\mathcal{M}(\mathsf{P}\mu)\\ \mathsf{P}\mu[h]\end{pmatrix}.$ (B.5)
@@
 The result then follows by applying Assumption (H3).
-We now turn our attention to establishing inequality (B.5a).
-To this end, let $\phi\colon^{d_{u}}\to^{d_{u}+d_{y}}$ define the map $\phi(u)=\bigl{(}u,h(u)\bigr{)}$, and let $\phi_{\sharp}$ denote the associated pushforward map on measures.
-We note that
+To obtain the covariance bounds, we proceed using analogous steps to the proof of [carrillo2022ensemble, Lemma B.2]; however, here different bounds reflecting the assumptions on $\Psi$ and $h$ in this paper are required.
+We begin by establishing inequality (B.4a).
+To this end, letting $\phi\colon^{d_{u}}\to^{d_{u}+d_{y}}$ define the map $\phi(u)=\bigl{(}u,h(u)\bigr{)}$, it holds that
 $\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)=\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)+\begin{pmatrix}0_{d_{u}\times d_{u}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&\Gamma\end{pmatrix}=\begin{pmatrix}\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)&\mathcal{C}^{uy}(\phi_{\sharp}\mathsf{P}\mu)\\ \mathcal{C}^{yu}(\phi_{\sharp}\mathsf{P}\mu)&\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)+\Gamma\end{pmatrix}.$ (B.7)
-It holds for any $(a,b)\in^{d_{u}}\times^{d_{y}}$ that
-$2\begin{pmatrix}aa^{\mathsf{T}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&bb^{\mathsf{T}}\end{pmatrix}-\begin{pmatrix}a\\ b\end{pmatrix}\otimes\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}a\\ -b\end{pmatrix}\otimes\begin{pmatrix}a\\ -b\end{pmatrix}\succcurlyeq 0_{(d_{u}+d_{y})\times(d_{u}+d_{y})}.$
-Hence, it follows that
+As shown in the proof of [carrillo2022ensemble, Lemma B.2] we have, for all $(a,b)\in^{d_{u}}\times^{d_{y}}$,
 $\begin{pmatrix}a\\ b\end{pmatrix}\otimes\begin{pmatrix}a\\ b\end{pmatrix}\preccurlyeq 2\begin{pmatrix}aa^{\mathsf{T}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&bb^{\mathsf{T}}\end{pmatrix},$
-which enables to deduce, using Lemmas A.1 and B.1 and Assumption H, that
+from which we establish, using Lemmas A.1 and B.1 and Assumptions (H3), (H4) and (H2), that
 $\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d_{u}}}\begin{pmatrix}u\\ h(u)\end{pmatrix}\otimes\begin{pmatrix}u\\ h(u)\end{pmatrix}\,\mathsf{P}\mu(u)\,\mathrm{d}u$
@@
 $\preccurlyeq 2\begin{pmatrix}\Sigma+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}I_{d_{u}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}\end{pmatrix},$ (B.8)
-where we used (B.3) and the calculation resulting in (B.7a) in the last inequality.
-Noting this inequality in combination with (B.8) yields the upper bound (B.5a).
-To establish the lower bound (B.5b), we first note that by the Cauchy–Schwarz inequality, for any $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
+where we applied (B.2) and the calculation resulting in (B.6a) in the last inequality. Noting this inequality in combination with (B.7) yields the upper bound (B.4a). We now show that
+$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}\geqslant(1-\varepsilon)\sigma\left\lvert a\right\rvert^{2}+\left(\gamma-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\right)\lvert b\rvert^{2}.$ (B.9)
+in order to establish the lower bound (B.4b).
+The argument is similar to that in [carrillo2022ensemble, Lemma B.2] but is explicit in certain constants whose dependencies on moments need to be controlled in this paper.
+We first note that for any $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
 $\left\lvert a^{\mathsf{T}}\mathcal{C}^{uy}(\pi)b\right\rvert$ =×$=\int_{{}^{d_{u}}\times^{d_{y}}}\Bigl{(}a^{\mathsf{T}}\bigl{(}u-\mathcal{M}^{u}(\pi)\bigr{)}\Bigr{)}\Bigl{(}b^{\mathsf{T}}\bigl{(}y-\mathcal{M}^{y}(\pi)\bigr{)}\Bigr{)}\pi(\mathrm{d}u\mathrm{d}y)$
-$\leqslant\sqrt{a^{\mathsf{T}}\mathcal{C}^{uu}(\pi)a}\,\sqrt{b^{\mathsf{T}}\mathcal{C}^{yy}(\pi)b}.$
-Hence, by Young’s inequality, we deduce that for all $\varepsilon\in(0,1)$ and for all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
+$\leqslant\sqrt{a^{\mathsf{T}}\mathcal{C}^{uu}(\pi)a}\,\sqrt{b^{\mathsf{T}}\mathcal{C}^{yy}(\pi)b},$
+by the Cauchy–Schwarz inequality. Hence, we deduce that for all $\varepsilon\in(0,1)$ and for all $(a,b)\in^{d_{u}}\times^{d_{y}}$
 $\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)a-\left(\frac{1}{\varepsilon}-1\right)b^{\mathsf{T}}\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)b$
 $\geqslant(1-\varepsilon)a^{\mathsf{T}}\Sigma a-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]},$
-where we applied (B.1) and the bound $\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)\preccurlyeq\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}I_{d_{y}}$ in the last inequality. We then use (B.8) to deduce that
+where we applied Young’s inequality for the bound in the first line, and (B.1) and the bound $\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)\preccurlyeq\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}I_{d_{y}}$ for the bound in the second line.
+We then apply (B.7) to find
 $\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\Sigma a-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}+b^{\mathsf{T}}\Gamma b$
 $\geqslant(1-\varepsilon)\sigma\left\lvert a\right\rvert^{2}+\left(\gamma-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\right)\lvert b\rvert^{2}.$
-Taking $\varepsilon=\frac{2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}$, so that the coefficient of $\lvert b\rvert^{2}$ is $\frac{\gamma}{2}$, we obtain
+Choosing $\varepsilon=\frac{2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}$, so that the coefficient of the $\lvert b\rvert^{2}$ term is $\frac{\gamma}{2}$, we obtain
 $\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}\geqslant\frac{\gamma\sigma}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}\lvert a\rvert^{2}+\frac{\gamma}{2}\lvert b\rvert^{2}\geqslant\min\left\{\frac{\gamma\sigma}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}},\frac{\gamma}{2}\right\}\left(\lvert a\rvert^{2}+\lvert b\rvert^{2}\right).$ (B.10)
@@
 $\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\succcurlyeq\gamma\frac{\min\left\{2\sigma,\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\right\}}{2\gamma+4\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}I_{d_{u}+d_{y}}.$
-In order to finally obtain (B.5b), we use (B.7b) to deduce that
+In order to deduce (B.4b), we use (B.6b) to find that
 $\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\leqslant 2\kappa_{h}^{2}\Bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\Bigr{)},$
-which concludes the proof. ∎
+from which we conclude the desired result. ∎
 Lemma B.3 (Moment Bound for the True Filtering Distribution).
+Let $q\geqslant 2$ be an integer.
 Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (1.9) initialized at the probability measure $\mu_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th order polynomial moment $\mathcal{M}_{q}(\mu_{0})<\infty$.
-If Assumption H holds then there exists $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$ such that
+If Assumptions (H1), (H3), (H4) and (H2) hold then there exists $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$ such that
 $\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}(\mu_{j})\leqslant C.$
@@
 $\leqslant C\Bigl{(}1+\mathcal{M}_{q}(\mu_{j})\Bigr{)},$ (B.13c)
-where (B.13a) follows from Fubini’s theorem, the inequality (B.13b) from properties of Gaussians, and (B.13c) from application of Assumption H.
+where (B.13a) follows from Fubini’s theorem, the inequality (B.13b) from properties of Gaussians and Assumption (H2), and (B.13c) from application of Assumption (H3).
 We note that in (B.13c) the constant $C$ depends on $\kappa_{\Psi},\Sigma$.
@@
 Lemma B.5 (Moment Bound for the Approximate Filtering Distribution).
+Let $q\geqslant 2$ be an integer.
 Assume that the probability measures $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (2.4) initialized at the probability measure $\mu^{{\rm EK}}_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th polynomial order moment $\mathcal{M}_{q}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)}<\infty$.
-If Assumption H holds then there exists $C=C\Bigl{(}\mathcal{M}_{q}\bigl{(}\mu_{0}^{\rm EK}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ such that
+If Assumptions (H1), (H3), (H4) and (H2) hold then there exists $C=C\Bigl{(}\mathcal{M}_{q}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ such that
 $\max_{j\in\llbracket 0,J\rrbracket}\mathcal{M}_{q}\bigl{(}\mu^{{\rm EK}}_{j}\bigr{)}\leqslant C.$
@@
 Lemma B.6.
-For $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ with finite second moments, it holds that
+Let $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ have finite second moments, then the following bounds hold:
 $\bigl{\lvert}\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\bigr{\rvert}$ $\leqslant\frac{1}{2}d_{g}(\mu_{1},\mu_{2}),$
@@
 Proof.
-The lemma as stated may be found in [carrillo2022ensemble, Lemma 20], where a complete proof is given. ∎
+The lemma as stated may be found in [carrillo2022ensemble, Lemma B.4], where a complete proof is given. ∎
 B.2 Action of $\mathsf{T}_{j}$ on Gaussians
@@
 Fix $y^{\dagger}_{j+1}\in^{d_{y}}$ and let $\pi$ be a Gaussian measure over ×${}^{d_{u}}\times^{d_{y}}$.
-Then the probability measure $\mathsf{B}_{j}\pi$, with $\mathsf{B}_{j}$ defined in (1.10), coincides with the probability measure $\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi$, where $\mathscr{T}$ is defined in (2.3).
+Then the probability measure $\mathsf{B}_{j}\pi$, with $\mathsf{B}_{j}$ defined in (1.8b) is equivalent to the probability measure $\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi$, where $\mathscr{T}$ is defined in (2.3b).
 Proof.
-The lemma as stated may be found in [carrillo2022ensemble, Lemma 21], where a complete proof is given. ∎
+The lemma as stated may be found in [carrillo2022ensemble, Lemma B.5], where a complete proof is given. ∎
 B.3 Stability Results
 Lemma B.8 (Map $\mathsf{P}$ is Lischitz).
-Suppose that $\Psi$ and $\Sigma$ satisfy Assumptions (H2) and (H5), respectively. Then it holds that
+Suppose that $\Sigma$ and $\Psi$ satisfy Assumptions (H3) and (H2), respectively. Then it holds that
 $\forall(\mu,\nu)\in\mathcal{P}(^{d_{u}})\times\mathcal{P}(^{d_{u}}),\qquad d_{g}(\mathsf{P}\mu,\mathsf{P}\nu)\leqslant\Bigl{(}1+2\kappa_{\Psi}^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}\,d_{g}(\mu,\nu).$
@@
 $=\Bigl{(}1+2\kappa_{\Psi}^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}\Bigl{\lvert}\mu[g]-\nu[g]\Bigr{\rvert}\leqslant\Bigl{(}1+2\kappa_{\Psi}^{2}+\operatorname*{tr}(\Sigma)\Bigr{)}\,d_{g}(\mu,\nu),$
-which concludes the proof. ∎
+thus concluding the proof. ∎
 Lemma B.9 (Map $\mathsf{Q}$ is Lipschitz).
-Suppose that $h$ and $\Gamma$ satisfy Assumptions (H3) and (H5), respectively. Then it holds that
+Suppose that $\Gamma$ and $h$ satisfy Assumptions (H4) and (H2), respectively. Then it holds that
 $\forall(\mu,\nu)\in\mathcal{P}(^{d_{u}})\times\mathcal{P}(^{d_{u}}),\qquad d_{g}(\mathsf{Q}\mu,\mathsf{Q}\nu)\leqslant\Bigl{(}1+2\kappa_{h}^{2}+\operatorname*{tr}(\Gamma)\Bigr{)}d_{g}(\mu,\nu).$ (B.17)
@@
 Lemma B.10 (Stability of Map $\mathsf{B}_{j}$).
-Under Assumption H, for any probability measure $\mu\in\mathcal{P}(^{d_{u}})$ with $\mathcal{M}_{2}(\mu)<\infty$ there exists a constant $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}>0$ such that
+Suppose that Assumptions (H1), (H3), (H4) and (H2) are satisfied.
+Then for any probability measure $\mu\in\mathcal{P}(^{d_{u}})$ with $\mathcal{M}_{2}(\mu)<\infty$ there exists a constant $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}>0$ such that
 $\forall j\in\llbracket 0,J\rrbracket,\qquad d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)\leqslant C_{\mathsf{B}}d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
 Proof.
-For ease of exposition we write $y^{\dagger}=y^{\dagger}_{j+1}$. We define the $y$-marginal densities
+The proof of this lemma follows the steps of the proof of [carrillo2022ensemble, Lemma B.9]; however, each step requires different bounds reflecting the assumptions on $\Psi$ and $h$ in this paper.
+For ease of exposition we write $y^{\dagger}=y^{\dagger}_{j+1}$.
+We define the $y$-marginal densities
 $\alpha_{\mu}(y):=\int_{{}^{d_{u}}}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\,,\qquad\beta_{\mu}(y):=\int_{{}^{d_{u}}}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\,.$
-It holds that
+By definition of $\mathsf{B}_{j}$, it holds that
 $d_{g}(\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{B}_{j}\mathsf{Q}\mathsf{P}\mu)=\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\left|\frac{\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})}{\alpha_{\mu}(y^{\dagger})}-\frac{\mathsf{Q}\mathsf{P}\mu(u,y^{\dagger})}{\beta_{\mu}(y^{\dagger})}\right|\,\mathrm{d}u$
@@
 The distribution $\alpha_{\mu}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ is Gaussian with mean $\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)$ and covariance
-$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)=\Gamma+\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\otimes h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]}-\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]}\otimes\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]},$ (B.19)
-which by Assumption H is bounded from below by $\Gamma$ and from above by $\Gamma+2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}$.
-To deduce this we have again made use of the fact that $aa^{\mathsf{T}}\preccurlyeq(a^{\mathsf{T}}a)I_{d}$ for any vector $a\in^{d_{u}}$, by the Cauchy-Schwarz inequality.
+$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)=\Gamma+\mathsf{P}\mu\bigl{[}h\otimes h\bigr{]}-\mathsf{P}\mu\bigl{[}h\bigr{]}\otimes\mathsf{P}\mu\bigl{[}h\bigr{]}.$ (B.19)
+Clearly $\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)\succcurlyeq\Gamma$.
+Furthermore, it holds that $\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)\preccurlyeq\Gamma+2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}$ by (B.8).
 Therefore, noting that (B.3b) implies that $\left\lvert\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)\right\rvert\leqslant C$ for a constant $C$ depending only on the parameters $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma$, we obtain
@@
 $\geqslant\frac{\exp\Bigl{(}-\frac{1}{2}\bigl{(}|y|+C\bigr{)}^{2}\left\lVert\Gamma^{-1}\right\rVert\Bigr{)}}{\sqrt{(2\pi)^{d_{y}}\det\bigl{(}\Gamma+2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}\bigr{)}}}.$ (B.20)
-The function $\beta_{\mu}$ can be bounded from below using similar reasoning. Indeed, by Assumption H we have that for all $y\in^{d_{y}}$,
+The function $\beta_{\mu}$ can be bounded from below using similar reasoning. Indeed, by Assumptions (H3), (H4) and (H2) we have that for all $y\in^{d_{y}}$,
 $\beta_{\mu}(y)$ $=\int_{{}^{d_{u}}}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u=\int_{{}^{d_{u}}}\frac{\exp\left(-\frac{1}{2}\bigl{(}y-h(u)\bigr{)}^{\mathsf{T}}\Gamma^{-1}\bigl{(}y-h(u)\bigr{)}\right)}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma)}}\mathsf{P}\mu(u)\,\mathrm{d}u$
-$\geqslant\frac{C\exp\Bigl{(}-\left\lVert\Gamma^{-1}\right\rVert|y|^{2}-\left\lVert\Gamma^{-1}\right\rVert\kappa^{2}_{h}\int_{{}^{d_{u}}}|u|^{2}\mathsf{P}\mu\mathrm{d}u\Bigr{)}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma)}}$ (B.21a)
+$\geqslant\frac{1}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma)}}\exp\Bigl{(}-\left\lVert\Gamma^{-1}\right\rVert|y|^{2}-\left\lVert\Gamma^{-1}\right\rVert\kappa^{2}_{h}\int_{{}^{d_{u}}}\bigl{(}2+|u|^{2}\bigr{)}\mathsf{P}\mu(u)\,\mathrm{d}u\Bigr{)}$ (B.21a)
 $\geqslant C\exp\Bigl{(}-\left\lVert\Gamma^{-1}\right\rVert|y|^{2}\Bigr{)},$ (B.21b)
-where we applied Jensen’s inequality in (B.21a) and the constant in (B.21b) depends on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$.
+where we applied Jensen’s inequality in (B.21a) and the constant in (B.21b) depends on $\mathcal{M}_{2}(\mu),\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
 Step 2: bounding the first term in (B.18)
-We begin by noting that for fixed $u\in^{d_{u}}$, the functions $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)$ and $y\mapsto\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)$ are Gaussians up to constant factors.
-The covariance matrix of the former is $\Gamma$; it is possible to deduce that the covariance of the latter is given by
+First note that, for fixed $u\in^{d_{u}}$, the functions $\mathsf{Q}\mathsf{P}\mu(u,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ and $\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ are Gaussians up to constant factors, with covariance matrices given respectively by $\Gamma$ and
 $\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)-\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu),$ (B.22)
@@
 We note that $\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ is positive semi-definite, hence by (B.19) and its upper bound it follows that the matrix (B.22) is bounded from above by ${2}\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}+\Gamma$.
-With the same notation as in (B.8), we find that
-$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)-\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ (B.23)
-$\qquad=\Gamma+\Bigl{(}\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)-\mathcal{C}^{yu}(\phi_{\sharp}\mathsf{P}\mu)\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\phi_{\sharp}\mathsf{P}\mu)\Bigr{)}\succcurlyeq\Gamma;$ (B.24)
-here, to deduce that the term in brackets is positive semi-definite, we observe that the term is the Schur complement of the block $\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)$ of the matrix $\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$.
-Hence, (B.22) is bounded from below by $\Gamma$.
-Therefore the integral in the first term of (B.18) may be bounded from above by applying the result of Lemma A.2 in Appendix A with parameter $\alpha=\alpha\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma\bigr{)}$, yielding
-$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\left|\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\right|\mathrm{d}u$
-$\qquad\leqslant C\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
-$\qquad\leqslant C\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\left(1+|u|^{2}+|y|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
-$\qquad=Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu),$ (B.25)
-where $C$ depends on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$.
+As shown in the proof of [carrillo2022ensemble, Lemma B.9], the matrix (B.22) is bounded from below by $\Gamma$ (see [carrillo2022ensemble, Equation (B.20)]).
+It follows from using Lemma A.2 in Appendix A, with parameter $\tau=\tau\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma\bigr{)}$, that
+$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\left|\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\right|\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu),$ (B.23)
+where $C$ is a constant depending on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$.
+We refer to [carrillo2022ensemble, Equation (B.21)] for the detailed steps used to establish this bound.
 Step 3: bounding the second term in (B.18)
-By (B.25), it holds that
-$\bigl{\lvert}\alpha_{\mu}(y)-\beta_{\mu}(y)\bigr{\rvert}$ $=\left|\int_{{}^{d_{u}}}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\right|$
-$\leqslant\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
-Now, since $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)/\mathsf{P}\mu(u)$ is a Gaussian density with covariance $\Gamma$, which is bounded uniformly from above by $\bigl{(}(2\pi)^{d_{y}}\det(\Gamma)\bigr{)}^{-1/2}$, we also have that
-$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$ $=\beta_{\mu}(y)+\int_{{}^{d_{u}}}|u|^{2}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$
-$\leqslant\beta_{\mu}(y)+\int_{{}^{d_{u}}}\frac{|u|^{2}\mathsf{P}\mu(u)}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}\,\mathrm{d}u$
-$=\beta_{\mu}(y)+\frac{\operatorname*{tr}\bigl{(}\mathcal{C}(\mathsf{P}\mu)\bigr{)}+\left\lvert\mathcal{M}(\mathsf{P}\mu)\right\rvert^{2}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}.$
-We proceed by noting the result of Lemma B.1, which gives that the first and second moments of $\mathsf{P}\mu$ are bounded from above by a constant which depends on $\mathcal{M}_{2}(\mu),\kappa_{\Psi},\Sigma$.
+In view of (B.23), it holds that
+$\bigl{\lvert}\alpha_{\mu}(y)-\beta_{\mu}(y)\bigr{\rvert}\leqslant\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
+Now, since $\mathsf{Q}\mathsf{P}\mu(u,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})/\mathsf{P}\mu(u)$ defines a Gaussian density which has covariance $\Gamma$ and is bounded uniformly from above by $\bigl{(}(2\pi)^{d_{y}}\det(\Gamma)\bigr{)}^{-1/2}$, we also have that
+$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\leqslant\int_{{}^{d_{u}}}\frac{\left(1+|u|^{2}\right)\mathsf{P}\mu(u)}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}\,\mathrm{d}u=\frac{1+\operatorname*{tr}\bigl{(}\mathcal{C}(\mathsf{P}\mu)\bigr{)}+\left\lvert\mathcal{M}(\mathsf{P}\mu)\right\rvert^{2}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}.$
+By the moment bounds in Lemma B.1, the right-hand side is bounded from above by a constant which depends on $\mathcal{M}_{2}(\mu),\kappa_{\Psi},\Sigma$.
 Step 4: concluding the proof
@@
 Suppose that Assumption H is satisfied.
-Then, for all $R\geqslant 1$, it holds that for any $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, there is $L_{\mathsf{T}}:=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$, such that
+Then, for all $R\geqslant 1$, it holds that for any $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, there is $L_{\mathsf{T}}:=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$, such that
 $\forall j\in\llbracket 1,J\rrbracket,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p).$
 Proof.
-By the results of Lemma B.2, it holds that the probability measure $\mathsf{Q}\mathsf{P}\mu$ belongs to $\mathcal{P}_{\widetilde{R}}(^{d_{u}}\times^{d_{y}})$ for some $\widetilde{R}\geqslant 1$.
-Define
+By the results of Lemma B.2, it holds that $\mathsf{Q}\mathsf{P}\mu\in\mathcal{P}_{\widetilde{R}}(^{d_{u}}\times^{d_{y}})$ for some $\widetilde{R}(\mathcal{M}_{2}(\mu),\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma)\geqslant 1$.
+Using analogous notation to the one found in the proof of [carrillo2022ensemble, Lemma B.10] we define
 $r=\max\Bigl{\{}R,\widetilde{R},\kappa_{y}\Bigr{\}}.$
-We denote by $\mathscr{T}^{\pi}$ and $\mathscr{T}^{p}$ the affine maps corresponding to evaluation of covariance information at the probability measures $\pi$ and $p=\mathsf{Q}\mathsf{P}\mu$, respectively.
-Namely, letting $K=\mathcal{C}(\pi)$, $S=\mathcal{C}(p)$ and $y^{\dagger}=y^{\dagger}_{j+1}$, we write
+Letting $K=\mathcal{C}(\pi)$, $S=\mathcal{C}(p)$ and $y^{\dagger}=y^{\dagger}_{j+1}$, we also define the affine maps $\mathscr{T}^{\pi}$ and $\mathscr{T}^{p}$ corresponding to use of covariance information at the probability measures $\pi$ and $p=\mathsf{Q}\mathsf{P}\mu$,
 $\mathscr{T}^{\pi}(u,y)$ $=u+A_{\pi}(y^{\dagger}-y),\qquad$ $A_{\pi}:=K_{uy}K_{yy}^{-1},$
@@
 $d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant d_{g}(\mathscr{T}^{\pi}_{\sharp}\pi,\mathscr{T}^{\pi}_{\sharp}p)+d_{g}(\mathscr{T}^{\pi}_{\sharp}p,\mathscr{T}^{p}_{\sharp}p).$ (B.24)
-Before bounding each term on the right-hand side of B.26, we derive some inequalities that we will use in the proof. It holds that
-$\lVert A_{\pi}\rVert\leqslant\lVert K_{uy}\rVert\lVert K_{yy}^{-1}\rVert\leqslant\lVert K\rVert\lVert K^{-1}\rVert\leqslant r^{4},$ (B.27)
-and similarly for $A_{p}$.
-This can easily be seen by noting that the matrix norm (induced by the Euclidean vector norm) of any submatrix is bounded from above by the norm of the full matrix.
-Using this bound again and assuming without loss of generality that $r>1$, we deduce that
-$\lVert A_{\pi}-A_{p}\rVert=\lVert K_{uy}K_{yy}^{-1}-S_{uy}S_{yy}^{-1}\rVert$ $\leqslant\lVert(K_{uy}-S_{uy})K_{yy}^{-1}\rVert+\lVert S_{uy}\left(K_{yy}^{-1}-S_{yy}^{-1}\right)\rVert$
-$\leqslant r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}-S_{yy}^{-1}\rVert\Bigr{)}$
-$=r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}(S_{yy}-K_{yy})S_{yy}^{-1}\rVert\Bigr{)}$
-$\leqslant r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}\rVert\lVert S_{yy}-K_{yy}\rVert\lVert S_{yy}^{-1}\rVert\Bigr{)}$
-$\leqslant 2r^{6}\lVert K-S\rVert\leqslant 2r^{6}(1+2r)\,d_{g}(\pi,p),$ (B.28)
-where the last line follows from the result in Lemma B.6.
+In the following steps we will separately bound the two terms on the right hand side of (B.24).
+Before proceeding we outline two auxiliary bounds that will be used in the rest of the proof.
+These bounds are identical to the ones found in the proof of [carrillo2022ensemble, Lemma B.10]; we include statements here for expository purposes.
+Noting that the operator 2-norm of any submatrix is bounded from above by the operator 2-norm of the full matrix, we observe (see [carrillo2022ensemble, Equation (B.23)]) that
+$\lVert A_{\pi}\rVert\leqslant\lVert K_{uy}\rVert\lVert K_{yy}^{-1}\rVert\leqslant\lVert K\rVert\lVert K^{-1}\rVert\leqslant r^{4}.$ (B.25)
+Note that the above bound similarly holds for $A_{p}$. Using (B.25) and assuming without loss of generality that $r>1$, we deduce (see [carrillo2022ensemble, Equation (B.24)]) that
+$\lVert A_{\pi}-A_{p}\rVert\leqslant\lVert(K_{uy}-S_{uy})K_{yy}^{-1}\rVert+\lVert S_{uy}\left(K_{yy}^{-1}-S_{yy}^{-1}\right)\rVert\leqslant 2r^{6}\lVert K-S\rVert\leqslant 2r^{6}(1+2r)\,d_{g}(\pi,p),$ (B.26)
+where the second inequality follows again from fact that the 2-norm of any submatrix is bounded from above by the 2-norm of the full matrix, while the last inequality follows from the result in Lemma B.6.
 Bounding the first term in (B.24)
-Let $f\colon^{d_{u}}\to$ be a function such that $\lvert f\rvert\leqslant g$ with $g(u)=1+\lvert u\rvert^{2}$. It follows from the definition of the pushforward measure that
-$\bigl{\lvert}\mathscr{T}^{\pi}_{\sharp}\pi[f]-\mathscr{T}^{\pi}_{\sharp}p[f]\bigr{\rvert}=\bigl{\lvert}\pi[f\circ\mathscr{T}^{\pi}]-p[f\circ\mathscr{T}^{\pi}]\bigr{\rvert}.$
-For any $(u,y)\in^{d_{u}}\times^{d_{y}}$, we have that
-$\bigl{\lvert}f\circ\mathscr{T}^{\pi}(u,y)\bigr{\rvert}$ $=\left\lvert f\Bigl{(}u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}\right\rvert\leqslant g\Bigl{(}u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}$
-$=1+\bigl{\lvert}u+A_{\pi}(y^{\dagger}-y)\bigr{\rvert}^{2}\leqslant 1+3\lvert u\rvert^{2}+3\lvert A_{\pi}y^{\dagger}\rvert^{2}+3\lvert A_{\pi}y\rvert^{2}$
-$\leqslant 3\left(1+\lvert A_{\pi}y^{\dagger}\rvert^{2}\right)\max\left\{1,\lVert A_{\pi}\rVert^{2}\right\}g(u,y).$
-Hence, by (B.27) we conclude that
+With an identical argument to the one used in the proof of [carrillo2022ensemble, Lemma B.10] we have that
 $d_{g}(\mathscr{T}^{\pi}_{\sharp}\pi,\mathscr{T}^{\pi}_{\sharp}p)\leqslant 3\left(1+r^{10}\right)r^{8}\,d_{g}(\pi,p).$ (B.27)
 Bounding the second term in (B.24)
-Let $f\colon^{d_{u}}\to$ again be a function such that the inequality $\lvert f\rvert\leqslant g$, with $g(u)=1+\lvert u\rvert^{2}$, is satisfied. We note that
+Let $f$ again satisfy $\lvert f\rvert\leqslant g$. We note that
 $\Bigl{\lvert}\mathscr{T}^{\pi}_{\sharp}p[f]-\mathscr{T}^{p}_{\sharp}p[f]\Bigr{\rvert}=\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}]-p[f\circ\mathscr{T}^{p}]\Bigr{\rvert}=\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}.$
-The right-hand side may be rewritten as
-$\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}$ $=\left\lvert\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f\left(u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\right)-f\left(u+A_{p}\bigl{(}y^{\dagger}-y\bigr{)}\right)\,p(u,y)\,\mathrm{d}u\,\mathrm{d}y\right\rvert.$
-A change of variable yields
-$\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\bigl{(}f(u+A_{\pi}z)-f(u+A_{p}z)\bigr{)}\,p(u,y^{\dagger}-z)\,\mathrm{d}u\,\mathrm{d}z$
-$=\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f(v)\Bigl{(}p(v-A_{\pi}z,y^{\dagger}-z)-p(v-A_{p}z,y^{\dagger}-z)\Bigr{)}\mathrm{d}v\,\mathrm{d}z.$ (B.30)
-Since $\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty$ by assumption, from Lemma A.3 it follows that
+The last term may be expressed as
+$\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}$ $=\left\lvert\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\Bigl{(}f\left(u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\right)-f\left(u+A_{p}\bigl{(}y^{\dagger}-y\bigr{)}\right)\Bigr{)}\,p(u,y)\,\mathrm{d}u\,\mathrm{d}y\right\rvert$
+$=\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\bigl{(}f(u+A_{\pi}z)-f(u+A_{p}z)\bigr{)}\,p(u,y^{\dagger}-z)\,\mathrm{d}u\,\mathrm{d}z$
+$=\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f(v)\Bigl{(}p(v-A_{\pi}z,y^{\dagger}-z)-p(v-A_{p}z,y^{\dagger}-z)\Bigr{)}\,\mathrm{d}v\,\mathrm{d}z,$ (B.28)
+where we used a change of variables in the second equality.
+Since $\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty$ by assumption, it follows from Lemma A.3 and from the inequality $\min\{a,b\}\leqslant\sqrt{a}\sqrt{b}$ that there is a constant $C$ so that
 $\Bigl{\lvert}p(v-A_{\pi}z,y^{\dagger}-z)-p(v-A_{p}z,y^{\dagger}-z)\Bigr{\rvert}$
-$\qquad\leqslant C\lvert A_{\pi}z-A_{p}z\rvert\cdot\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\cdot$
-$\qquad\qquad\cdot\frac{1}{1+\lvert y^{\dagger}-z\rvert^{\max\{3+d_{u},4+d_{y}\}}}.$
-We apply this inequality to bound for fixed $z\in^{d_{y}}$ the inner integral in (B.30). Keeping only the terms that depend on $v$, we obtain that
-$\int_{{}^{d_{u}}}\lvert f(v)\rvert\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\,\mathrm{d}v$
-$\qquad\leqslant\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\mathrm{d}v+\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\mathrm{d}v.$
-Since $\lvert f(v)\rvert\leqslant 1+\lvert v\rvert^{2}$, proceeding by substitution gives that this expression is bounded from above by
+$\qquad\leqslant C\lvert A_{\pi}z-A_{p}z\rvert\cdot\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}$
+×$\qquad\qquad\times\frac{1}{1+\lvert y^{\dagger}-z\rvert^{\max\{3+d_{u},4+d_{y}\}}}.$
+We apply this inequality to bound for fixed $z\in^{d_{y}}$ the inner integral in (B.28). Considering only the terms that depend on $v$ gives
+$\int_{{}^{d_{u}}}\lvert f(v)\rvert\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\,\mathrm{d}v$ (B.29)
+$\qquad\leqslant\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v+\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v$
+$\qquad\leqslant\int_{{}^{d_{u}}}\frac{1+\lvert v\rvert^{2}}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v+\int_{{}^{d_{u}}}\frac{1+\lvert v\rvert^{2}}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v$
+$\qquad\leqslant\int_{{}^{d_{u}}}\frac{1+\lvert w+A_{\pi}z\rvert^{2}}{1+\lvert w\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}w+\int_{{}^{d_{u}}}\frac{1+\lvert w+A_{p}z\rvert^{2}}{1+\lvert w\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}w,$
+where we used that $\lvert f(v)\rvert\leqslant 1+\lvert v\rvert^{2}$, as well as a change of variable in the last line. It follows that the integral in (B.29) is bounded from above by
 $C\Bigl{(}1+\lvert A_{\pi}z\rvert^{2}+\lvert A_{p}z\rvert^{2}\Bigr{)}\leqslant Cr^{8}\left(1+\lvert z\rvert^{2}\right),$
-where the last inequality follows from (B.27). The remaining integral in the $z$ variable can be bounded in a similar way:
+where the inequality follows from (B.25). Finally, the resulting integral in the $z$ variable can be bounded analogously, which gives
 $d_{g}(\mathscr{T}^{\pi}_{\sharp}p,\mathscr{T}^{p}_{\sharp}p)$ $\leqslant Cr^{8}\int_{{}^{d_{y}}}\frac{1+\lvert z\rvert^{2}}{1+\lvert y^{\dagger}-z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\lvert A_{\pi}z-A_{p}z\rvert\,\mathrm{d}z$
@@
 Proof.
-The lemma as stated may be found in [carrillo2022ensemble, Lemma 28], where a complete proof is given. ∎
+The lemma as stated may be found in [carrillo2022ensemble, Lemma B.12], where a complete proof is given. ∎
 Lemma C.2.
@@
 We use such results as a component in proving Theorem 2.4 and, in so-doing, explicit dependence on moments must be tracked.
-In this section we detail proofs of the result in  [le2009large, Proposition 4.4] and [le2009large, Theorem 5.2] , following the analysis closely and, in addition, tracking dependence on moments.
+In this section we give a self-contained proof of [le2009large, Theorem 5.2], following the analysis closely [^1] The result of [le2009large] holds more generally for functions $\Psi$ that are locally Lipschitz and that grow at most polynomially at infinity [le2009large, Assumption B].
+However, our proof uses that $\Psi$ is globally Lipschitz, for instance in (D.12).
+and, in addition, tracking dependence on moments; this moment dependence may be useful in the context of future work generalizing what we do in this paper.
 This leads to the following error estimate stating the desired Monte Carlo error estimate.
@@
 Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H2).
-We assume that the vector field $h$ is a linear transformation, i.e.
-that Assumption (V2) is satisfied.
-Furthermore, if the vector field $\Psi$ additionally satisfies $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, then for all $\phi$ satisfying Assumption P1, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu^{\rm EK}_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\Sigma,\Gamma\bigr{)}$ where $q\coloneqq\max\{16\cdot 3^{J-1},8(\varsigma+1)3^{J-1}\}$ such that
-$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}}.$
+We assume that the vector field $\Psi$ satisfies Assumption (H3) and that $h$ is a linear transformation, i.e.
+that Assumption (V2) is satisfied and hence also Assumption (H4).
+Furthermore, if the vector field $\Psi$ additionally satisfies $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, then for all $\phi$ satisfying Assumption P1, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu^{{\rm EK}}_{0}),J,R_{\phi},L_{\phi},\varsigma,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\Sigma,\Gamma\bigr{)}$ where $q\coloneqq\max\{4^{J+1},4\varsigma,2(\varsigma+1)\}$ such that
+$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}\leqslant\frac{C}{\sqrt{N}}.$
 Proof.
 To prove the proposition, we apply the coupling argument used in [le2009large].
-Using similar notation to the one in [le2009large], to the interacting $N$-particle system $\{u^{(i)}_{j}\}_{n=1}^{N}$ evolving according to the ensemble Kalman dynamics (2.5) , we couple $N$ copies of the mean field dynamics $\{\overline{u}^{(j)}_{j}\}_{n=1}^{N}$ evolving according to the mean field ensemble Kalman dynamics (2.1).
+Using similar notation to the one in [le2009large], to the interacting $N$-particle system $\bigl{\{}u^{(i)}_{j}\bigr{\}}_{n=1}^{N}$ evolving according to the ensemble Kalman dynamics (2.5), we couple $N$ copies of the mean field dynamics $\bigl{\{}\overline{u}^{(j)}_{j}\bigr{\}}_{n=1}^{N}$ evolving according to the mean field ensemble Kalman dynamics (2.1).
 The mean field replicas are synchronously coupled to the interacting particle system, in the sense that they are initialized at the same initial condition and driven by the same noises; namely, the two particle systems are initialized at i.i.d.
@@
 Furthermore, we add a bar $\overline{\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}}$ to all the variables related to the synchronously coupled mean field particles, including the probability measures $\overline{\pi}^{{\rm EK}}_{j}$.
-We define for $\pi\in\mathcal{P}(^{d_{u}\times d_{y}})$ the Kalman gain
+We also define for $\pi\in\mathcal{P}(^{d_{u}\times d_{y}})$ the Kalman gain
 $\mathcal{K}(\pi):=\operatorname{\mathcal{C}}^{uh}\left(\pi\right)\left(\operatorname{\mathcal{C}}^{hh}\left(\pi\right)+\Gamma\right)^{-1}.$
@@
 […]
-Note that the synchronously coupled particles are independent and identically distributed. With this set-up we proceed with the proof. We first define
-$D_{j}^{N,p}\coloneqq\Bigl{(}\mathbb{E}\left\lvert u^{(i)}_{j}-\overline{u}^{(i)}_{j}\right\rvert^{p}\Bigr{)}^{1/p}.$
-We now let $\phi$ satisfy Assumption P1. By applying the triangle inequality, we deduce that
-$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{1/2}+\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}.$ (D.3)
-The proof of [le2009large, Theorem 5.2] uses the law of large numbers in $L^{2}$ to establish the bound for the second term of the right hand side of D.3
-$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{J}^{\rm EK}\bigr{)},R_{\phi}\Bigr{)}}{\sqrt{N}};$ (D.4)
-furthermore, in the proof of [le2009large, Theorem 5.2], the following bound on the first term of the right hand side of D.3 may be deduced:
-$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{1/2}\leqslant L_{\phi}D_{J}^{N,4}\cdot\Bigl{(}1+\bigl{|}\mathcal{M}^{1/4}_{4\varsigma}\bigl{(}\mu_{J}^{\rm EK}\bigr{)}\bigr{|}\Bigr{)}+L_{\phi}|D_{J}^{N,2(\varsigma+1)}|^{\varsigma+1}.$ (D.5)
-Under the assumptions of the proposition, the result of [le2009large, Proposition 4.4] shows by induction that for any $j\in\mathbb{N}$, it holds that  [^1] The result of [le2009large, Proposition 4.4] holds more generally for functions $\Psi$ that are locally Lipschitz and that grow at most polynomially at infinity [le2009large, Assumption B].
-However, to obtain (D.6) from the result of [le2009large, Proposition 4.4] we use that $\Psi$ is assumed to be globally Lipschitz.
-$D_{j}^{N,p}\leqslant CD_{j-1}^{N,p}$ $+CD_{j-1}^{N,2\vee(pr)}\Bigl{(}D_{j-1}^{N,2\vee(pr^{\prime})}+\bigl{(}\mathcal{M}\bigl{(}\mathsf{P}\mu_{j-1}^{\rm EK}\bigr{)}\bigr{)}^{2\vee(pr^{\prime})}\Bigr{)}\cdot\Bigl{(}D_{j-1}^{N,2\vee(pr^{\prime\prime})}+\overline{S}_{j}^{pr^{\prime\prime}}\Bigr{)}$ (D.6)
-$\quad+C\Bigl{(}\mathbb{E}\bigl{|}Z_{j}^{N}\bigr{|}^{pq}\Bigr{)}^{1/pq}\cdot\bigl{(}D_{j-1}^{N,pq^{\prime}}+\overline{S}_{j}^{pq^{\prime}}\bigr{)},$
-for any $j\in\mathbb{N}$, where $C$ is a constant depending on $\Sigma,\Gamma,\ell_{\Psi},\kappa_{\Psi},\kappa_{h},\mathcal{M}_{2}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)}$.
-We note that the specific form of $C$ may be deduced using the proof of [le2009large, Lemma 4.1].
-Furthermore, we note that
-$\frac{1}{q}+\frac{1}{q^{\prime}}=1,\qquad\frac{1}{r}+\frac{1}{r^{\prime}}+\frac{1}{r^{\prime\prime}}=1,$
-and that $\overline{S}_{j}^{t}$ and $Z_{j}^{N}$ are defined as
-$\overline{S}_{j}^{t}$ $\coloneqq\Bigl{(}\mathbb{E}\bigl{|}{y}^{\dagger}_{j}-h\Bigl{(}\overline{v}^{(i)}_{j}\Bigr{)}-\eta^{(i)}_{j}\bigr{|}^{t}\Bigr{)}^{1/t},$
-$Z_{j}^{N}$ $\coloneqq\left\|\mathcal{C}\Bigl{(}\frac{1}{N}\sum_{i=1}^{N}\delta_{\overline{v}^{(i)}_{j}}\Bigr{)}-\mathcal{C}\bigl{(}\mathsf{P}\mu_{j-1}^{\rm EK}\bigr{)}\right\|.$
-By applying the triangle inequality it is straightforward to deduce that
-$\overline{S}_{j}^{t}\leqslant\left(\int_{{}^{d_{u}}}\bigl{|}{y}^{\dagger}_{j}-h(u)\bigr{|}^{t}\mathsf{P}\mu_{j-1}^{\rm EK}(\mathrm{d}u)\right)^{1/t}+\bigl{(}\mathbb{E}|\eta_{j}|^{t}\bigr{)}^{1/t}\leqslant C\Bigl{(}\mathcal{M}_{t}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)},\kappa_{\Psi},\kappa_{h},\kappa_{y},\Sigma,\Gamma\Bigr{)},$ (D.7)
-for any $t\in\mathbb{N}$. On the other hand, [le2009large, Lemma 3.3] shows that
-$\Bigl{(}\mathbb{E}\bigl{|}Z_{j}^{N}\bigr{|}^{t}\Bigr{)}^{1/t}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2t}\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)},\kappa_{\Psi},\Sigma\Bigr{)}}{\sqrt{N}},$ (D.8)
-for any $N\in\mathbb{N}$ and any $t\in\mathbb{N}$.
-Combining the results from (D.7) and (D.8) with the one in (D.6), and choosing for convenience $q=q^{\prime}=2$ and $r=r^{\prime}=r^{\prime\prime}=3$ for any $j\in\mathbb{N}$ we obtain
-$D_{j}^{N,p}\leqslant CD_{j-1}^{N,p}+CD_{j-1}^{N,3p}\Bigl{(}D_{j-1}^{N,3p}+1\Bigr{)}\cdot\Bigl{(}D_{j-1}^{N,3p}+1\Bigr{)}+\frac{C}{\sqrt{N}}\cdot\bigl{(}D_{j-1}^{N,2p}+1\bigr{)},$ (D.9)
-for a constant $C$ depending on $\mathcal{M}_{4p}{\bigl{(}\mu_{j-1}^{\rm EK}\bigr{)}},\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$.
-Since $D_{0}^{N,p}=0$ for any $p\in\mathbb{N}$, by iterating in $j$ the result in (D.9), we find that
-$D_{J}^{N,4}\leqslant\frac{C}{\sqrt{N}},$
-where $C$ is a constant depending on $\mathcal{M}_{16\cdot 3^{J-1}}{\bigl{(}\mu_{0}^{\rm EK}\bigr{)}},J,\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$.
-Hence, by combining the results of (D.5) and (D.4) with the decomposition (D.3), we deduce that
-$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}},$
-where $C$ is a constant depending on $\mathcal{M}_{q}{\bigl{(}\mu_{0}^{\rm EK}\bigr{)}},J,R_{\phi},L_{\phi},\varsigma,\kappa_{\Psi},\kappa_{h},\kappa_{y},\ell_{\Psi},\Sigma,\Gamma$ for a constant $q$ defined as $q\coloneqq\max\{16\cdot 3^{J-1},8(\varsigma+1)3^{J-1}\}$.
+Note that the synchronously coupled particles are independent and identically distributed.
+With this set-up we proceed with the proof.
+Let $\phi$ satisfy Assumption P1.
+By applying the triangle inequality, we deduce that
+$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}$ $\leqslant\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$
+$\qquad+\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}.$ (D.3)
+Step 1: bounding the second term in (D.3)
+Noting that the random variables $\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu_{J}^{{\rm EK}}[\phi]$ are i.i.d. and expanding the square, we obtain
+$\left(\mathbb{E}~{}\biggl{\lvert}\sum_{i=1}^{N}\frac{1}{N}\Bigl{(}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{{\rm EK}}_{J}[\phi]\Bigr{)}\biggr{\rvert}^{2}\right)^{\frac{1}{2}}=\frac{1}{\sqrt{N}}\Bigl{(}\mathbb{E}\bigl{|}\phi\bigl{(}\overline{u}^{(1)}_{J}\bigr{)}-\mu_{J}^{{\rm EK}}[\phi]\bigr{|}^{2}\Bigr{)}^{\frac{1}{2}},$
+Since by Assumption P1 it holds that $|\phi(u)|\leqslant R_{\phi}\bigl{(}1+|u|^{{\varsigma}+1}\bigr{)}$ for any $u\in^{d_{u}}$, it follows that
+$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\bigl{(}\overline{u}^{(i)}_{J}\bigr{)}-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}\leqslant\frac{C\Bigl{(}\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},R_{\phi},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}}{\sqrt{N}},$ (D.5)
+where we used Lemma B.5 to bound $\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{J}^{{\rm EK}}\bigr{)}$ in terms of $\mathcal{M}_{2(\varsigma+1)}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)}$.
+Step 2: bounding the first term in (D.3)
+For the first term, by Jensen’s inequality, exchangeability, and finally by Assumption P1 on $\phi$ together with the Cauchy–Schwarz inequality, it holds without loss of generality that
+$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\Bigl{(}u^{(i)}_{J}\Bigr{)}-\phi\Bigl{(}\overline{u}^{(i)}_{J}\Bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$ $\leqslant\left(\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\Bigl{|}\phi\Bigl{(}u^{(i)}_{J}\Bigr{)}-\phi\Bigl{(}\overline{u}^{(i)}_{J}\Bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$
+$=\left(\mathbb{E}\Bigl{|}\phi\Bigl{(}u^{(1)}_{J}\Bigr{)}-\phi\Bigl{(}\overline{u}^{(1)}_{J}\Bigr{)}\Bigr{|}^{2}\right)^{\frac{1}{2}}$
+$\leqslant 3L_{\phi}\left(\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{4}\right)^{\frac{1}{4}}\left(\mathbb{E}\left[1+\left\lvert u^{(1)}_{J}\right\rvert^{4\varsigma}+\left\lvert\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\right]\right)^{\frac{1}{4}}.$ (D.6)
+By Lemma B.5, there exists $C=C\Bigl{(}\mathcal{M}_{4\varsigma}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ so that
+$\mathbb{E}\left\lvert\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\leqslant C\quad\text{and}\quad\mathbb{E}\left\lvert u^{(1)}_{J}\right\rvert^{4\varsigma}\leqslant 2^{4\varsigma-1}\left(\mathbb{E}\left\lvert\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}+\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\right)\leqslant 2^{4\varsigma-1}\left(C+\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{4\varsigma}\right).$
+Hence it remains to bound terms of the form $\mathbb{E}\bigl{\lvert}u^{(1)}_{J}-\overline{u}^{(1)}_{J}\bigr{\rvert}^{p}$ with $p=4$ and $p=4\varsigma$.
+Such a bound will follow from a propagation of chaos result; which will be the object of Step 4.
+In Step 3 we show auxiliary moment bounds for the mean field dynamics.
+Step 3: moment bounds for the mean field dynamics
+By Lemma B.5, for any $j\in\llbracket 0,J-1\rrbracket$ there exists a constant $C=C\Bigl{(}\mathcal{M}_{p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ so that
+$\left(\mathbb{E}\left\lvert\overline{u}_{j}^{(1)}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C.$ (D.7)
+From this it immediately follows, by Assumptions (H3) and (H4), that there exist constants $C_{v},C_{y}$ depending on $\mathcal{M}_{p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$ so that
+$\left(\mathbb{E}\left\lvert\overline{v}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{v},\qquad\left(\mathbb{E}\left\lvert\overline{y}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{y}.$ (D.8)
+Furthermore, it holds that for the Kalman gain that
+$\Bigl{\lVert}\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\Bigr{\rVert}^{p}\leqslant\left\lVert\mathcal{C}^{uh}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\right\rVert^{p}\left\lVert\Gamma^{-1}\right\rVert^{p}=\left\lVert\Gamma^{-1}\right\rVert^{p}\left\lVert\mathcal{C}^{uu}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}H^{\mathsf{T}}\right\rVert^{p}\leqslant C,$ (D.9)
+where $C$ depends on $\mathcal{M}_{2}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
+Step 4: propagation of chaos
+In this step, we prove a propagation of chaos result stating for all $p$, there exists a constant $C_{p}=C_{p}\Bigl{(}\mathcal{M}_{4^{J}\cdot p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma\Bigr{)}$ such that
+$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert u^{(1)}_{J}-\overline{u}^{(1)}_{J}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{p}.$ (D.10)
+To prove this result, we follow the strategy in [le2009large] and reason by induction.
+Fix $j\in\llbracket 0,J-1\rrbracket$ and assume that for all $p\in\mathbb{N}$ there is $C_{j,p}=C_{j,p}\Bigl{(}\mathcal{M}_{4^{j}\cdot p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma\Bigr{)}$ such that
+$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert u^{(1)}_{j}-\overline{u}^{(1)}_{j}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant C_{j,p}.$ (D.11)
+By Lipschitz continuity of $\Psi$ and $h$, we deduce immediately that
+$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert v^{(1)}_{j+1}-\overline{v}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant\ell_{\Psi}C_{j,p},\qquad\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lvert y^{(1)}_{j+1}-\overline{y}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}\leqslant\ell_{h}\ell_{\Psi}C_{j,p}.$ (D.12)
+Now we write
+$u^{(1)}_{j+1}-\overline{u}^{(1)}_{j+1}$ $=\left(v^{(1)}_{j+1}-\overline{v}^{(1)}_{j+1}\right)+\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\left(\overline{y}^{(1)}_{j+1}-y^{(1)}_{j+1}\right)$
+$\qquad+\Bigl{(}\mathcal{K}\bigl{(}\pi^{{\rm EK},N}_{j+1}\bigr{)}-\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\Bigr{)}\left(y^{\dagger}_{j+1}-y^{(1)}_{j+1}\right).$
+Thus, by the triangle and Hölder inequalities,
+$\sqrt{N}\left(\mathbb{E}\left\lvert u^{(1)}_{j+1}-\overline{u}^{(1)}_{j+1}\right\rvert^{p}\right)^{\frac{1}{p}}$ $\leqslant\ell_{\Psi}C_{j,p}+\left\lVert\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\right\rVert\ell_{h}\ell_{\Psi}C_{j,p}$
+$\qquad+\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\bigl{(}\pi^{{\rm EK},N}_{j+1}\bigr{)}-\mathcal{K}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}\left(\mathbb{E}\left\lvert y^{\dagger}_{j+1}-y^{(1)}_{j+1}\right\rvert^{2p}\right)^{\frac{1}{2p}}.$
+The first two terms on the right-hand side are bounded uniformly in $N$ in view of (D.9).
+In order to bound the last term, we note that the term $\mathbb{E}\bigl{\lvert}y^{\dagger}_{j+1}-y^{(1)}_{j+1}\bigr{\rvert}^{2p}$ is bounded uniformly in $N$ by a constant which depends on $\mathcal{M}_{4^{j}\cdot 2p}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)}$; this follows by (D.8) and (D.12).
+To complete the inductive step, it remains to show that
+$\sup_{N\in\mathbb{N}}\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}<\infty.$
+To this end, in line with the classical propagation of chaos approach, we decompose
+$\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}$ $\leqslant\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}$
+$\qquad+\sqrt{N}\left(\mathbb{E}\left\lVert\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}},$ (D.13)
+where we introduced the empirical measure associated with the mean field particles:
+$\overline{\pi}^{{\rm EK},N}_{j+1}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{\left(\overline{v}^{(i)}_{j+1},\overline{y}^{(i)}_{j+1}\right)}.$
+To conclude the proof of propagation of chaos, we must prove that
+  • The first term on the right-hand side of (D.13) is bounded uniformly in $N$, which is achieved by employing the induction hypothesis (D.11), as well as (D.12).
+  • The second term on the right-hand side of (D.13) is also bounded uniformly in $N$. This follows from classical arguments based on the law of large number in $L^{p}$ spaces.
+Let us first bound the second term. To this end, note that
+$\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}$ $=\operatorname{\mathcal{C}}^{uh}\left(\overline{\pi}^{{\rm EK}}_{j+1}\right)\left(\left(\Gamma+\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right)^{-1}-\left(\Gamma+\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right)^{-1}\right)$
+$\qquad+\left(\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right)\left(\Gamma+\operatorname{\mathcal{C}}^{hh}\left(\overline{\pi}^{{\rm EK},N}_{j+1}\right)\right)^{-1}.$
+Therefore, using that $A^{-1}-B^{-1}=B^{-1}(B-A)A^{-1}$, we have that
+$\left\lVert\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert$ $\leqslant\left\lVert\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert\lVert\Gamma^{-1}\rVert^{2}\left\lVert\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{hh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert$
+$\qquad+\lVert\Gamma^{-1}\rVert\left\lVert\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uh}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert$
+$\leqslant C(\Gamma,H)\Bigl{(}1+\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert\Bigr{)}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert.$
+The term $\bigl{\lVert}\operatorname{\mathcal{C}}^{uu}\bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\bigr{)}\bigr{\rVert}$ is bounded by (D.8), and the $L^{p}$ norm of the second term tends to zero with rate $N^{-\frac{p}{2}}$ by classical law of large number arguments, see for example [MR4234152, Lemma 3] and [vaes2024, Lemma 3]; indeed it holds that
+$\mathbb{E}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK}}_{j+1}\Bigr{)}\right\rVert^{2p}\leqslant\frac{C\Bigl{(}\mathcal{M}_{4p}\bigl{(}\mu^{{\rm EK}}_{0}\bigr{)}\Bigr{)}}{N^{p}}.$
+It remains to bound the other term. Reasoning similarly to above, we have that
+$\left\lVert\mathcal{K}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\mathcal{K}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert\leqslant C(\Gamma,H)\Bigl{(}1+\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert\Bigr{)}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert.$ (D.14)
+By [vaes2024, Lemma 2], it holds that
+$\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert^{2}\leqslant 2\left(\frac{1}{N}\sum_{i=1}^{N}\left\lvert v^{(i)}_{j+1}-\overline{v}^{(i)}_{j+1}\right\rvert^{2}\right)\left(\frac{1}{N}\sum_{i=1}^{N}\left\lvert v^{(i)}_{j+1}\right\rvert^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\lvert\overline{v}^{(i)}_{j+1}\right\rvert^{2}\right).$
+Using the Cauchy-Schwarz and Jensen’s inequalities, moment bounds and finally the induction hypothesis (D.10) enables to conclude that
+$\left(\mathbb{E}\left\lVert\operatorname{\mathcal{C}}^{uu}\Bigl{(}\pi^{{\rm EK},N}_{j+1}\Bigr{)}-\operatorname{\mathcal{C}}^{uu}\Bigl{(}\overline{\pi}^{{\rm EK},N}_{j+1}\Bigr{)}\right\rVert^{2p}\right)^{\frac{1}{2p}}\leqslant\frac{C\Bigl{(}\mathcal{M}_{4^{j+1}p}\bigl{(}\mu^{{\rm EK}}_{0}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma\Bigr{)}}{\sqrt{N}}.$
+Step 5: concluding the proof
+Using the propagation of chaos result from (D.10) in (D.6) gives that
+$\left(\mathbb{E}~{}\Bigl{|}\frac{1}{N}\sum_{i=1}^{N}\phi\left(u^{(i)}_{J}\right)-\phi\left(\overline{u}^{(i)}_{J}\right)\Bigr{|}^{2}\right)^{\frac{1}{2}}\leqslant\frac{C}{\sqrt{N}},$
+for some constant $C$ depending on $\mathcal{M}_{q}\bigl{(}\mu^{{\rm EK}}_{0}\bigr{)},J,L_{\phi},\varsigma,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\ell_{h},\Sigma,\Gamma$ with $q=\max\{4^{J+1},4\varsigma\}$.
+Combining this result with (D.5) yields the desired result.
 ∎
-[…]